### v3

# Computational Complexity of Physical Information Sufficiency

Abstract
We study physical information sufficiency as a decision-theoretic meta-problem. For $\mathcal{D}=(A,S,U)$ with factored state space $S=X_{1}\times\cdots\times X_{n}$, a coordinate set $I$ is sufficient when
$s_{I}=s^{\prime}_{I}\implies\text{Opt}(s)=\text{Opt}(s^{\prime}).$
This asks whether projected information preserves the full optimal-action correspondence.
Complexity landscape.
  • SUFFICIENCY-CHECK is coNP-complete; MINIMUM-SUFFICIENT-SET is coNP-complete; ANCHOR-SUFFICIENCY is $\Sigma_{2}^{\textsf{P}}$-complete.
  • Under explicit-state encoding, polynomial-time algorithms hold for bounded actions, separable utility, and tree-structured utility.
  • Under succinct encoding, hardness is regime-dependent: with ETH, there are worst-case families with $k^{*}=n$ requiring $2^{\Omega(n)}$ time.
  • Under query access, the finite-state core has worst-case Opt-oracle complexity $\Omega(|S|)$, with Boolean value-entry and state-batch refinements preserving the obstruction.
Physical grounding. The paper formalizes a physical-to-core encoding $E:\mathcal{P}\to\mathcal{D}$ and a transport rule: declared physical assumptions transfer to core assumptions, and core claims lift back to encoded physical instances. Encoded physical counterexamples induce core failures on the encoded slice. Discrete-time interface semantics (decision event = one tick) and budgeted thermodynamic lifts (bit lower bounds to energy/carbon lower bounds under declared constants) are formalized in the same assumption-typed framework. [C:P2.34,2.36,2.28,2.29,2.30,2.31,5.18,5.28;C2.35,5.22;T2.37;R:AR] (L: CT6,CT8,CT4,CT1,DQ548,DQ537,DQ539,DQ541,PBC1,CC65,DQ569.)
All theorem-level claims are machine-checked in Lean 4 (13391 lines, 597 theorem/lemma statements). Complexity-class completeness follows by composition with standard complexity results; regime-dependent and physical-transport consequences are proved as assumption-explicit closures.
Keywords: computational complexity, physical information, decision theory, polynomial hierarchy, formal verification
© 2026 Tristan Simas. This work is licensed under CC BY 4.0. License: https://creativecommons.org/licenses/by/4.0/
1 Introduction
Consider any decision problem $\mathcal{D}$ with actions $A$ and factored state space $S=X_{1}\times\cdots\times X_{n}$. A coordinate set $I\subseteq\{1,\ldots,n\}$ is sufficient if knowing only coordinates in $I$ determines the optimal action:
$s_{I}=s^{\prime}_{I}\implies\text{Opt}(s)=\text{Opt}(s^{\prime})$
This is coordinate sufficiency, the formalization of information sufficiency (whether available information determines the optimal action) over a factored state space.
In information-theoretic terms, this is a task-conditional coding problem: the object preserved by projection is the decision boundary $s\mapsto\text{Opt}(s)$, not the full source state. The same object appears in classical source-coding and fidelity viewpoints [57, 59, 14, 64], and in zero-error/confusability structure where exact recoverability depends on graph constraints rather than raw description length [58, 31, 38, 15]. The same separation also appears in model selection: short descriptions and certifiable adequacy are distinct [54, 24, 35, 8].
For physical instantiations, the same tuple has a direct substrate reading: coordinates are measurable state degrees of freedom, actions are physically realizable interventions or transitions, and the optimizer map is the decision boundary induced by the declared objective contract. The physical-to-core encoding discipline in Section 2.7 makes this bridge theorem-level rather than metaphorical.[C:P2.34,2.36,2.28,2.29,2.30,2.31;C2.35;T2.37;R:AR] (L: CT6,CT8,CT4,CT1,DQ548,DQ537,DQ539,DQ541.)
Compact claim-stamp notation.
We use compact provenance stamps of the form $C:\cdots;R:\cdots]$: claim references are grouped by type ($D/P/C/T/L/R$ for Definition/Proposition/Corollary/Theorem/Lemma/Remark), and regime tags are abbreviated (for example ARAR=active declared regime, DCDC=declared contract, CRCR=declared class/regime, QfQf=query-finite, QbQb=query-Boolean). Theorem-strength narrative statements are stamped and, where available, paired with compact Lean handle IDs; unstamped narrative remains explanatory or citation-only context.
SUFFICIENCY-CHECK, MINIMUM-SUFFICIENT-SET, and ANCHOR-SUFFICIENCY are meta-problems: each is parameterized by an arbitrary decision problem $\mathcal{D}$ and asks a structural question about $\mathcal{D}$’s coordinate space. The complexity results therefore apply to every problem with factored state space, not to a specific problem domain.[C:T4.6,4.7,4.9;R:S] (L: CC61,CC36,CC5.)
Remark 1.1 (Universality of the Decision-Problem Tuple).
The formal object of this paper is the decision-problem tuple $\mathcal{D}=(A,S,U)$: actions, factored states, utility. Within this model family, any problem requiring a choice among alternatives given structured information instantiates the tuple: the alternatives are $A$, the structured information is $S=X_{1}\times\cdots\times X_{n}$, and the objective is $U$. Thus the qualifier “decision problem” does not narrow scope inside the declared coordinate model. The meta-problem results below are universal over problems representable with this coordinate structure; the $(A,S,U)$ formalism supplies the language for stating the sufficiency predicate.
This paper characterizes the complexity landscape of these meta-problems within the formal model:
The paper keeps two layers explicit throughout: a substrate-neutral decision core (for complexity classification) and an assumption-typed physical instantiation layer (for transport to concrete systems). This separation prevents conflating representational hardness with physical assumptions while preserving a formal bridge between them.[C:P2.34;T2.37;R:AR] (L: CT6,CT4.)
Section 2.4 fixes the computational model and input encodings used for all complexity claims. Section 2.5 gives the model contract and regime tags used to type every strong claim. Completeness claims are class-internal: complete for the static sufficiency class fixed by C1–C4 and the declared access regimes. Adjacent objective classes that violate the bridge conditions are typed separately (Section 2.5).
Problem Complexity When Tractable
SUFFICIENCY-CHECK coNP-complete Bounded actions (explicit-state), separable utility, tree-structured utility
MINIMUM-SUFFICIENT-SET coNP-complete Same conditions
ANCHOR-SUFFICIENCY $\Sigma_{2}^{\textsf{P}}$-complete Open
The tractable cases are stated with explicit encoding assumptions (Section 2.4). Outside those regimes, the succinct model yields hardness.
1.1 Hardness–Recovery Reading Map
Theorems in this paper are intended to be read as hardness/recovery pairs for the same fixed decision question:
  1. 1.
    Sufficiency checking: [S] coNP-complete in the general succinct regime (Theorem 4.6), paired with polynomial-time recovery under explicit-state and structured-access regimes (Theorem 6.1).
  2. 2.
    Minimal sufficient sets: [S] coNP-complete in general (Theorem 4.7), paired with the collapse criterion and relevance-computation recovery route (Theorems 4.14, 6.1).
  3. 3.
    Anchor sufficiency: $\Sigma_{2}^{\textsf{P}}$-complete in the general anchored formulation (Theorem 4.9); no general polynomial-time recovery condition is established in this model.
(L: CC60, CC36, CC34, CC4, CC71, S2P2.)
An operational criterion follows later from the same chain: once decision-site count exceeds the amortization threshold $n^{*}$, avoiding structural identification is more expensive than paying the one-time centralization cost (Theorem 9.6).[C:T9.6;R:LC] (L: HD3, HD21.)
1.2 Landscape Summary
When is sufficiency checking tractable? We identify three sufficient conditions:
  1. 1.
    Bounded actions ($|A|\leq k$) under explicit-state encoding: with constantly many actions, we enumerate action pairs over the explicit utility table.
  2. 2.
    Separable utility ($U(a,s)=f(a)+g(s)$): The optimal action depends only on $f$, making all coordinates irrelevant to the decision.
  3. 3.
    Tree-structured utility: With explicit local factors over a tree, dynamic programming yields polynomial algorithms in the input length.
Each condition is stated with its encoding assumption. Outside these regimes, the general problem remains coNP-hard (Theorem 4.6). [C:T4.6,6.1;R:E,S] (L: CC61,CC69,CC70,CC71.)
When is it intractable? The general problem is coNP-complete (Theorem 4.6), with a separation between explicit-state tractability and succinct worst-case hardness:
  • In the explicit-state model: SUFFICIENCY-CHECK is solvable in polynomial time in $|S|$ by explicitly computing $\text{Opt}(s)$ for all $s\in S$ and checking all pairs $(s,s^{\prime})$ with equal $I$-projection. In particular, instances with $k^{*}=O(\log|S|)$ are tractable in this model.
  • In the intermediate query-access model: the finite-state core yields an Opt-oracle lower bound $\Omega(|S|)$ (Boolean instantiation: $\Omega(2^{n})$), and the mechanized Boolean interface refinements preserve the obstruction for value-entry and state-batch interfaces, with explicit subproblem-to-full transfer, weighted randomized robustness, and oracle-lattice transfer/strictness theorems (Propositions 5.35.16).
  • In the succinct model: under ETH there exist worst-case instances produced by our strengthened gadget in which the minimal sufficient set has size $\Omega(n)$ (indeed $n$) and SUFFICIENCY-CHECK requires $2^{\Omega(n)}$ time.
The explicit ETH lower bound is still a succinct worst-case statement; Proposition 5.3 provides the finite-state query core and Propositions 5.75.16 provide theorem-level Boolean-interface refinements. [C:P5.3,5.7,5.8,5.11,5.15,5.16;R:Qf,Qb] (L: CC50,L9,L11,L8,L21,L20.)
1.3 Main Theorems
  1. 1.
    Theorem 4.6: SUFFICIENCY-CHECK is coNP-complete via reduction from TAUTOLOGY.
  2. 2.
    Theorem 4.7: MINIMUM-SUFFICIENT-SET is coNP-complete (the $\Sigma_{2}^{\textsf{P}}$ structure collapses).
  3. 3.
    Theorem 4.9: ANCHOR-SUFFICIENCY is $\Sigma_{2}^{\textsf{P}}$-complete via reduction from $\exists\forall$-SAT.
  4. 4.
    Theorem 5.1: Encoding-regime separation: explicit-state polynomial-time (polynomial in $|S|$), and under ETH a succinct worst-case lower bound witnessed by a hard family with $k^{*}=n$.
  5. 5.
    Theorem 6.1: Polynomial algorithms for bounded actions, separable utility, tree structure.
  6. 6.
    Theorem 2.37: physical-system bridge composition (law-objective semantics, behavior-preserving configuration reduction, and required-work lower bound) is proved as one bundled mechanized statement under declared encoding assumptions.
  7. 7.
    Propositions 2.282.31: the interface-time model is discrete, each decision event advances time by one unit, run-length equals elapsed interface time, and the unit-time law is substrate-invariant under the declared interface-consistency condition.
[C:T4.6,4.7,4.9,5.1,6.1,2.37;P2.28,2.29,2.30,2.31;R:E,S,S+ETH,AR] (L: CC61,CC36,CC5,CC16,CC69,CC70,CC71,CT4,DQ548,DQ537,DQ539,DQ541.)
1.4 Claim-Typing Theorems
  1. 1.
    Proposition 3.16: report admissibility is typed by certificate class ($\mathsf{ABSTAIN}$, $\mathsf{EXACT}$, $\mathsf{EPSILON}(\varepsilon)$).
  2. 2.
    Proposition 3.18: first-class evidence-object existence is equivalent to typed report admissibility.
  3. 3.
    Propositions 2.5 and 2.6: the $I=\emptyset$ query is exactly universal constancy of the decision boundary, and non-sufficiency is exactly existence of a counterexample pair.
  4. 4.
    Proposition 3.21: exact admissibility is equivalent to exact evidence existence.
  5. 5.
    Proposition 3.23 and Corollary 3.25: certified confidence is evidence-gated; for exact reports in no-competence regimes, certified confidence must be zero.
  6. 6.
    Proposition 3.24: under signal consistency, no evidence implies zero certified confidence.
  7. 7.
    Proposition 3.29 and Proposition 3.30: abstain may carry tentative guesses and self-reflected confidence, but self-confidence alone does not certify exactness.
  8. 8.
    Propositions 3.26, 3.27, 3.28: scalar step-count witnesses are always exact/falsifiable, while completion percentages are meaningful only under explicit declared bounds.
  9. 9.
    Theorem 3.56: outside declared carve-outs, admissible exact-claim policy requires explicit assumptions.
  10. 10.
    Proposition 3.61: under non-collapse, integrity and exact competence cannot both hold universally on hardness-blocked declared regimes.
  11. 11.
    Proposition 3.62 and Corollary 3.63: schema-level consequence for declared class/regime objects and exact-report inadmissibility.
[C:T3.56;P3.16,3.18,2.5,2.6,3.21,3.23,3.24,3.29,3.30,3.26,3.27,3.28,3.61,3.62;C3.25,3.63;R:DC,AR] (L: CC75,IC2,DP6,DP7,IC35,IC38,IC40,IC46,IC47,IC42,IC44,IC45,CC30,CC15,CC38.)
1.5 Regime Separation Theorems
  1. 1.
    Theorem 5.1: same-question separation between explicit-state and succinct regimes (with [S+ETH] hard-family lower bound).
  2. 2.
    Proposition 5.3: finite-state Opt-oracle lower bound in [Q_fin].
  3. 3.
    Propositions 5.75.16: Boolean-interface refinements and transfer/strictness in [Q_bool].
[C:T5.1;P5.3,5.7,5.11,5.8,5.15,5.16;R:E,S+ETH,Qf,Qb] (L: CC16,CC50,L9,L8,L11,L21,L20,L23.)
1.6 Machine-Checked Proofs
The reduction constructions and key equivalence theorems are machine-checked in Lean 4 [16, 17, 67, 21, 32] (13391 lines, 597 theorem/lemma statements). The formalization verifies that the TAUTOLOGY reduction correctly maps tautologies to sufficient coordinate sets. Complexity class membership (coNP-completeness, $\Sigma_{2}^{P}$-completeness) follows by composition with standard complexity-theoretic results.
What is new.
We contribute (i) formal definitions of decision-theoretic sufficiency in Lean; (ii) machine-checked proofs of reduction correctness for the TAUTOLOGY and $\exists\forall$-SAT reductions; (iii) a complete complexity landscape for coordinate sufficiency with explicit encoding assumptions; and (iv) a formal separation between structural complexity and representational hardness used to derive theorem-indexed engineering corollaries, including theorem-level typed coverage/completeness closures for the declared static class (Theorems 2.38, 3.4). Prior work establishes hardness informally; we formalize the constructions and their regime-typed consequences. [C:T2.38,3.4,4.6,4.9;R:CR,S] (L: CC14,CC76,CC61,CC5.)
1.7 Paper Structure
[…]
Section 2: decision-problem foundations and encoding model. Section 3: structural vs representational hardness; integrity vs competence. Section 4: hardness proofs. Section 5: regime separation. Section 6: tractable cases. Section 7: regime-conditional corollaries. Section 8: dominance theorems for hardness placement. Section 9: conservation law for decision-relevant coordinates. Section 10: related work. Appendix A: Lean listings.
Sections 79 are not applied commentary. Every claim in those sections is a formal corollary, regime-typed by the framework of Section 3 and indexed to a machine-checked theorem handle. They complete the complexity landscape by characterizing when exact minimization is rational and when it is not: the rationality matrix (Proposition 3.60), the dominance and conservation theorems (Theorems 8.16, 9.3), and the amortization threshold (Theorem 9.6) are all mechanized in Lean. [C:P3.60;T8.16,9.3,9.6;R:AR,LC] (L: IC27,HD1,HD4,HD3.)
2 Formal Foundations
We formalize decision problems with coordinate structure, sufficiency of coordinate sets, and the decision quotient, drawing on classical decision theory [55, 52]. The core definitions in this section are substrate-neutral: they specify the decision relation independently of implementation medium. Physical interpretation is introduced later through explicit encoding and assumption-transfer maps (Section 2.7).[C:D2.32;P2.34;T2.37;R:AR] (L: CT6,CT4.)
2.1 Decision Problems with Coordinate Structure
Definition 2.1 (Decision Problem).
A decision problem with coordinate structure is a tuple $\mathcal{D}=(A,X_{1},\ldots,X_{n},U)$ where:
  • AA is a finite set of actions (alternatives)
  • $X_{1},\ldots,X_{n}$ are finite coordinate spaces
  • $S=X_{1}\times\cdots\times X_{n}$ is the state space
  • $U:A\times S\to\mathbb{Q}$ is the utility function
Definition 2.2 (Projection).
For state $s=(s_{1},\ldots,s_{n})\in S$ and coordinate set $I\subseteq\{1,\ldots,n\}$:
$s_{I}:=(s_{i})_{i\in I}$
is the projection of $s$ onto coordinates in $I$.
Definition 2.3 (Optimizer Map).
For state $s\in S$, the optimal action set is:
$\text{Opt}(s):=\arg\max_{a\in A}U(a,s)=\{a\in A:U(a,s)=\max_{a^{\prime}\in A}U(a^{\prime},s)\}$
2.2 Sufficiency and Relevance
Definition 2.4 (Sufficient Coordinate Set).
A coordinate set $I\subseteq\{1,\ldots,n\}$ is sufficient for decision problem $\mathcal{D}$ if:
$\forall s,s^{\prime}\in S:\quad s_{I}=s^{\prime}_{I}\implies\text{Opt}(s)=\text{Opt}(s^{\prime})$
Equivalently, the optimal action depends only on coordinates in $I$. In this paper, this set-valued invariance of the full optimal-action correspondence is the primary decision-relevance target.
Proposition 2.5 (Empty-Set Sufficiency Equals Constant Decision Boundary).
For any decision problem,
$\text{sufficient}(\emptyset)\iff\forall s,s^{\prime}\in S,\ \text{Opt}(s)=\text{Opt}(s^{\prime}).$
Hence, the $I=\emptyset$ query is exactly a universal constancy check of the decision boundary. (L: DP6.)
Proposition 2.6 (Insufficiency Equals Counterexample Witness).
For any coordinate set $I$,
$\neg\text{sufficient}(I)\iff\exists s,s^{\prime}\in S:\ s_{I}=s^{\prime}_{I}\ \wedge\ \text{Opt}(s)\neq\text{Opt}(s^{\prime}).$
In particular,
$\neg\text{sufficient}(\emptyset)\iff\exists s,s^{\prime}\in S:\ \text{Opt}(s)\neq\text{Opt}(s^{\prime}).$
(L: DP7, DP8.)
[C:D2.4;P2.5,2.6;R:DM]
Definition 2.7 (Deterministic Selector).
A deterministic selector is a map $\sigma:2^{A}\to A$ that returns one action from an optimal-action set.
Definition 2.8 (Selector-Level Sufficiency).
For fixed selector $\sigma$, a coordinate set $I$ is selector-sufficient if
$\forall s,s^{\prime}\in S:\quad s_{I}=s^{\prime}_{I}\implies\sigma(\text{Opt}(s))=\sigma(\text{Opt}(s^{\prime})).$
Proposition 2.9 (Set-Level Target Dominates Selector Targets).
If $I$ is sufficient in the set-valued sense (Definition 2.4), then $I$ is selector-sufficient for every deterministic selector $\sigma$ (Definition 2.8). (L: DP5)
Proof.
If $I$ is sufficient, then $s_{I}=s^{\prime}_{I}$ implies $\text{Opt}(s)=\text{Opt}(s^{\prime})$. Applying any deterministic selector $\sigma$ to equal optimal-action sets yields equal selected actions.  $\blacksquare$
Definition 2.10 ($\varepsilon$-Optimal Set and $\varepsilon$-Sufficiency).
For $\varepsilon\geq 0$, define
$\text{Opt}_{\varepsilon}(s):=\{a\in A:\forall a^{\prime}\in A,\ U(a^{\prime},s)\leq U(a,s)+\varepsilon\}.$
Coordinate set $I$ is $\varepsilon$-sufficient if
$\forall s,s^{\prime}\in S:\ s_{I}=s^{\prime}_{I}\implies\text{Opt}_{\varepsilon}(s)=\text{Opt}_{\varepsilon}(s^{\prime}).$
Proposition 2.11 (Zero-$\varepsilon$ Reduction).
Set-valued sufficiency is exactly the $\varepsilon=0$ case:
[…]
(L: DP1, DP4)
Proposition 2.12 (Selector-Level Separation Witness).
The converse of Proposition 2.9 fails in general: there exists a decision problem, selector, and coordinate set $I$ such that $I$ is selector-sufficient but not sufficient in the full set-valued sense. (L: CC54.)
Definition 2.13 (Minimal Sufficient Set).
A sufficient set $I$ is minimal if no proper subset $I^{\prime}\subsetneq I$ is sufficient.
Definition 2.14 (Relevant Coordinate).
Coordinate $i$ is relevant if there exist states that differ only at coordinate $i$ and induce different optimal action sets:
$i\text{ is relevant}\iff\exists s,s^{\prime}\in S:\;\Big(\forall j\neq i,\;s_{j}=s^{\prime}_{j}\Big)\ \wedge\ \text{Opt}(s)\neq\text{Opt}(s^{\prime}).$
Proposition 2.15 (Minimal-Set/Relevance Equivalence).
For any minimal sufficient set $I$ and any coordinate $i$:
$i\in I\iff i\text{ is relevant}.$
Hence every minimal sufficient set is exactly the relevant-coordinate set. (L: DP2, DP3.)
Proof.
The “only if” direction follows by minimality: if $i\in I$ were irrelevant, removing $i$ would preserve sufficiency, contradicting minimality. The “if” direction follows from sufficiency: every sufficient set must contain each relevant coordinate.  $\blacksquare$
Definition 2.16 (Exact Relevance Identifiability).
For a decision problem $\mathcal{D}$ and candidate coordinate set $I$, we say $I$ is exactly relevance-identifying if
$\forall i\in\{1,\ldots,n\}:\quad i\in I\iff i\text{ is relevant for }\mathcal{D}.$
Equivalently, $I$ is exactly relevance-identifying iff $I$ equals the full relevant-coordinate set.
Example 2.17 (Weather Decision).
Consider deciding whether to carry an umbrella:
  • Actions: $A=\{\text{carry},\text{don't carry}\}$
  • Coordinates: $X_{1}=\{\text{rain},\text{no rain}\}$, $X_{2}=\{\text{hot},\text{cold}\}$, $X_{3}=\{\text{Monday},\ldots,\text{Sunday}\}$
[…]
The minimal sufficient set is $I=\{1\}$ (only rain forecast matters). Coordinates 2 and 3 (temperature, day of week) are irrelevant.
2.3 The Decision Quotient
Definition 2.18 (Decision Equivalence).
For coordinate set $I$, states $s,s^{\prime}$ are $I$-equivalent (written $s\sim_{I}s^{\prime}$) if $s_{I}=s^{\prime}_{I}$.
Definition 2.19 (Decision Quotient).
The decision quotient for state $s$ under coordinate set $I$ is:
$\text{DQ}_{I}(s)=\frac{|\{a\in A:a\in\text{Opt}(s^{\prime})\text{ for some }s^{\prime}\sim_{I}s\}|}{|A|}$
This measures the fraction of actions that are optimal for at least one state consistent with $I$.
Proposition 2.20 (Sufficiency Characterization).
[C:P2.20;R:H=succinct-hard] Coordinate set $I$ is sufficient if and only if $\text{DQ}_{I}(s)=|\text{Opt}(s)|/|A|$ for all $s\in S$. (L: CC62, CC63)
Proof.
If $I$ is sufficient, then $s\sim_{I}s^{\prime}\implies\text{Opt}(s)=\text{Opt}(s^{\prime})$, so the set of actions optimal for some $s^{\prime}\sim_{I}s$ is exactly $\text{Opt}(s)$.
Conversely, if the condition holds, then for any $s\sim_{I}s^{\prime}$, the optimal actions form the same set (since $\text{DQ}_{I}(s)=\text{DQ}_{I}(s^{\prime})$ and both equal the relative size of the common optimal set).  $\blacksquare$
2.4 Computational Model and Input Encoding
We fix the computational model used by the complexity claims. In physical deployments, these encodings correspond to different access conditions on the same underlying system: explicit tables correspond to fully exposed state/utility structure, succinct encodings correspond to compressed generative descriptions, and query regimes correspond to constrained measurement interfaces.[C:T5.1;P5.3,2.34;R:E,S+ETH,Qf,AR] (L: CC16,CC50,CT6.)
Succinct encoding (primary for hardness).
This succinct circuit encoding is the standard representation for decision problems in complexity theory; hardness is stated with respect to the input length of the circuit description [5, 47]. An instance is encoded as:
  • a finite action set $A$ given explicitly,
  • coordinate domains $X_{1},\ldots,X_{n}$ given by their sizes in binary,
  • a Boolean or arithmetic circuit $C_{U}$ that on input $(a,s)$ outputs $U(a,s)$.
The input length is $L=|A|+\sum_{i}\log|X_{i}|+|C_{U}|$. Polynomial time and all complexity classes (coNP, $\Sigma_{2}^{P}$, ETH) are measured in $L$. All hardness results in Section 4 use this encoding.
Explicit-state encoding (used for enumeration algorithms and experiments).
The utility is given as a full table over $A\times S$. The input length is $L_{\text{exp}}=\Theta(|A||S|)$ (up to the bitlength of utilities). Polynomial time is measured in $L_{\text{exp}}$. Results stated in terms of $|S|$ use this encoding.
Query-access regime (intermediate black-box access).
The solver is given oracle access to decision information at queried states (e.g., $\text{Opt}(s)$, or $U(a,s)$ with $\text{Opt}(s)$ reconstructed from finitely many value queries). Complexity is measured by oracle-query count (optionally paired with per-query evaluation cost). This separates representational access from full-table availability and from succinct-circuit input length.
Unless explicitly stated otherwise, “polynomial time” refers to the succinct encoding.
Remark 2.21 (Decision Questions vs Decision Problems).
We distinguish between a decision question and a decision problem. A decision question is the encoding-independent predicate: “is coordinate set $I$ sufficient for $\mathcal{D}$?” This is a mathematical relation, fixed by the model contract (C1–C4). A decision problem is a decision question together with a specific encoding of its inputs. The same decision question yields different decision problems under different encodings, and these problems may belong to different complexity classes. Formally, each encoding regime $e$ induces a language $L_{\mathcal{D},e}$; complexity classification is a property of $L_{\mathcal{D},e}$, not of $\mathcal{D}$ alone.
Proposition 2.22 (Under-Resolution Forces Representation Collision).
[C:P2.22;R:DM] Let $\mathrm{repr}:\mathcal{L}\to\mathcal{C}$ be any representation map from logical possibilities to code states. If $|\mathcal{C}|<|\mathcal{L}|$, then there exist distinct logical states with the same representation:
$\exists s,s^{\prime}\in\mathcal{L},\ s\neq s^{\prime}\wedge\mathrm{repr}(s)=\mathrm{repr}(s^{\prime}).$
This is the finite-cardinality collision statement underlying under-resolved exact-decision claims. (L: DQ582, DQ583.)
Definition 2.23 (Regime Simulation).
Fix a decision question $Q$ and two regimes $R_{1},R_{2}$ for that question. We say $R_{1}$ simulates $R_{2}$ if there exists a polynomial-time transformer that maps any $R_{2}$ instance to an $R_{1}$ instance while preserving the answer to $Q$; in oracle settings, this includes a polynomial-time transcript transducer that preserves yes/no outcomes for induced solvers. (L: CC1, CC52.)
This paper uses that simulation spine in two explicit forms: restriction-map simulation (subproblem-to-full transfer; Proposition 5.8) and oracle-transducer simulation (batch-to-entry transfer; Proposition 5.15).
2.5 Model Contract and Regime Tags
All theorem statements in this paper are typed by the following model contract:
  • C1 (finite actions): $A$ is finite.
  • C2 (finite coordinate domains): each $X_{i}$ is finite, so $S=\prod_{i}X_{i}$ is finite.
  • C3 (evaluable utility): $U(a,s)$ is computable from the declared input encoding.
  • C4 (fixed decision semantics): optimality is defined by $\text{Opt}(s)=\arg\max_{a}U(a,s)$.
We use short regime tags for applied corollaries:
  • [E] explicit-state encoding,
  • [Q] query-access (oracle) regime,
  • [Q_fin] finite-state query-access core (state-oracle),
  • [S] succinct encoding,
  • [S+ETH] succinct encoding with ETH,
  • [Q_bool] mechanized Boolean query submodel,
  • [S_bool] mechanized Boolean-coordinate submodel.
This tagging is a claim-typing convention: each strong statement is attached to the regime where it is proven.
Remark 2.24 (Uniform Assumption Discipline).
All theorem-level claims in this paper are conditional on their stated premises and regime tags; no assumption family is given special status. For each claim, the active assumptions are exactly those appearing in its statement, its regime tag, and its cited mechanized handles.
Remark 2.25 (Modal Convention: Contract-Conditional Necessity).
In this paper, normative modal statements denote contract-conditional necessity: once an agent declares an objective/regime/assumption contract, admissible policy is fixed by typed admissibility and integrity under that contract. No unconditional external moral prescription is asserted.
2.6 Discrete-Time Event Semantics
Definition 2.26 (Timed Interface State).
For any interface state space $S$, a timed interface state is a pair
$x=(s,t)\in S\times\mathbb{N}.$
The time coordinate is the natural-number component $t$.
Definition 2.27 (Decision Tick and Decision Event).
Fix an interface decision process $(\mathrm{decide},\mathrm{transition})$ with
$\mathrm{decide}:S\to A,\qquad\mathrm{transition}:S\times A\to S.$
Define one tick by
$\mathrm{tick}(s,t)=\bigl(\mathrm{transition}(s,\mathrm{decide}(s)),\,t+1\bigr).$
Define $\mathrm{DecisionEvent}(x,y)$ to mean that $y$ is the one-tick successor of $x$ under this rule.
Proposition 2.28 (Time Coordinate Is Discrete).
In Definition 2.26, time is discrete:
$\forall x\in S\times\mathbb{N},\ \exists k\in\mathbb{N}\ \text{such that}\ x_{t}=k.$
Pointwise time equality is decidable for every $k\in\mathbb{N}$:
$x_{t}=k\ \vee\ x_{t}\neq k.$
(L: DQ548,DQ547.)
Proposition 2.29 (Decision Event Is One Unit of Time).
Under Definition 2.27, for all timed states $x,y$:
$\mathrm{DecisionEvent}(x,y)\ \Longleftrightarrow\ y=\mathrm{tick}(x),$
and therefore
$\mathrm{DecisionEvent}(x,y)\ \Longrightarrow\ y_{t}=x_{t}+1.$
(L: DQ535,DQ537,DQ545,DQ546.)
Proposition 2.30 (Run Accounting Identity).
Let $\mathrm{run}(n,x)$ be $n$ repeated ticks from $x$, and let $\mathrm{trace}(n,x)$ be the emitted decision sequence. Then:
$\mathrm{run}(n,x)_{t}=x_{t}+n,\qquad\mathrm{run}(n,x)_{t}-x_{t}=n,\qquad|\mathrm{trace}(n,x)|=n.$
Hence decision count equals elapsed time:
$|\mathrm{trace}(n,x)|=\mathrm{run}(n,x)_{t}-x_{t}.$
(L: DQ540,DQ539,DQ533,DQ534.)
Proposition 2.31 (Substrate Realization Preserves Unit-Time Events).
For any substrate model that is consistent with the declared interface transition rule, each one-step substrate evolution realizes one interface decision event and advances interface time by one unit. The time update law is invariant under substrate kind. (L: DQ542,DQ541,DQ549.)
2.7 Physical-System Encoding and Theorem Transport
Definition 2.32 (Physical-to-Core Encoding).
Fix a physical instance class $\mathcal{P}$ and a core decision class $\mathcal{D}$. A physical-to-core encoding is a map
$E:\mathcal{P}\to\mathcal{D}.$
All physical claims below are typed relative to a declared encoding map $E$ and declared assumption transfer map from physical assumptions to core assumptions.
Proposition 2.33 (Heisenberg-Strong Binding Yields Nontrivial Decision Boundary).
[C:P2.33;R:AR] Fix a declared noisy physical encoding interface and assume a Heisenberg-strong binding witness: one physical instance is compatible with two distinct interface states whose encoded optimal-action sets differ. Then there exists a core decision problem with non-constant optimizer map. Equivalently, under this declared physical ambiguity assumption, decision nontriviality is derivable at the core level. (L: DQ584, DQ585, DQ586. )
Proposition 2.34 (Conditional Physical Claim Transport).
Let $E:\mathcal{P}\to\mathcal{D}$ be an encoding. If
$\forall d\in\mathcal{D},\ \mathrm{CoreAssm}(d)\Rightarrow\mathrm{CoreClaim}(d)$
and
$\forall p\in\mathcal{P},\ \mathrm{PhysAssm}(p)\Rightarrow\mathrm{CoreAssm}(E(p)),$
then
$\forall p\in\mathcal{P},\ \mathrm{PhysAssm}(p)\Rightarrow\mathrm{CoreClaim}(E(p)).$
(L: CT6, CT5, DQ572.)
Corollary 2.35 (Physical Counterexample Implies Core Failure on Encoded Slice).
Under the same encoding and assumption-transfer map:
  1. 1.
    any encoded physical counterexample induces a core counterexample on the encoded slice;
  2. 2.
    if a physical counterexample exists, then the purported core rule is invalid;
  3. 3.
    if the lifted claim holds on all physically admissible encoded instances, no such physical counterexample exists.
(L: CT8, CT7, CT3, DQ571.)
Proposition 2.36 (Law-Instance Objective Bridge).
For the mechanized law-gap physical instance class, the encoded optimizer claim holds at theorem level and admits no counterexample:
$\forall x\in\mathcal{P}_{\mathrm{law}},\ \mathrm{lawGapPhysicalClaim}(x),\qquad\neg\exists x\in\mathcal{P}_{\mathrm{law}},\ \neg\,\mathrm{lawGapPhysicalClaim}(x).$
(L: CT1, CT2.)
Theorem 2.37 (Physical Bridge Bundle).
In one bundled mechanized statement, the paper’s physical bridge layer composes:
  1. 1.
    law-induced objective semantics ($\text{Opt}=$ feasible actions for law-gap instances),
  2. 2.
    behavior-preserving configuration reduction equivalence,
  3. 3.
    irreducible required-work lower bound under site-count scaling.
(L: CT4.)
[C:D2.32;P2.34,2.36;C2.35;T2.37;R:AR]
Theorem 2.38 (Declared Regime Coverage for the Static Class).
Let
$\mathcal{R}_{\mathrm{static}}=\{\text{[E]},\ \text{[S+ETH]},\ \text{[Q\_fin:Opt]},\ \text{[Q\_bool:value-entry]},\ \text{[Q\_bool:state-batch]}\}.$
For each declared regime in $\mathcal{R}_{\mathrm{static}}$, the paper has a theorem-level mechanized core claim. Equivalently, regime typing is complete over the declared static family. (L: CC14, CC51.)
Scope Lattice (typed classes and transfer boundary).
Layer Transfer Status Lean handles (representative)
Static sufficiency class (C1–C4; declared regimes) Internal landscape complete CC76
CC51
Bridge-admissible adjacent class (one-step deterministic) Transfer licensed CC42
CC10
Non-admissible represented adjacent classes (horizon, stochastic, transition-coupled) Transfer blocked by witness CC9
CC8
2.8 Adjacent Objective Regimes and Bridge
Definition 2.39 (Adjacent Sequential Objective Regime).
An adjacent sequential objective instance consists of:
  • finite action set $A$,
  • finite coordinate state space $S=X_{1}\times\cdots\times X_{n}$,
  • horizon $T\in\mathbb{N}_{\geq 1}$ and history-dependent policy class,
  • reward process $r_{t}$ and objective functional $J(\pi)$ (e.g., cumulative reward or regret).
Proposition 2.40 (One-Step Deterministic Bridge).
Consider an instance of Definition 2.39 satisfying:
  1. 1.
    $T=1$,
  2. 2.
    deterministic rewards $r_{1}(a,s)=U(a,s)$ for some evaluable $U$,
  3. 3.
    objective $J(\pi)=U(\pi(s),s)$ (single-step maximization),
  4. 4.
    no post-decision state update relevant to the objective.
Then the induced optimization problem is exactly the static decision problem of Definition 2.1, and coordinate sufficiency in the sequential formulation is equivalent to Definition 2.4. (L: CC42)
Proof.
Under (1)–(3), optimizing $J$ at state $s$ is identical to choosing an action in $\arg\max_{a\in A}U(a,s)=\text{Opt}(s)$. Condition (4) removes any dependence on future transition effects. Therefore the optimal-policy relation in the adjacent formulation coincides pointwise with Opt from Definition 2.3, and the invariance condition “same projection implies same optimal choice set” is exactly Definition 2.4.  $\blacksquare$
Proposition 2.41 (Bridge Transfer Rule).
Under conditions (1)–(4), any sufficiency statement formulated over Definition 2.4 is equivalent between the adjacent sequential formulation and the static model. (L: CC42)
Proof.
By Proposition 2.40, the adjacent sequential objective induces exactly the same sufficiency predicate as Definition 2.4 when (1)–(4) hold. Equivalence of any sufficiency statement then follows by substitution.  $\blacksquare$
If any bridge condition fails, direct transfer from this paper’s static complexity theorems is not licensed by this rule.
Remark 2.42 (Extension Boundary).
Beyond Proposition 2.40 (multi-step horizon, stochastic transitions/rewards, or regret objectives), the governing complexity objects change. Those regimes are natural extensions, but they are distinct formal classes from the static sufficiency class analyzed in this paper.
Proposition 2.43 (Horizon->1>1 Bridge-Failure Witness).
Dropping the one-step condition can break sufficiency transfer: there exists a two-step objective where sufficiency in the immediate-utility static projection does not match sufficiency in the two-step objective. (L: CC26, CC25.)
Proposition 2.44 (Stochastic-Criterion Bridge-Failure Witness).
If optimization is performed against a stochastic criterion not equal to deterministic utility maximization, bridge transfer to Definition 2.1 can fail even when an expected-utility projection is available. (L: CC57.)
Proposition 2.45 (Transition-Coupled Bridge-Failure Witness).
If post-decision transition effects are objective-relevant, bridge transfer to the static one-step class can fail. (L: CC73.)
Theorem 2.46 (Exact Bridge Boundary in the Represented Adjacent Family).
Within the represented adjacent family
$\{\text{one-step deterministic},\ \text{horizon-extended},\ \text{stochastic-criterion},\ \text{transition-coupled}\},$
direct transfer of static sufficiency claims is licensed if and only if the class is one-step deterministic. Each represented non-one-step class has an explicit mechanized bridge-failure witness. (L: CC10, CC9, CC8.)
Definition 2.47 (Agent Snapshot vs Agent Process Typing).
Within the represented adjacent family, we type:
  • agent snapshot: fixed-parameter inference view (mapped to one-step deterministic class),
  • agent process: online/dynamical update views (horizon-extended, stochastic-criterion, or transition-coupled classes).
Proposition 2.48 (Snapshot/Process Transfer Typing).
In the represented adjacent family, direct transfer of static sufficiency claims is licensed if and only if the agent is typed as an agent snapshot. Every process-typed represented class has an explicit mechanized bridge-failure witness. (L: CC3, CC48, CC56.)
Operationally: a trained model evaluated with fixed parameters during inference is an agent snapshot for this typing; once parameters or transition-relevant internals are updated online, the object is process-typed and transfer from static sufficiency theorems is blocked unless the one-step bridge conditions are re-established.[C:P2.48;R:RA]
3 Interpretive Foundations: Hardness and Solver Claims
The claims in later applied sections are theorem-indexed consequences of this section and Sections 46.
3.1 Structural Complexity vs Representational Hardness
Definition 3.1 (Structural Complexity).
For a fixed decision question (e.g., “is $I$ sufficient for $\mathcal{D}$?”; see Remark 2.21), structural complexity means its placement in standard complexity classes within the formal model (coNP, $\Sigma_{2}^{P}$, etc.), as established by class-membership arguments and reductions.
Definition 3.2 (Representational Hardness).
For a fixed decision question and an encoding regime $E$ (Section 2.4), representational hardness is the worst-case computational cost incurred by solvers whose input access is restricted to $E$.
In the mechanized architecture model, this is made explicit as a required-work observable and intrinsic lower bound: ‘requiredWork‘ is total realized work, ‘hardnessLowerBound‘ is the intrinsic floor, and ‘hardness_is_irreducible_required_work‘ proves the lower-bound relation for all nonzero use-site counts; ‘requiredWork_eq_affine_in_sites‘ and ‘workGrowthDegree‘ make the site-count growth law explicit.(L: DQ562,DQ560,DQ561,DQ563,DQ564,DQ565.)
Remark 3.3 (Interpretation Contract).
This paper keeps the decision question fixed and varies the encoding regime explicitly. Thus, later separations are read as changes in representational hardness under fixed structural complexity, not as changes to the underlying sufficiency semantics. Accordingly, “complete landscape” means complete for this static sufficiency class under the declared regimes; adjacent objective classes are distinct typed objects, not unproved remainder cases of the same class.
Theorem 3.4 (Typed Completeness for the Static Sufficiency Class).
For the static sufficiency class fixed by C1–C4 and declared regime family $\mathcal{R}_{\mathrm{static}}$, the paper provides:
  1. 1.
    conditional class-label closures for SUFFICIENCY-CHECK, MINIMUM-SUFFICIENT-SET, and ANCHOR-SUFFICIENCY;
  2. 2.
    a conditional explicit/succinct dichotomy closure;
  3. 3.
    a regime-indexed mechanized core claim for every declared regime; and
  4. 4.
    declared-family exhaustiveness in the typed regime algebra.
(L: CC76.)
3.2 Solver Integrity and Regime Competence
To keep practical corollaries type-safe, we separate integrity (what a solver is allowed to assert) from competence (what it can cover under a declared regime), following the certifying-algorithms schema [39, 43]. The induced-relation view used below is standard in complexity/computability presentations: algorithms are analyzed as machines deciding languages or computing functions/relations over encodings [47, 5, 21].
Definition 3.5 (Certifying Solver).
Fix a decision question $\mathcal{R}\subseteq\mathcal{X}\times\mathcal{Y}$ and an encoding regime $E$ over $\mathcal{X}$. A certifying solver is a pair $(Q,V)$ where:
  • $Q$ maps each input $x\in\mathcal{X}$ to either $\mathsf{ABSTAIN}$ or a candidate pair $(y,w)$,
  • $V$ is a polynomial-time checker (in $|{\rm enc}_{E}(x)|$) with output in $\{0,1\}$.
Definition 3.6 (Induced Solver Relation).
For any deterministic program $M$ computing a partial map $f_{M}:\mathcal{X}\rightharpoonup\mathcal{Y}$ on encoded inputs, define
$\mathcal{R}_{M}:=\{(x,y)\in\mathcal{X}\times\mathcal{Y}:f_{M}(x)\downarrow\ \wedge\ y=f_{M}(x)\}.$
Proposition 3.7 (Universality of Solver Framing).
Every deterministic computer program that computes a (possibly partial) map on encoded inputs can be framed as a solver for its induced relation (Definition 3.6). In this formal sense, all computers are solvers of specific problems. (L: CC77.)
Proof.
Immediate from Definition 3.6: each program induces a relation given by its graph on the domain where it halts.  $\blacksquare$
Definition 3.8 (Solver Integrity).
A certifying solver $(Q,V)$ has integrity for relation $\mathcal{R}$ if:
  • (assertion soundness) $Q(x)=(y,w)\implies V(x,y,w)=1$,
  • (checker soundness) $V(x,y,w)=1\implies(x,y)\in\mathcal{R}$.
The output $\mathsf{ABSTAIN}$ (equivalently, $\mathsf{UNKNOWN}$) is first-class and carries no assertion about membership in $\mathcal{R}$.
Definition 3.9 (Competence Under a Regime).
Fix a regime $\Gamma=(\mathcal{X}_{\Gamma},E_{\Gamma},\mathcal{C}_{\Gamma})$ with instance family $\mathcal{X}_{\Gamma}\subseteq\mathcal{X}$, encoding assumptions $E_{\Gamma}$, and resource bound $\mathcal{C}_{\Gamma}$. A certifying solver $(Q,V)$ is competent on $\Gamma$ for relation $\mathcal{R}$ if:
  • it has integrity for $\mathcal{R}$ (Definition 3.8),
  • (coverage) $\forall x\in\mathcal{X}_{\Gamma},\;Q(x)\neq\mathsf{ABSTAIN}$,
[…]
Corollary 3.10 (Universality of Integrity Applicability).
Let $M$ be any deterministic program, viewed as a certifying solver pair $(Q,V)$ for some externally fixed target relation $\mathcal{R}$. Then Definition 3.8 applies unchanged. Thus epistemic integrity is architecture-universal over computing systems once they are cast as solvers for declared tasks. (L: CC31.)
Proof.
Definition 3.8 quantifies over a relation $\mathcal{R}$ and a certifying pair $(Q,V)$; it does not assume any special implementation substrate.  $\blacksquare$
[C:P3.7;C3.10;R:TR]
Remark 3.11 (External-Task Non-Vacuity).
If the target relation is chosen post hoc from the program’s own behavior (for example, $\mathcal{R}=\mathcal{R}_{M}$ from Definition 3.6), integrity can become tautological and competence claims can be vacuous. Non-vacuous competence claims therefore require the target relation, encoding regime, and resource bound to be declared independently of observed outputs. [C:D3.8,3.9;P3.58,3.61;R:AR]
Definition 3.12 ($\varepsilon$-Objective Relation Family).
An $\varepsilon$-objective relation family is a map
$\varepsilon\mapsto\mathcal{R}_{\varepsilon}\subseteq\mathcal{X}\times\mathcal{Y}$
that assigns a certified target relation to each tolerance level $\varepsilon\geq 0$. The exact target is $\mathcal{R}_{0}$.
Definition 3.13 ($\varepsilon$-Competence Under a Regime).
Fix a relation family $(\mathcal{R}_{\varepsilon})_{\varepsilon\geq 0}$ and regime $\Gamma$. A certifying solver $(Q,V)$ is $\varepsilon$-competent on $\Gamma$ if it is competent on $\Gamma$ for relation $\mathcal{R}_{\varepsilon}$ in the sense of Definition 3.9.
Proposition 3.14 (Zero-$\varepsilon$ Competence Reduction).
If $\mathcal{R}_{0}=\mathcal{R}$, then
[…]
Moreover, $\varepsilon$-competence implies integrity for the corresponding $\mathcal{R}_{\varepsilon}$. (L: IC33, IC19.)
Definition 3.15 (Typed Claim Report).
For a declared objective family and regime, a solver-side report is one of:
  • $\mathsf{ABSTAIN}$,
  • $\mathsf{EXACT}$ (claiming exact competence/certification),
  • $\mathsf{EPSILON}(\varepsilon)$ (claiming $\varepsilon$-competence/certification).
Proposition 3.16 (Typed Claim Admissibility Discipline).
Under Definitions 3.9, 3.13, and 3.15:
  • $\mathsf{ABSTAIN}$ is always admissible;
  • $\mathsf{EXACT}$ is admissible iff exact competence holds;
  • $\mathsf{EPSILON}(\varepsilon)$ is admissible iff $\varepsilon$-competence holds.
Hence claim type and certificate type must match (no cross-typing of uncertified assertions as certified guarantees). (L: CC75.)
Definition 3.17 (Evidence Object for a Typed Report).
For declared $(\mathcal{R},(\mathcal{R}_{\varepsilon})_{\varepsilon\geq 0},\Gamma,Q)$ and report type $r\in\{\mathsf{ABSTAIN},\mathsf{EXACT},\mathsf{EPSILON}(\varepsilon)\}$, evidence is a first-class witness object:
  • for $\mathsf{ABSTAIN}$: trivial witness;
  • for $\mathsf{EXACT}$: exact-competence witness;
  • for $\mathsf{EPSILON}(\varepsilon)$: $\varepsilon$-competence witness.
(L: IC2.)
Proposition 3.18 (Evidence–Admissibility Equivalence).
For any declared contract and report type $r$:
$\text{Evidence object for }r\ \text{exists}\iff\text{Claim }r\ \text{is admissible}.$
(L: IC17, IC21, IC20.)
[C:D3.17;P3.16,3.18;R:AR]
Security-game reading (derived).
The typed-report layer can be read as a standard claim-verification game under a declared contract: a reporting algorithm emits a report token $r$ and optional evidence, and a checker accepts if and only if the evidence is valid for $r$. Within this game view, existing results already provide the core security properties: completeness for admissible report classes (Proposition 3.18), soundness against overclaiming (Propositions 3.16, 3.21), and evidence-gated confidence (Propositions 3.23, 3.24). So integrity is enforced as verifiable claim discipline, not as unchecked payload declaration. [C:P3.16,3.18,3.21,3.23,3.24;R:AR]
Definition 3.19 (Signaled Typed Report (Guess + Confidence)).
A signaled typed report augments the typed report token $r\in\{\mathsf{ABSTAIN},\mathsf{EXACT},\mathsf{EPSILON}(\varepsilon)\}$ with:
  • an optional guess payload $g$,
  • self-reflected confidence $p_{\mathrm{self}}\in[0,1]$,
  • certified confidence $p_{\mathrm{cert}}\in[0,1]$.
  • scalar execution witness $t_{\mathrm{run}}\in\mathbb{N}$ (steps actually run),
  • optional declared bound $B\in\mathbb{N}$.
Here $p_{\mathrm{self}}$ is a self-signal, while $p_{\mathrm{cert}}$ is an evidence-gated signal under the declared contract. (L: IC34.)
Definition 3.20 (Signal Consistency Discipline).
For a signaled typed report $(r,g,p_{\mathrm{self}},p_{\mathrm{cert}},t_{\mathrm{run}},B)$, require:
$p_{\mathrm{cert}}>0\Rightarrow\exists\ \text{evidence object for }r.$
Proposition 3.21 (Exact Claims Require Exact Evidence).
In the typed claim discipline,
$\text{ClaimAdmissible}(\mathsf{EXACT})\iff\exists\ \text{exact evidence object}.$
Equivalently, admissible exact claims are exactly those with an evidence witness. (L: IC35, IC36.)
Definition 3.22 (Bound-Dependent Completion Fraction).
Completion-fraction semantics are defined only when a positive declared bound exists and covers observed runtime:
$\mathrm{CompletionFractionDefined}\iff\exists b>0:\ B=b\ \wedge\ t_{\mathrm{run}}\leq b.$
(L: IC37.)
Proposition 3.23 (Certified-Confidence Gate).
Under Definition 3.20, positive certified confidence implies typed admissibility:
$p_{\mathrm{cert}}>0\Rightarrow\text{ClaimAdmissible}(r).$
Thus certified confidence is not a free scalar; it is evidence-gated by the same typed discipline as the report itself. (L: IC38, IC39.)
Proposition 3.24 (No Evidence Forces Zero Certified Confidence).
Under Definition 3.20, absence of evidence for the reported type forces:
$\neg\exists\ \text{evidence object for }r\Rightarrow p_{\mathrm{cert}}=0.$
(L: IC40.)
Corollary 3.25 (Exact-No-Competence Zero-Certified Constraint).
For exact reports, if exact competence is unavailable on the declared regime/objective, then any signal-consistent report must satisfy:
$p_{\mathrm{cert}}=0.$
(L: IC41.)
Proposition 3.26 (Scalar Step Witness Is Always Defined and Falsifiable).
For every signaled report, the execution witness is an exact natural number and any equality claim about it is decidable:
$\exists k\in\mathbb{N}:\ t_{\mathrm{run}}=k,\qquad\forall k\in\mathbb{N},\ (t_{\mathrm{run}}=k)\ \vee\ (t_{\mathrm{run}}\neq k).$
(L: IC42, IC43.)
Proposition 3.27 (No Completion Fraction Without Declared Bound).
If no declared bound is provided ($B=\varnothing$), completion-fraction semantics are undefined:
$B=\varnothing\Rightarrow\neg\,\mathrm{CompletionFractionDefined}.$
(L: IC44.)
Proposition 3.28 (Completion Fraction Under Explicit Bound).
If a positive declared bound is provided and observed runtime is within bound, completion-fraction semantics are defined:
$B=b,\ b>0,\ t_{\mathrm{run}}\leq b\Rightarrow\mathrm{CompletionFractionDefined}.$
(L: IC45.)
Proposition 3.29 (Abstain Can Carry Guess and Self-Confidence).
For any optional guess payload $g$ and any self-reflected confidence $p_{\mathrm{self}}\in[0,1]$, there exists a signal-consistent abstain report
$(\mathsf{ABSTAIN},g,p_{\mathrm{self}},0,0,\varnothing).$
Hence the framework is non-binary at the report-signal layer: abstention can carry a tentative answer and self-reflection without upgrading to certified exactness. (L: IC46.)
Proposition 3.30 (Self-Reflected Confidence Is Not Certification).
Self-reflected confidence alone does not certify exactness: if exact competence is unavailable, there exist exact-labeled reports with arbitrary $p_{\mathrm{self}}$ that remain inadmissible under zero certified confidence. (L: IC47.)
Definition 3.31 (Declared Budget Slice).
For a declared regime $\Gamma$ over state objects and horizon/budget cut $H\in\mathbb{N}$, define the in-scope slice
$\mathcal{S}_{\Gamma,H}:=\{s:\ s\in\Gamma.\mathrm{inScope}\ \wedge\ \mathrm{encLen}_{\Gamma}(s)\leq H\}.$
(L: CC13.)
Proposition 3.32 (Bounded-Slice Irrelevance of a Meta-Coordinate).
Fix coordinate $i_{\infty}$ and declared slice $\mathcal{S}_{\Gamma,H}$ from Definition 3.31. If optimizer sets are invariant to $i_{\infty}$ on that slice, i.e.,
$\forall s,s^{\prime}\in\mathcal{S}_{\Gamma,H},\ \Big(\forall j\neq i_{\infty},\ s_{j}=s^{\prime}_{j}\Big)\Rightarrow\text{Opt}(s)=\text{Opt}(s^{\prime}),$
then $i_{\infty}$ is irrelevant on $\mathcal{S}_{\Gamma,H}$, and hence not relevant for the declared task slice. (L: CC32, CC33.)
[C:D3.31;P3.32;R:AR]
Definition 3.33 (Goal Class).
A goal class is a non-empty set of admissible evaluators over a fixed action/state structure:
$\mathcal{G}=(\mathcal{U}_{\mathcal{G}},A,S),\qquad\varnothing\neq\mathcal{U}_{\mathcal{G}}\subseteq\{U:A\times S\to\mathbb{Q}\}.$
The solver need not know which $U\in\mathcal{U}_{\mathcal{G}}$ is active; it only knows class-membership constraints. (L: IV1.)
Definition 3.34 (Interior Pareto Dominance).
For a goal class $\mathcal{G}$ and score model $\sigma:S\times\{1,\ldots,n\}\to\mathbb{Q}$, let $J_{\mathcal{G}}$ be the class-tautological coordinate set. State $s$ interior-Pareto-dominates $s^{\prime}$ on $J_{\mathcal{G}}$ when:
$\forall j\in J_{\mathcal{G}},\ \sigma(s^{\prime},j)\leq\sigma(s,j)\quad\text{and}\quad\exists j\in J_{\mathcal{G}},\ \sigma(s^{\prime},j)<\sigma(s,j).$
(L: IV2, IV3.)
Proposition 3.35 (Interior Dominance Implies Universal Non-Rejection).
If a state $s$ interior-Pareto-dominates $s^{\prime}$ on a checked coordinate set $J$, and every admissible evaluator in the goal class is monotone on $J$, then no admissible evaluator strictly prefers $s^{\prime}$ over $s$ on that checked slice:
$\forall U\in\mathcal{U}_{\mathcal{G}},\ \forall a\in A,\ U(a,s^{\prime})\leq U(a,s).$
(L: IV5.)
Proposition 3.36 (Interior Verification Tractability).
[C:P3.36;R:H=tractable-structured] If membership in the checked coordinate set is decidable and interior dominance is decidable, then interior verification yields a polynomial-time checker with exact acceptance criterion:
$\exists\ \mathrm{verify}:S\times S\to\{0,1\},\quad\mathrm{verify}(s,s^{\prime})=1\iff\text{interior-dominance}(s,s^{\prime}).$
(L: IV7, IV4.)
Proposition 3.37 (Interior Dominance Is One-Sided).
Interior dominance is one-sided and does not imply full coordinate sufficiency: a counterexample pair outside the checked slice can still violate global optimizer equivalence. (L: IV6, IV8.)
Corollary 3.38 (Singleton Interior Certificate).
[C:C3.38;R:H=tractable-structured] For a singleton checked coordinate $i$, strict improvement on $i$ with class-monotonicity on $\{i\}$ is already a valid interior certificate of non-rejection. (L: IV9.)
[C:D3.33,3.34;P3.35,3.36,3.37;C3.38;R:AR]
Definition 3.39 (RLFF Objective (Evidence-Gated Report Reward)).
Given report type $r\in\{\mathsf{ABSTAIN},\mathsf{EXACT},\mathsf{EPSILON}(\varepsilon)\}$ and declared contract, RLFF assigns base reward by report type when evidence exists and applies an explicit inadmissibility penalty otherwise:
$\mathrm{Reward}(r)=\begin{cases}\mathrm{Base}(r),&\text{if evidence for }r\text{ exists},\\ -\mathrm{Penalty},&\text{otherwise}.\end{cases}$
(L: IC4, IC28, IC29.)
Proposition 3.40 (RLFF Maximizer Is Admissible).
If $\mathsf{ABSTAIN}$ reward strictly exceeds the inadmissibility branch, then any global RLFF maximizer must be evidence-backed and therefore admissible under the typed-claim discipline. (L: IC31, IC32.)
Corollary 3.41 (RLFF Abstain Preference Without Certificates).
If $\mathsf{EXACT}$ and $\mathsf{EPSILON}(\varepsilon)$ have no evidence objects while $\mathsf{ABSTAIN}$ beats the inadmissibility branch, then RLFF strictly prefers $\mathsf{ABSTAIN}$ to both report types. (L: IC30, DQ570.)
[C:D3.39;P3.40;C3.41;R:AR]
Definition 3.42 (Certainty Inflation).
For a typed report $r$, certainty inflation is the state of emitting $r$ without a matching evidence object. (L: IC1, IC3.)
Proposition 3.43 (Certainty Inflation Equals Report Inadmissibility).
For every typed report $r$:
$\text{CertaintyInflation}(r)\iff\neg\ \text{ClaimAdmissible}(r).$
For $r=\mathsf{EXACT}$ this is equivalently failure of exact competence. (L: IC8, IC22.)
Corollary 3.44 (Hardness-Blocked Regimes Induce Exact Certainty Inflation).
Under the declared hardness/non-collapse premises that block universal exact admissibility, every exact report is certainty-inflated. (L: CC19.)
[C:D3.42;P3.43;C3.44;R:CR]
Definition 3.45 (Raw-vs-Certified Bit Accounting).
Fix a declared contract and report type $r$. A report-bit model declares:
  • raw report bits $B_{\mathrm{raw}}(r)$ (payload bits in the asserted report token),
  • certificate bits $B_{\mathrm{cert}}(r)$ (bits allocated to the matching evidence class).
Evidence-gated overhead and certified total are then:
$B_{\mathrm{over}}(r)=\begin{cases}B_{\mathrm{cert}}(r),&\text{if evidence for }r\text{ exists},\\ 0,&\text{otherwise},\end{cases}\qquad B_{\mathrm{total}}(r)=B_{\mathrm{raw}}(r)+B_{\mathrm{over}}(r).$
(L: IC5, IC9, IC12.)
Proposition 3.46 (Raw/Certified Bit Split).
For every typed report $r$:
$B_{\mathrm{total}}(r)=B_{\mathrm{raw}}(r)+B_{\mathrm{over}}(r),$
with
$\neg\exists\ \text{evidence for }r\Rightarrow B_{\mathrm{over}}(r)=0,\qquad\exists\ \text{evidence for }r\Rightarrow B_{\mathrm{over}}(r)=B_{\mathrm{cert}}(r).$
Hence $B_{\mathrm{raw}}(r)\leq B_{\mathrm{total}}(r)$ always, and strict inequality holds when evidence exists with positive certificate-bit allocation. (L: CC11, IC11, IC10, IC16, IC15, IC13, IC14, DQ567.)
Theorem 3.47 (Exact Report: Certified-Bit Gap iff Admissibility).
Under a report-bit model with positive exact certificate-bit allocation:
$\text{ClaimAdmissible}(\mathsf{EXACT})\iff B_{\mathrm{raw}}(\mathsf{EXACT})<B_{\mathrm{total}}(\mathsf{EXACT}).$
Equivalently:
$B_{\mathrm{total}}(\mathsf{EXACT})=B_{\mathrm{raw}}(\mathsf{EXACT})\iff\text{ExactCertaintyInflation}.$
(L: CC18, CC20, CC17, DQ566.)
Corollary 3.48 (Finite Budget Eventually Blocks Exact Admissibility).
[C:C3.48;R:H=exp-lb-conditional] Fix a declared physical computer model with finite budget and positive per-bit cost, a declared operation-requirement lower-bound family, and a declared typed-report bit model. If:
  1. 1.
    exact admissibility implies certified-bit footprint at least the declared operation requirement at size $n$, and
  2. 2.
    exact admissibility implies physical feasibility of that certified footprint under the declared budget model,
then there exists $n_{0}$ such that for all $n\geq n_{0}$, exact admissibility fails. The same conclusion holds in the canonical exponential-requirement specialization. (L: DQ573, DQ574.)
Corollary 3.49 (Hardness-Blocked Exact Reports Are Raw-Only).
If exact reporting is inadmissible on the declared contract, then exact accounting collapses to raw-only:
$B_{\mathrm{total}}(\mathsf{EXACT})=B_{\mathrm{raw}}(\mathsf{EXACT}).$
(L: CC21, IC23.)
[C:D3.45;T3.47;P3.46;C3.49;R:AR]
Corollary 3.50 (No Uncertified Exact-Exhaustion Claim).
If exact competence does not hold on the declared regime/objective, then an $\mathsf{EXACT}$ report is inadmissible. (L: CC41.)
[C:P3.16;C3.50;R:AR]
Proposition 3.51 (Declared-Contract Separation: Selection vs. Validity).
For solver-side reporting, there are two formally distinct layers:
  • Declared-contract selection layer: choosing objective family $(\mathcal{R}_{\varepsilon})_{\varepsilon\geq 0}$, regime $\Gamma$, and assumption profile;
  • Validity layer: once declared, admissibility of $\mathsf{EXACT}/\mathsf{EPSILON}(\varepsilon)/\mathsf{ABSTAIN}$ reports is fixed by certificate-typed rules.
Thus the framework treats contract selection as an external declaration choice, and then enforces mechanically checked admissibility within that declared contract. (L: CC75, CC22, CC39.)
Proof.
The validity layer is Proposition 3.16: report admissibility is exactly tied to the matching competence certificate type. For exact physical claims outside carve-outs, Proposition 3.55 and Theorem 3.56 require explicit assumptions, and Corollary 3.57 blocks uncertified exact reports under declared non-collapse assumptions. Therefore, once the contract is declared, admissibility is mechanically determined by the typed rules and attached assumption profile.  $\blacksquare$
[C:T3.56;P3.16,3.55;C3.57;R:DC]
Definition 3.52 (Exact-Claim Excuse Carve-Outs).
For exact physical claims in this framework, we declare four explicit carve-outs:
  1. 1.
    trivial scope (empty in-scope family),
  2. 2.
    tractable class (exact competence available in the active regime),
  3. 3.
    stronger oracle/regime shift (the claim is made under strictly stronger access assumptions),
  4. 4.
    unbounded resources (budget bounds intentionally removed).
Definition 3.53 (Explicit Hardness-Assumption Profile).
An explicit hardness-assumption profile for declared class/regime $(\mathcal{R},\Gamma)$ consists of:
  • a non-collapse hypothesis,
  • a collapse implication from universal exact competence on $(\mathcal{R},\Gamma)$.
Definition 3.54 (Well-Typed Exact Physical Claim Policy).
An exact physical claim is well-typed only if either:
  • at least one carve-out from Definition 3.52 is explicitly declared, or
  • an explicit hardness-assumption profile (Definition 3.53) is explicitly declared.
Proposition 3.55 (Outside Carve-Outs, Explicit Assumptions Are Required).
If an exact physical claim is well-typed (Definition 3.54) and no carve-out from Definition 3.52 applies, then an explicit hardness-assumption profile must be present. (L: CC22.)
Theorem 3.56 (Claim-Integrity Meta-Theorem).
For declared class/regime objects, every admissible exact-claim policy outside the carve-outs requires explicit assumptions:
[…]
(L: CC22.)
Proof.
Immediate from Proposition 3.55; this is the universally quantified theorem-level restatement over the typed claim/regime objects.  $\blacksquare$
Corollary 3.57 (Outside Carve-Outs, Declared Hardness Blocks Exact Reports).
If no carve-out applies and the declared hardness-assumption profile holds, then $\mathsf{EXACT}$ reports are inadmissible for every solver on the declared class/regime. (L: CC39.)
[C:T3.56;P3.55;C3.57;R:CR]
Proposition 3.58 (Integrity–Competence Separation).
Integrity and competence are distinct predicates: integrity constrains asserted outputs, while competence adds non-abstaining coverage under resource bounds. (L: IC18, IC25.)
Proof.
Take the always-abstain solver $Q_{\bot}(x)=\mathsf{ABSTAIN}$ with any polynomial-time checker $V$. Definition 3.8 holds vacuously, so $(Q_{\bot},V)$ is integrity-preserving, but it fails Definition 3.9 whenever $\mathcal{X}_{\Gamma}\neq\emptyset$ because coverage fails. Hence integrity does not imply competence. The converse is immediate because competence includes integrity as a conjunct.  $\blacksquare$
Definition 3.59 (Attempted Competence Failure).
Fix an exact objective under regime $\Gamma$. A solver state is an attempted competence failure if:
  • integrity holds (Definition 3.8),
  • exact competence was actually attempted for the active scope/objective,
  • competence on $\Gamma$ fails for that exact objective (Definition 3.9).
Proposition 3.60 (Attempted-Competence Rationality Matrix).
Let $I,A,C\in\{0,1\}$ denote integrity, attempted exact competence, and competence available in the active regime. Policy verdict for persistent over-specification is:
  • if $I=0$: inadmissible,
  • if $I=1$ and $(A,C)=(1,0)$: conditionally rational,
  • if $I=1$ and $(A,C)\in\{(0,0),(0,1),(1,1)\}$: irrational for the same verified-cost objective.
Hence, in the integrity-preserving plane ($I=1$), exactly one cell is rational and three are irrational. (L: IC27, IC7, IC6)
This separation plus Proposition 3.60 is load-bearing for the regime-conditional trilemma used later: if exact competence is blocked by hardness in a declared regime after an attempted exact procedure, integrity forces one of three responses: abstain, weaken guarantees, or change regime assumptions.[C:P3.60;R:AR]
Mechanized status.
This separation is machine-checked in DecisionQuotient/IntegrityCompetence.lean via: IC18 and IC25; the attempted-competence matrix is mechanized via IC27, IC7, and IC6.
Proposition 3.61 (Integrity-Resource Bound).
Let $\Gamma$ be a declared regime whose in-scope family includes all instances of SUFFICIENCY-CHECK and whose declared resource bound is polynomial in the encoding length. If $\textsf{P}\neq\textsf{coNP}$, then no certifying solver is simultaneously integrity-preserving and competent on $\Gamma$ for exact sufficiency. Equivalently, for every integrity-preserving solver, at least one of the competence conjuncts must fail on $\Gamma$: either full non-abstaining coverage fails or the declared budget bound fails. (L: IC26, IC24, CC30.)
Proof.
[…]
Proposition 3.62 (Declared-Physics No-Universal-Exact-Certifier Schema).
Fix any declared physical/task class representable by relation $\mathcal{R}$ and regime $\Gamma$ in the certifying-solver model. If
$\big(\exists Q,\ \mathrm{CompetentOn}(\mathcal{R},\Gamma,Q)\big)\ \Rightarrow\ (\textsf{P}=\textsf{coNP}),$
then under $\textsf{P}\neq\textsf{coNP}$ there is no universally exact-competent solver for that declared class. (L: CC15.)
Corollary 3.63 (No Universal Exact-Claim Admissibility in Hardness-Blocked Classes).
Under the same premises as Proposition 3.62, and under the typed-claim discipline of Definition 3.15, an $\mathsf{EXACT}$ report is inadmissible for every solver on that declared class. Hence admissible reporting must use $\mathsf{ABSTAIN}$ or explicitly weakened guarantees. (L: CC38.)
[C:P3.62;C3.63;R:CR]
4 Computational Complexity of Decision-Relevant Uncertainty
This section establishes the computational complexity of information sufficiency, formalized as coordinate sufficiency, for an arbitrary decision problem $\mathcal{D}$ with factored state space. Because the problems below are parameterized by $\mathcal{D}$, and the $(A,S,U)$ tuple captures any problem with choices under structured information (Remark 1.1), the hardness results are universal: they apply to every problem with coordinate structure, not to a specific problem domain. We prove three main results:
  1. 1.
    SUFFICIENCY-CHECK is coNP-complete
  2. 2.
    MINIMUM-SUFFICIENT-SET is coNP-complete (the $\Sigma_{2}^{P}$ structure collapses)
  3. 3.
    ANCHOR-SUFFICIENCY (fixed coordinates) is $\Sigma_{2}^{P}$-complete
Under the model contract of Section 2.5 and the succinct encoding of Section 2.4, these results place exact relevance identification beyond NP-completeness: in the worst case, finding (or certifying) minimal decision-relevant sets is computationally intractable.
Reading convention for this section.
Each hardness theorem below is paired with a recovery boundary for the same decision question: when structural-access assumptions in Theorem 6.1 hold, polynomial-time exact procedures are recovered; when they fail in [S], the stated hardness applies. (L: CC60, CC36, CC4, CC71.)
4.1 Problem Definitions
Definition 4.1 (Decision Problem Encoding).
A decision problem instance is a tuple $(A,X_{1},\ldots,X_{n},U)$ where:
  • AA is a finite set of alternatives
  • $X_{1},\ldots,X_{n}$ are the coordinate domains, with state space $S=X_{1}\times\cdots\times X_{n}$
  • $U:A\times S\to\mathbb{Q}$ is the utility function (in the succinct encoding, $U$ is given as a Boolean circuit)
Succinct/circuit encoding.
When $U$ is given succinctly as a Boolean or arithmetic circuit, the instance size refers to the circuit description rather than the (possibly exponentially large) truth table. Evaluating $U(a,s)$ on a single pair $(a,s)$ can be done in time polynomial in the circuit size, but quantification over all states (for example, "for all ss") ranges over an exponentially large domain and is the source of the hardness results below.
Definition 4.2 (Optimizer Map).
For state $s\in S$, define:
$\text{Opt}(s):=\arg\max_{a\in A}U(a,s)$
Definition 4.3 (Sufficient Coordinate Set).
A coordinate set $I\subseteq\{1,\ldots,n\}$ is sufficient if:
$\forall s,s^{\prime}\in S:\quad s_{I}=s^{\prime}_{I}\implies\text{Opt}(s)=\text{Opt}(s^{\prime})$
where $s_{I}$ denotes the projection of $s$ onto coordinates in $I$.
Problem 4.4 (SUFFICIENCY-CHECK).
Input: Decision problem $(A,X_{1},\ldots,X_{n},U)$ and coordinate set $I\subseteq\{1,\ldots,n\}$
Question: Is $I$ sufficient?
Problem 4.5 (MINIMUM-SUFFICIENT-SET).
Input: Decision problem $(A,X_{1},\ldots,X_{n},U)$ and integer $k$
Question: Does there exist a sufficient set $I$ with $|I|\leq k$?
4.2 Hardness of SUFFICIENCY-CHECK
Classical coNP-completeness methodology follows standard NP/coNP reduction frameworks and polynomial-time many-one reducibility [47, 5]. We instantiate that framework for SUFFICIENCY-CHECK with a mechanized TAUTOLOGY reduction.
Theorem 4.6 (coNP-completeness of SUFFICIENCY-CHECK).
SUFFICIENCY-CHECK is coNP-complete. (L: CC61, CC60, L18.)
Source Target Key property preserved
TAUTOLOGY SUFFICIENCY-CHECK Tautology iff $\emptyset$ sufficient
$\exists\forall$-SAT ANCHOR-SUFFICIENCY Witness anchors iff formula true
Proof.
extbfMembership in coNP: The complementary problem INSUFFICIENCY is in NP. A nondeterministic verifier can check a witness pair $(s,s^{\prime})$ as follows:
  1. 1.
    Verify $s_{I}=s^{\prime}_{I}$ by projecting and comparing coordinates; this takes time polynomial in the input size.
  2. 2.
    Evaluate $U$ on every alternative for both $s$ and $s^{\prime}$. When $U$ is given as a circuit each evaluation runs in time polynomial in the circuit size; using these values the verifier checks whether $\text{Opt}(s)\neq\text{Opt}(s^{\prime})$.
Hence INSUFFICIENCY is in NP and SUFFICIENCY-CHECK is in coNP.
coNP-hardness: We reduce from TAUTOLOGY.
Given Boolean formula $\varphi(x_{1},\ldots,x_{n})$, construct a decision problem with:
  • Alternatives: $A=\{\text{accept},\text{reject}\}$
  • State space: $S=\{\text{reference}\}\cup\{0,1\}^{n}$ (equivalently, encode this as a product space with one extra coordinate $r\in\{0,1\}$ indicating whether the state is the reference state)
  • Utility:
    $U(\text{accept},\text{reference})$ $=1$
    $U(\text{reject},\text{reference})$ $=0$
    $U(\text{accept},a)$ $=\varphi(a)$
    $U(\text{reject},a)$ $=0\quad\text{for assignments }a\in\{0,1\}^{n}$
  • Query set: $I=\emptyset$
Claim: $I=\emptyset$ is sufficient $\iff$ $\varphi$ is a tautology.
($\Rightarrow$) Suppose $I$ is sufficient. Then $\text{Opt}(s)$ is constant over all states. Since $U(\text{accept},a)=\varphi(a)$ and $U(\text{reject},a)=0$:
  • $\text{Opt}(a)=\text{accept}$ when $\varphi(a)=1$
  • $\text{Opt}(a)=\{\text{accept},\text{reject}\}$ when $\varphi(a)=0$
For Opt to be constant, $\varphi(a)$ must be true for all assignments $a$; hence $\varphi$ is a tautology.
($\Leftarrow$) If $\varphi$ is a tautology, then $U(\text{accept},a)=1>0=U(\text{reject},a)$ for all assignments $a$. Thus $\text{Opt}(s)=\{\text{accept}\}$ for all states $s$, making $I=\emptyset$ sufficient.  $\blacksquare$
Immediate recovery boundary (SUFFICIENCY-CHECK).
Theorem 4.6 is the [S] general-case hardness statement. For the same sufficiency relation, Theorem 6.1 gives polynomial-time recovery under explicit-state, separable, and tree-structured regimes. (L: CC60, CC71, CC69, CC70, CC72.)
Mechanized strengthening (all coordinates relevant).
The reduction above establishes coNP-hardness using a single witness set $I=\emptyset$. For the ETH-based lower bound in Theorem 5.1, we additionally need worst-case instances where the minimal sufficient set has linear size.
We formalized a strengthened reduction in Lean 4: given a Boolean formula $\varphi$ over $n$ variables, construct a decision problem with $n$ coordinates such that if $\varphi$ is not a tautology then every coordinate is decision-relevant (so $k^{*}=n$). Intuitively, the construction places a copy of the base gadget at each coordinate and makes the global “accept” condition hold only when every coordinate’s local test succeeds; a single falsifying assignment at one coordinate therefore changes the global optimal set, witnessing that coordinate’s relevance. This strengthening is mechanized in Lean; see Appendix A.
4.3 Complexity of MINIMUM-SUFFICIENT-SET
Theorem 4.7 (MINIMUM-SUFFICIENT-SET is coNP-complete).
MINIMUM-SUFFICIENT-SET is coNP-complete. (L: CC34, CC36.)
Proof.
Structural observation: The $\exists\forall$ quantifier pattern suggests $\Sigma_{2}^{P}$:
$\exists I\,(|I|\leq k)\;\forall s,s^{\prime}\in S:\quad s_{I}=s^{\prime}_{I}\implies\text{Opt}(s)=\text{Opt}(s^{\prime})$
However, this collapses because sufficiency has a simple characterization.
Key lemma: In the Boolean-coordinate collapse model, a coordinate set $I$ is sufficient if and only if $I$ contains all relevant coordinates (proven formally as sufficient_iff_relevant_subset / sufficient_iff_relevantFinset_subset in Lean):
$\text{sufficient}(I)\iff\text{Relevant}\subseteq I$
[…]
Consequence: The minimum sufficient set is exactly the relevant-coordinate set. Thus MINIMUM-SUFFICIENT-SET asks: “Is the number of relevant coordinates at most kk?”
coNP membership: A witness that the answer is NO is a set of $k+1$ coordinates, each proven relevant (by exhibiting $s,s^{\prime}$ pairs). Verification is polynomial.
coNP-hardness: The $k=0$ case asks whether no coordinates are relevant, i.e., whether $\emptyset$ is sufficient. This is exactly SUFFICIENCY-CHECK, which is coNP-complete by Theorem 4.6.  $\blacksquare$
Immediate recovery boundary (MINIMUM-SUFFICIENT-SET).
Theorem 4.7 pairs with Theorem 4.14: exact minimization complexity is governed by relevance-cardinality once the collapse is applied. Recovery then depends on computing relevance efficiently, with structured-access assumptions from Theorem 6.1 giving the corresponding route for the underlying sufficiency computations. (L: CC36, CC34, CC35, S2P2, CC71.)
4.4 Anchor Sufficiency (Fixed Coordinates)
We also formalize a strengthened variant that fixes the coordinate set and asks whether there exists an assignment to those coordinates that makes the optimal action constant on the induced subcube.
Problem 4.8 (ANCHOR-SUFFICIENCY).
Input: Decision problem $(A,X_{1},\ldots,X_{n},U)$ and fixed coordinate set $I\subseteq\{1,\ldots,n\}$
Question: Does there exist an assignment $\alpha$ to $I$ such that $\text{Opt}(s)$ is constant for all states $s$ with $s_{I}=\alpha$?
The second-level hardness template used here follows canonical $\Sigma_{2}^{P}$ polynomial-hierarchy characterizations [47, 5], instantiated to the anchor formulation below.
Theorem 4.9 (ANCHOR-SUFFICIENCY is $\Sigma_{2}^{P}$-complete).
ANCHOR-SUFFICIENCY is $\Sigma_{2}^{P}$-complete (already for Boolean coordinate spaces). (L: CC5, CC4, L2.)
Proof.
Membership in $\Sigma_{2}^{P}$: The problem has the form
$\exists\alpha\;\forall s\in S:\;(s_{I}=\alpha)\implies\text{Opt}(s)=\text{Opt}(s_{\alpha}),$
which is an $\exists\forall$ pattern.
$\Sigma_{2}^{P}$-hardness: Reduce from $\exists\forall$-SAT. Given $\exists x\,\forall y\,\varphi(x,y)$ with $x\in\{0,1\}^{k}$ and $y\in\{0,1\}^{m}$, if $m=0$ we first pad with a dummy universal variable (satisfiability is preserved), construct a decision problem with:
  • Actions $A=\{\text{YES},\text{NO}\}$
  • State space $S=\{0,1\}^{k+m}$ representing $(x,y)$
  • Utility
[…]
  • Fixed coordinate set $I$ = the $x$-coordinates.
If $\exists x^{\star}\,\forall y\,\varphi(x^{\star},y)=1$, then for any $y$ we have $U(\text{YES})=2$ and $U(\text{NO})\leq 1$, so $\text{Opt}(x^{\star},y)=\{\text{YES}\}$ is constant. Conversely, fix $x$ and suppose $\exists y_{f}$ with $\varphi(x,y_{f})=0$.
  • If $\varphi(x,0^{m})=1$, then $\text{Opt}(x,0^{m})=\{\text{YES}\}$. The falsifying assignment must satisfy $y_{f}\neq 0^{m}$, where $U(\text{YES})=U(\text{NO})=0$, so $\text{Opt}(x,y_{f})=\{\text{YES},\text{NO}\}$.
  • If $\varphi(x,0^{m})=0$, then $\text{Opt}(x,0^{m})=\{\text{NO}\}$. After padding we have $m>0$, so choose any $y^{\prime}\neq 0^{m}$: either $\varphi(x,y^{\prime})=1$ (then $\text{Opt}(x,y^{\prime})=\{\text{YES}\}$) or $\varphi(x,y^{\prime})=0$ (then $\text{Opt}(x,y^{\prime})=\{\text{YES},\text{NO}\}$). In both subcases the optimal set differs from $\{\text{NO}\}$.
Hence the subcube for this $x$ is not constant. Thus an anchor assignment exists iff the $\exists\forall$-SAT instance is true.  $\blacksquare$
Immediate boundary statement (ANCHOR-SUFFICIENCY).
Theorem 4.9 remains a second-level hardness statement in the anchored formulation; unlike MINIMUM-SUFFICIENT-SET, no general collapse to relevance counting is established here, so the corresponding tractability status remains open in this model. (L: CC5, CC4, L2.)
4.5 Tractable Subcases
Despite the general hardness, several natural subcases admit efficient algorithms:
Proposition 4.10 (Small State Space).
When $|S|$ is polynomial in the input size (i.e., explicitly enumerable), MINIMUM-SUFFICIENT-SET is solvable in polynomial time.
Proof.
Compute $\text{Opt}(s)$ for all $s\in S$. The minimum sufficient set is exactly the set of coordinates that “matter” for the resulting function, computable by standard techniques.  $\blacksquare$
Proposition 4.11 (Linear Utility).
When $U(a,s)=w_{a}\cdot s$ for weight vectors $w_{a}\in\mathbb{Q}^{n}$, MINIMUM-SUFFICIENT-SET reduces to identifying coordinates where weight vectors differ.
4.6 Implications
Corollary 4.12 (Computational Bottleneck for Exact Minimization).
Under the succinct encoding, exact minimization of sufficient coordinate sets is coNP-hard via the $k=0$ case, and fixed-anchor minimization is $\Sigma_{2}^{P}$-complete. (L: CC36, CC4)
Proof.
The $k=0$ case of MINIMUM-SUFFICIENT-SET is SUFFICIENCY-CHECK (Theorem 4.6), giving coNP-hardness for exact minimization. The fixed-anchor variant is exactly Theorem 4.9.  $\blacksquare$
The modeling budget for deciding what to model is therefore a computationally constrained resource under this encoding.[C:Thm. 4.6, Thm. 4.9;R:[S]]
4.7 The succinct–minimal gap and physical agents
In complexity theory, a succinct representation of a combinatorial object is a Boolean circuit of size $\mathrm{poly}(n)$ that computes a function whose explicit (truth-table) form has size $2^{\Theta(n)}$. The term is standard but informal: it denotes a size class (polynomial in input variables), not a structural property. In particular, “succinct” does not mean minimal. The minimal circuit for a function is the smallest circuit that computes it, a rigorous, unique object. A succinct circuit may be far larger than the minimal one while still qualifying as “succinct” by the poly-size criterion.
The gap.
Let $C$ be a succinct circuit computing a utility function $U$, and let $C^{*}$ be the minimal circuit for the same function. The representation gap $\varepsilon=|C|-|C^{*}|$ measures how far the agent’s encoding is from the tightest possible encoding. This gap has no standard formal definition in the literature; it is the unnamed distance between an informal size class and a rigorous structural minimum.
Proposition 4.13 (Coordinate Representation Gap as a Formal Object).
In the Boolean-coordinate model, let $R_{\mathrm{rel}}=\texttt{relevantFinset}(dp)$ and define
$\varepsilon(I)\;:=\;|I\setminus R_{\mathrm{rel}}|+|R_{\mathrm{rel}}\setminus I|.$
Then:
$\varepsilon(I)=0\iff I=R_{\mathrm{rel}}\iff I\text{ is minimal sufficient}.$
(L: S2P3, S2P4, S2P5.)
The results of this paper give $\varepsilon$ formal complexity consequences. By Theorem 4.7, determining the minimal sufficient coordinate set is coNP-complete under the succinct encoding. By Theorem 5.1, under ETH there exist succinct instances requiring $2^{\Omega(n)}$ time to verify sufficiency. In the formal Boolean-coordinate collapse model, size-bounded $\varepsilon=0$ search is exactly the relevance-cardinality predicate (Lean: S2P.min_representationGap_zero_iff_relevant_card). Together: for an agent encoded as a succinct circuit, the worst-case cost of determining whether its own coordinate attention set is minimal is exponential in the number of input coordinates (under ETH). The gap $\varepsilon$ is not merely unnamed; it is computationally irreducible: no polynomial-time algorithm solves all succinct instances (assuming $\mathrm{P}\neq\textsf{coNP}$), and the hard instance family requires $2^{\Omega(n)}$ time.
Physical agents as circuits.
Any physical agent that selects actions based on sensory state can be modeled, in the static snapshot, as a circuit: sensory inputs are coordinates $X_{1},\ldots,X_{n}$; the agent’s internal computation maps states to actions; and the quality of that mapping is measured by a utility function UU.[C:Prop. 2.48, Prop. 2.41;R:[bridge-admissible snapshot]] The explicit-state encoding (truth table) of this mapping has size exponential in the number of sensory dimensions. No physical agent can store it.[C:Thm. 5.1;R:[E] vs [S+ETH]] Every physical agent is therefore a succinct encoding, a circuit whose size is bounded by physical resources (neurons, synapses, transistors, parameters) rather than by the size of the state space it navigates.[C:Thm. 5.1;R:[S]]
This is not a metaphor. A biological nervous system is a feedforward or recurrent circuit that takes sensory state as input and produces motor commands as output.[C:Prop. 2.48;R:[represented adjacent class]] A trained neural network is a parameterized circuit.[C:Prop. 2.48;R:[represented adjacent class]] In both cases, the agent is a succinct encoding of a utility-to-action mapping, and the question “is this agent attending to the right inputs?” is exactly SUFFICIENCY-CHECK under the succinct encoding.[C:Thm. 7.2, Thm. 4.6;R:[S]] The compression that makes a circuit physically realizable, with polynomial size rather than exponential size, is the same compression that makes self-verification exponentially costly in the worst case: the state space the circuit compressed away is precisely the domain that must be exhaustively checked to certify sufficiency.[C:Thm. 5.1, Prop. 5.3;R:[S+ETH],[Q_fin]]
The simplicity tax as $\varepsilon$.
The simplicity tax (Definition 9.2) measures $|\mathrm{Gap}(M,P)|=|R(P)\setminus A(M)|$: the number of decision-relevant coordinates that the agent’s model does not handle natively. In the $\varepsilon$ decomposition above, this is exactly the unattended-relevant component $|R_{\mathrm{rel}}\setminus I|$ (Lean: S2P.representationGap_missing_eq_gapCard, S2P.representationGap_eq_waste_plus_missing). For a physical agent modeled as a circuit, this is the coordinate-level manifestation of $\varepsilon$: the agent attends to a superset of what is necessary, or fails to attend to coordinates that are relevant, and faces worst-case cost exponential in $n$ to verify which case holds (Theorem 4.6). Theorem 9.3 then says the total relevance budget is conserved: unhandled coordinates are not eliminated, only redistributed as external per-site cost.
Consequences for learned circuits.
Large language models and deep networks are succinct circuits with $\mathrm{poly}(n)$ parameters.[C:Thm. 5.1;R:[S]] Mechanistic interpretability asks which internal components (attention heads, neurons, layers) are necessary for a given capability; this is SUFFICIENCY-CHECK applied to the circuit’s internal coordinate structure. Pruning research asks for the minimal subcircuit that preserves performance; this is MINIMUM-SUFFICIENT-SET. The results of this paper imply that these questions are coNP-hard in the worst case under succinct encoding.[C:Thm. 4.6, Thm. 4.7;R:[S]] Empirical progress on interpretability and pruning therefore depends on exploiting the structural properties of specific trained circuits (analogous to the tractable subcases of Theorem 6.1), not on solving the general problem.[C:Thm. 4.6, Thm. 6.1;R:[S] vs [tractable regimes]]
Scope and caveats.
The formal results above apply to the static, deterministic model fixed by C1–C4 (Section 2.5). Biological neural systems add noise, temporal dynamics, and synaptic plasticity; trained networks add stochastic optimization and distribution shift. These factors may alter empirical tractability relative to the static worst-case bounds. The bridge conditions of Proposition 2.40 delineate when static sufficiency results transfer to sequential settings; beyond those conditions, the governing complexity objects change (Propositions 2.432.45). The claims in this subsection are therefore regime-typed consequences of the formal results, not universal assertions about all possible agent architectures.[C:Prop. 2.40, Thm. 2.46, Prop. 2.48;R:[represented adjacent class]]
4.8 Quantifier Collapse for MINIMUM-SUFFICIENT-SET
Theorem 4.14 (Collapse of the Apparent $\exists\forall$ Structure).
The apparent second-level predicate
$\exists I\,(|I|\leq k)\;\forall s,s^{\prime}\in S:\;s_{I}=s^{\prime}_{I}\implies\text{Opt}(s)=\text{Opt}(s^{\prime})$
is equivalent to the coNP predicate $|\text{Relevant}|\leq k$, where
[…]
Consequently, MINIMUM-SUFFICIENT-SET is governed by coNP certificates rather than a genuine $\Sigma_{2}^{P}$ alternation. (L: CC34, CC35, S2P2.)
Proof.
By the formal equivalence sufficient_iff_relevant_subset (finite-set form sufficient_iff_relevantFinset_subset), a coordinate set $I$ is sufficient iff $\text{Relevant}\subseteq I$. Therefore:
$\exists I\;(|I|\leq k\wedge\text{sufficient}(I))\iff\exists I\;(|I|\leq k\wedge\text{Relevant}\subseteq I)\iff|\text{Relevant}|\leq k.$
So the existential-over-universal presentation collapses to counting the relevant coordinates.
A NO certificate for $|\text{Relevant}|\leq k$ is a list of $k+1$ distinct relevant coordinates, each witnessed by two states that differ only on that coordinate and yield different optimal sets; this is polynomially verifiable. Hence the resulting predicate is in coNP, matching Theorem 4.7.
This also clarifies why ANCHOR-SUFFICIENCY remains $\Sigma_{2}^{P}$-complete: once an anchor assignment is existentially chosen, the universal quantifier over the residual subcube does not collapse to a coordinate-counting predicate.  $\blacksquare$
5 Encoding-Regime Separation
The hardness results of Section 4 apply to worst-case instances under the succinct encoding. This section states an encoding-regime separation: an explicit-state upper bound versus a succinct-encoding worst-case lower bound, and an intermediate query-access obstruction family. The complexity class is a property of the encoding measured against a fixed decision question (Remark 2.21), not a property of the question itself: the same sufficiency question admits polynomial-time algorithms under one encoding and worst-case exponential cost under another.
Physical reading: the separation is an access-and-budget separation for the same decision semantics. Explicit access assumes full boundary observability; succinct access assumes compressed structure that may require exponential expansion to certify universally; query access assumes bounded measurement interfaces.[C:T5.1;P5.3,2.34;R:E,S+ETH,Qf,AR] (L: CC16,CC50,CT6.)
Model note.
Theorem 5.1 has Part 1 in [E] (time polynomial in $|S|$) and Part 2 in [S+ETH] (time exponential in $n$). Proposition 5.3 is [Q_fin] (finite-state Opt-oracle core), while Propositions 5.7 and 5.11 are [Q_bool] interface refinements (value-entry and state-batch). Relative to Definition 2.23, Propositions 5.8 and 5.15 are explicit simulation-transfer instances. The [E] vs [S+ETH] split in Theorem 5.1 is a same-question encoding/cost-model separation, not a bidirectional simulation claim between those regimes. These regimes are defined in Section 2.4 and are not directly comparable as functions of one numeric input-length parameter.
Theorem 5.1 (Explicit–Succinct Regime Separation).
Let $\mathcal{D}=(A,X_{1},\ldots,X_{n},U)$ be a decision problem with $|S|=N$ states. Let $k^{*}$ be the size of the minimal sufficient set.
  1. 1.
    [E] Explicit-state upper bound: Under the explicit-state encoding, SUFFICIENCY-CHECK is solvable in time polynomial in $N$ (e.g. $O(N^{2}|A|)$).
  2. 2.
    [S+ETH] Succinct lower bound (worst case): Assuming ETH, there exists a family of succinctly encoded instances with $n$ coordinates and minimal sufficient set size $k^{*}=n$ such that SUFFICIENCY-CHECK requires time $2^{\Omega(n)}$.
(L: CC16, CC23, CC24.)
Proof.
Part 1 (Explicit-state upper bound): Under the explicit-state encoding, SUFFICIENCY-CHECK is decidable in time polynomial in $N$ by direct enumeration: compute $\text{Opt}(s)$ for all $s\in S$ and then check all pairs $(s,s^{\prime})$ with $s_{I}=s^{\prime}_{I}$.
Equivalently, for any fixed $I$, the projection map $s\mapsto s_{I}$ has image size $|S_{I}|\leq|S|=N$, so any algorithm that iterates over projection classes (or over all state pairs) runs in polynomial time in $N$. Thus, in particular, when $k^{*}=O(\log N)$, SUFFICIENCY-CHECK is solvable in polynomial time under the explicit-state encoding.
Remark (bounded coordinate domains).
In the general model $S=\prod_{i}X_{i}$, for a fixed $I$ one always has $|S_{I}|\leq\prod_{i\in I}|X_{i}|$ (and $|S_{I}|\leq N$). If the coordinate domains are uniformly bounded, $|X_{i}|\leq d$ for all $i$, then $|S_{I}|\leq d^{|I|}$.
Part 2 (Succinct ETH lower bound, worst case): A strengthened version of the TAUTOLOGY reduction used in Theorem 4.6 produces a family of instances in which the minimal sufficient set has size $k^{*}=n$: given a Boolean formula $\varphi$ over $n$ variables, we construct a decision problem with $n$ coordinates such that if $\varphi$ is not a tautology then every coordinate is decision-relevant (thus $k^{*}=n$). This strengthening is mechanized in Lean (see Appendix A). Under the Exponential Time Hypothesis (ETH) [28, 5], TAUTOLOGY requires time $2^{\Omega(n)}$ in the succinct encoding, so SUFFICIENCY-CHECK inherits a $2^{\Omega(n)}$ worst-case lower bound via this reduction.  $\blacksquare$
Corollary 5.2 (Regime Separation (by Encoding)).
There is a clean separation between explicit-state tractability and succinct worst-case hardness (with respect to the encodings in Section 2.4):
  • Under the explicit-state encoding, SUFFICIENCY-CHECK is polynomial in $N=|S|$.
  • Under ETH, there exist succinctly encoded instances with $k^{*}=\Omega(n)$ (indeed $k^{*}=n$) for which SUFFICIENCY-CHECK requires $2^{\Omega(n)}$ time.
For Boolean coordinate spaces ($N=2^{n}$), the explicit-state bound is polynomial in $2^{n}$ (exponential in $n$), while under ETH the succinct lower bound yields $2^{\Omega(n)}$ time for the hard family in which $k^{*}=\Omega(n)$.
Proposition 5.3 (Finite-State Query Lower-Bound Core via Empty-Set Subproblem).
In the finite-state query regime [Q_fin], let $S$ be any finite state type with $|S|\geq 2$ and let $Q\subset S$ with $|Q|<|S|$. Then there exist two decision problems $\mathcal{D}_{\mathrm{yes}},\mathcal{D}_{\mathrm{no}}$ over the same state space such that:
  • their oracle views on all states in $Q$ are identical,
  • $\emptyset$ is sufficient for $\mathcal{D}_{\mathrm{yes}}$,
  • $\emptyset$ is not sufficient for $\mathcal{D}_{\mathrm{no}}$.
Consequently, no deterministic query procedure using fewer than $|S|$ state queries can solve SUFFICIENCY-CHECK on all such instances; the worst-case query complexity is $\Omega(|S|)$. (L: CC50, L7, L13.)
Proof.
This is exactly the finite-state indistinguishability theorem CC4. For any $|Q|<|S|$, there is an unqueried hidden state $s_{0}$ producing oracle-indistinguishable yes/no instances with opposite truth values on the $I=\emptyset$ subproblem. Since SUFFICIENCY-CHECK contains that subproblem, fewer than $|S|$ queries cannot be correct on both instances, yielding the $\Omega(|S|)$ worst-case lower bound.  $\blacksquare$
Proposition 5.4 (Checking–Witnessing Duality on the Empty-Set Core).
[C:P5.4;R:H=query-lb] Let $n\geq 1$ and let $P$ be a finite family of Boolean-state pairs. Assume $P$ is sound for refuting empty-set sufficiency in the sense that for every $\text{Opt}:\{0,1\}^{n}\to\{0,1\}$, whenever $\emptyset$ is not sufficient, some pair in $P$ witnesses disagreement of Opt. Then
$|P|\geq 2^{n-1}.$
Equivalently, any sound checker must inspect at least the witness budget $W(n)=2^{n-1}$, so checking budget $T(n)$ satisfies $T(n)\geq W(n)$. (L: DQ587, DQ588.)
Corollary 5.5 (Information Barrier in Query Regimes).
[C:C5.5;R:H=query-lb] For the fixed sufficiency question, finite query interfaces induce an information barrier:
  • Q_fin
    (Opt-oracle): with $|Q|<|S|$, there exist yes/no instances indistinguishable on all queried states but with opposite truth values on the $I=\emptyset$ subproblem.
  • Q_bool
    value-entry and state-batch interfaces preserve the same obstruction at cardinality $2^{n}$.
Hence the barrier is representational-access based: without querying enough of the hidden state support, exact certification is blocked by indistinguishability rather than by search-strategy choice. (L: CC27, CC29, CC28.)
Corollary 5.6 (Boolean-Coordinate Instantiation).
[C:C5.6;R:H=query-lb] In the Boolean-coordinate state space $S=\{0,1\}^{n}$, Proposition 5.3 yields the familiar $\Omega(2^{n})$ worst-case query lower bound for Opt-oracle access. (L: CC49; core: L6.)
Proposition 5.7 (Value-Entry Query Lower Bound).
In the mechanized Boolean value-entry query submodel [Q_bool], for any deterministic procedure using fewer than $2^{n}$ value-entry queries $(a,s)\mapsto U(a,s)$, there exist two queried-value-indistinguishable instances with opposite truth values for SUFFICIENCY-CHECK on the $I=\emptyset$ subproblem. Therefore worst-case value-entry query complexity is also $\Omega(2^{n})$. (L: L9, L19.)
Proof.
The theorem L2 constructs, for any query set of cardinality $2^{n}$, an unqueried hidden state $s_{0}$ and a yes/no instance pair that agree on all queried values but disagree on $\emptyset$-sufficiency. The auxiliary bound L6 ensures that fewer than $2^{n}$ value-entry queries cannot cover all states, so the indistinguishability argument applies in the worst case.  $\blacksquare$
Proposition 5.8 (Subproblem-to-Full Transfer Rule).
If every full-problem solver induces a solver for a fixed subproblem, then any lower bound for that subproblem lifts to the full problem. (L: CC58, CC59, CC52.)
This is the restriction-map instance of Definition 2.23: a solver for the full regime induces one for the restricted subproblem regime, so lower bounds transfer.
Proposition 5.9 (Randomized Robustness (Seedwise)).
[C:P5.9;R:H=query-lb] In [Q_bool], for any query set with cardinality $2^{n}$, the indistinguishable yes/no pair from Proposition 5.3 forces one decoding error per random seed for any seed-indexed decoder from oracle transcripts. Consequently, finite-support randomization does not remove the obstruction: averaging preserves a constant error floor on the hard pair. (L: L11, L12, L5.)
Proposition 5.10 (Randomized Robustness (Weighted Form)).
[C:P5.10;R:H=query-lb] For any finite-support seed weighting $\mu$, the same hard pair satisfies a weighted identity: the weighted sum of yes-error and no-error equals total seed weight. Hence randomization cannot collapse both errors simultaneously. (L: L23, L24.)
Proposition 5.11 (State-Batch Query Lower Bound).
In [Q_bool], the same $\Omega(2^{n})$ lower bound holds for a state-batch oracle that returns the full Boolean-action utility tuple at each queried state. (L: L8, L14.)
Proposition 5.12 (Finite-State Empty-Subproblem Generalization).
[C:P5.12;R:H=query-lb] The empty-subproblem indistinguishability lower-bound core extends from Boolean-vector state spaces to any finite state type with at least two states. (L: L7, L13.)
Proposition 5.13 (Adversary-Family Tightness by Full Scan).
[C:P5.13;R:H=query-lb] For the const/spike adversary family used in the query lower bounds, querying all states distinguishes the pair; thus the lower-bound family is tight up to constant factors under full-state scan. (L: L10.)
Proposition 5.14 (Weighted Query-Cost Transfer).
[C:P5.14;R:H=query-lb] Let $w(q)$ be per-query cost and $w_{\min}$ a lower bound on all queried costs. Any cardinality lower bound $|Q|\geq L$ lifts to weighted cost:
$\sum_{q\in Q}w(q)\ \geq\ w_{\min}\cdot L.$
(L: L21, L22.)
Proposition 5.15 (Oracle-Lattice Transfer: Batch $\Rightarrow$ Entry).
In [Q_bool], agreement on state-batch oracle views over touched states implies agreement on all value-entry views for the corresponding entry-query set. (L: L20, CC43.)
This is the oracle-transducer instance of Definition 2.23: batch transcripts induce entry transcripts while preserving indistinguishability for the target sufficiency question.
Proposition 5.16 (Oracle-Lattice Strictness: Opt vs Value Entries).
‘Opt‘-oracle views are strictly coarser than value-entry views: there exist instances with identical ‘Opt‘ views but distinguishable value entries. (L: L3, L4.)
Remark 5.17 (Instantiation of Definitions 3.1 and 3.2).
Theorem 5.1 and Propositions 5.35.7 keep the structural problem fixed (same sufficiency relation) and separate representational hardness by access regime: explicit-state access exposes the boundary $s\mapsto\text{Opt}(s)$, finite-state query access already yields $\Omega(|S|)$ lower bounds at the Opt-oracle level (Boolean instantiation: $\Omega(2^{n})$), value-entry/state-batch interfaces preserve the obstruction in [Q_bool], and succinct access can hide structure enough to force ETH-level worst-case cost on a hard family.
This regime-typed landscape identifies tractability under [E], a finite-state query lower-bound core under [Q_fin] with Boolean interface refinements under [Q_bool], and worst-case intractability under [S+ETH] for the same underlying decision question.
The structural reason for this separation is enumerable access. Under the explicit-state encoding, the mapping $s\mapsto\text{Opt}(s)$ is directly inspectable and sufficiency verification reduces to scanning state pairs, at cost polynomial in $|S|$. Under succinct encoding, the circuit compresses an exponential state space into polynomial size; universal verification over that compressed domain, namely “for all $s,s^{\prime}$ with $s_{I}=s^{\prime}_{I}$, does $\text{Opt}(s)=\text{Opt}(s^{\prime})$?”, carries worst-case cost exponential in $n$, because the domain the circuit compressed away cannot be reconstructed without undoing the compression.
5.1 Budgeted Physical Crossover
The [E] vs [S] split can be typed against an explicit physical budget. Let $E(n)$ and $S(n)$ denote explicit and succinct representation sizes for the same decision question at scale parameter $n$, and let $B$ be a hard representation budget. These crossover claims are interpreted through the physical-to-core encoding discipline of Section 2.7: physical assumptions are transferred explicitly into core assumptions, and encoded physical counterexamples map to core failures on the encoded slice.[C:P2.34;C2.35;R:AR] (L: CT6,CT8,CT4.)
Proposition 5.18 (Budgeted Crossover Witness).
[C:P5.18;R:H=exp-lb-conditional] If $E(n)>B$ and $S(n)\leq B$ for some $n$, then explicit representation is infeasible at $(B,n)$ while succinct representation remains feasible at $(B,n)$. (L: CC45, PBC1.)
Proof.
This is exactly the definitional split encoded in PBC1: explicit infeasibility is $B<E(n)$ and succinct feasibility is $S(n)\leq B$. The theorem CC2 returns both conjuncts.  $\blacksquare$
Proposition 5.19 (Above-Cap Crossover Existence).
[C:P5.19;R:H=exp-lb-conditional] Assume there is a global succinct-size cap $C$ and explicit size is unbounded:
$\forall n,\ S(n)\leq C,\qquad\forall B,\ \exists n,\ B<E(n).$
Then for every budget $B\geq C$, there exists a crossover witness at budget $B$. (L: CC44, PBC2.)
Proof.
Fix $B\geq C$. By unboundedness of $E$, choose $n$ with $B<E(n)$. By the cap assumption, $S(n)\leq C\leq B$, so $(B,n)$ is a crossover witness. The mechanized closure theorem is CC1.  $\blacksquare$
Proposition 5.20 (Crossover Does Not Imply Exact-Certification Competence).
[C:P5.20;R:H=exp-lb-conditional] Fix any collapse schema “exact certification competence $\Rightarrow$ complexity collapse” for the target exact-certification predicate. Under the corresponding non-collapse assumption, exact-certification competence is impossible even when a budgeted crossover witness exists. (L: CC46.)
Proof.
The mechanized theorem CC3 bundles the crossover split with the same logical closure used in Proposition 3.61: representational feasibility and certification hardness are distinct predicates.  $\blacksquare$
Proposition 5.21 (Crossover Policy Closure Under Integrity).
[C:P5.21;R:H=exp-lb-conditional] In the certifying-solver model (Definitions 3.83.9), assume:
  • a budgeted crossover witness at $(B,n)$,
  • the same collapse implication as Proposition 3.61,
  • a solver that is integrity-preserving for the declared relation.
Then either full in-scope non-abstaining coverage fails, or declared budget compliance fails. Equivalently, in uncertified hardness-blocked regions, integrity is compatible with $\mathsf{ABSTAIN}/\mathsf{UNKNOWN}$ but not with unconditional exact-competence claims. (L: CC47.)
Corollary 5.22 (Finite Budget Implies Eventual Exact-Certification Failure).
[C:C5.22;R:H=exp-lb-conditional] Fix a physical computer model with finite total budget and strictly positive per-bit physical cost. If required operation count for exact certification has an exponential lower bound, then there exists a threshold $n_{0}$ such that for all $n\geq n_{0}$, required physical work exceeds budget. For the canonical requirement model used in this paper, the same eventual-budget-failure conclusion holds. (L: DQ569, DQ568, DQ574.)
Proposition 5.23 (Least Divergence Point at Fixed Budget).
[C:P5.23;R:H=exp-lb-conditional] Fix a budgeted encoding model and budget $B$. If a crossover witness exists at budget $B$, then there is a least index $n_{\mathrm{crit}}$ such that crossover holds at $n_{\mathrm{crit}}$, and no smaller index has crossover. (L: DQ576.)
Proposition 5.24 (Eventual Explicit Infeasibility from Monotone Growth).
[C:P5.24;R:H=exp-lb-conditional] If explicit size is monotone in $n$ and already exceeds budget $B$ at some $n_{0}$, then explicit infeasibility holds for all $n\geq n_{0}$. (L: DQ577.)
Proposition 5.25 (Payoff Threshold as Eventual Split).
[C:P5.25;R:H=exp-lb-conditional] If, beyond $n_{0}$, explicit encoding is infeasible while succinct encoding remains feasible, then every $n\geq n_{0}$ is a crossover point. (L: DQ575, DQ579.)
Corollary 5.26 (No Universal Survivor Without Succinct Bound).
[C:C5.26;R:H=exp-lb-conditional] If both explicit and succinct encodings are unbounded, then for every fixed budget $B$, there exists an index where explicit is infeasible and an index where succinct is infeasible. (L: DQ578.)
Proposition 5.27 (Policy Closure at and Beyond Divergence).
[C:P5.27;R:H=exp-lb-conditional] Under the same integrity-resource collapse assumption as Proposition 5.21, policy closure holds at any fixed divergence point and uniformly beyond any threshold where crossover holds. (L: DQ580, DQ581.)
5.2 Thermodynamic Lift in Declared Model
The crossover and hardness statements are complexity-theoretic. The thermodynamic lift follows the same assumption discipline as the rest of the paper: claims are exactly those derivable from the declared conversion model and declared bit-operation lower-bound premises.
Proposition 5.28 (Bit-to-Energy/Carbon Lift).
[C:P5.28;R:H=cost-growth] Fix a declared thermodynamic conversion model with constants $(\lambda,\rho)$ where $\lambda$ lower-bounds energy per irreversible bit operation and $\rho$ lower-bounds carbon per joule. If a bit-operation lower bound $b_{\mathrm{LB}}\leq b_{\mathrm{used}}$ holds, then:
$E_{\mathrm{LB}}=\lambda b_{\mathrm{LB}}\leq\lambda b_{\mathrm{used}},\qquad C_{\mathrm{LB}}=\rho E_{\mathrm{LB}}\leq\rho(\lambda b_{\mathrm{used}}).$
(L: CC65, CC66.)
Proposition 5.29 (Hardness–Thermodynamics Bundle).
[C:P5.29;R:H=cost-growth] Under the same non-collapse/collapse schema used by Proposition 3.61, exact-certification competence is impossible; combined with a declared bit-operation lower bound, this yields simultaneous lower bounds on energy and carbon in the declared conversion model. (L: CC67.)
Proposition 5.30 (Mandatory Physical Cost).
[C:P5.30;R:H=cost-growth] Within the declared linear thermodynamic model, if
$\lambda>0,\quad\rho>0,\quad b_{\mathrm{LB}}>0,$
then both lower bounds are strictly positive:
$E_{\mathrm{LB}}=\lambda b_{\mathrm{LB}}>0,\qquad C_{\mathrm{LB}}=\rho E_{\mathrm{LB}}>0.$
(L: CC68.)
Proposition 5.31 (Conserved Additive Accounting).
[C:P5.31;R:H=cost-growth] Within the same declared linear model, lower-bound accounting is additive over composed bit-operation lower bounds:
$E_{\mathrm{LB}}(b_{1}+b_{2})=E_{\mathrm{LB}}(b_{1})+E_{\mathrm{LB}}(b_{2}),\qquad C_{\mathrm{LB}}(b_{1}+b_{2})=C_{\mathrm{LB}}(b_{1})+C_{\mathrm{LB}}(b_{2}).$
(L: CC64.)
6 Tractable Special Cases: When You Can Solve It
We distinguish the encodings of Section 2.4. The tractability results below state the model assumption explicitly. Structural insight: hardness dissolves exactly when the full decision boundary $s\mapsto\text{Opt}(s)$ is recoverable in polynomial time from the input representation; the three cases below instantiate this single principle. Concretely, each tractable regime corresponds to a specific structural insight (explicit boundary exposure, additive separability, or tree factorization) that removes the hardness witness; this supports reading the general-case hardness as missing structural access in the current representation rather than as an intrinsic semantic barrier. In physical terms, these are the regimes where required structure is directly observable or factorized so that verification work scales with exposed local structure rather than hidden global combinations on encoded physical instances.[C:T6.1,2.37;P2.34;R:E,TA,AR] (L: CC69,CC70,CC71,CC72,CT6,CT4.)
Theorem 6.1 (Tractable Subcases).
SUFFICIENCY-CHECK is polynomial-time solvable in the following cases:
  1. 1.
    Explicit-state encoding: The input contains the full utility table over $A\times S$. SUFFICIENCY-CHECK runs in $O(|S|^{2}|A|)$ time; if $|A|$ is constant, $O(|S|^{2})$.
  2. 2.
    Separable utility (any encoding): $U(a,s)=f(a)+g(s)$.
  3. 3.
    Tree-structured utility with explicit local factors (succinct structured encoding): There exists a rooted tree on coordinates and local functions $u_{i}$ such that
    $U(a,s)=\sum_{i}u_{i}\bigl(a,s_{i},s_{\mathrm{parent}(i)}\bigr),$
    with the root term depending only on $(a,s_{\mathrm{root}})$ and all $u_{i}$ given explicitly as part of the input.
(L: CC69, CC70, CC72, CC71.)
6.1 Explicit-State Encoding
Proof of Part 1.
Given the full table of $U(a,s)$, compute $\text{Opt}(s)$ for all $s\in S$ in $O(|S||A|)$ time. For SUFFICIENCY-CHECK on a given $I$, verify that for all pairs $(s,s^{\prime})$ with $s_{I}=s^{\prime}_{I}$, we have $\text{Opt}(s)=\text{Opt}(s^{\prime})$. This takes $O(|S|^{2}|A|)$ time by direct enumeration and is polynomial in the explicit input length. If $|A|$ is constant, the runtime is $O(|S|^{2})$.  $\blacksquare$
6.2 Separable Utility
Proof of Part 2.
If $U(a,s)=f(a)+g(s)$, then:
$\text{Opt}(s)=\arg\max_{a\in A}[f(a)+g(s)]=\arg\max_{a\in A}f(a)$
The optimal action is independent of the state. Thus $I=\emptyset$ is always sufficient.  $\blacksquare$
6.3 Tree-Structured Utility
Proof of Part 3.
Assume the tree decomposition and explicit local tables as stated. For each node $i$ and each value of its parent coordinate, compute the set of actions that are optimal for some assignment of the subtree rooted at $i$. This is a bottom-up dynamic program that combines local tables with child summaries; each table lookup is explicit in the input. A coordinate is relevant if and only if varying its value changes the resulting optimal action set. The total runtime is polynomial in $n$, $|A|$, and the size of the local tables.  $\blacksquare$
6.4 Practical Implications
Condition Examples
Explicit-state encoding Small or fully enumerated state spaces
Separable utility Additive costs, linear models
Tree-structured utility Hierarchies, causal trees
Proposition 6.2 (Heuristic Reusability).
[C:P6.2;R:H=tractable-structured] Let $\mathcal{C}$ be a structure class for which SUFFICIENCY-CHECK is polynomial (Theorem 6.1). If membership in $\mathcal{C}$ is decidable in polynomial time, then the combined detect-then-check procedure is a sound, polynomial-time heuristic applicable to all future instances in $\mathcal{C}$. For each subcase of Theorem 6.1, structure detection is polynomial (under the declared representation assumptions). (L: CC53, CC7, CC55, CC74.)
Remark 6.3 (Heuristics and Integrity).
Proposition 6.2 removes the complexity-theoretic obstruction to integrity-preserving action on detectable tractable instances: integrity no longer forces abstention. Whether the agent acts still requires competence (Definition 3.9): budget and coverage must also be satisfied. An agent that detects structure class $\mathcal{C}$, applies the corresponding polynomial checker, and abstains when detection fails maintains integrity; it never claims sufficiency without verification. Mistaking the heuristic for the general solution, claiming polynomial-time competence on a coNP-hard problem, violates integrity (Proposition 3.61).
7 Regime-Conditional Corollaries
This section derives regime-typed engineering corollaries from the core complexity theorems. Theorem 7.2 maps configuration simplification to SUFFICIENCY-CHECK; Theorems 4.6, 4.7, and 5.1 then yield exact minimization consequences under [S] and [S+ETH]. The same derivation applies to physical configuration/control systems once they are encoded into the C1–C4 decision interface: “configuration parameter” becomes any declared physical control or observation coordinate, and the same sufficiency predicate governs behavior preservation. [C:T7.2,2.37;P2.34;R:AR] (L: CR1,CT6,CT4.)
Regime tags used below follow Section 2.5: [S], [S+ETH], [E], [S_bool]. Any prescription that requires exact minimization is constrained by these theorem-level bounds.[C:T4.6,4.7,5.1;R:S,S+ETH] Theorem 7.4 implies that persistent failure to isolate a minimal sufficient set is a boundary-characterization signal in the current model, not a universal irreducibility claim.[C:T7.4;R:S]
Conditional rationality criterion.
For the objective “minimize verified total cost while preserving integrity,” persistent over-specification is admissible only in the attempted competence failure cell (Definition 3.59): after an attempted exact procedure, if exact irrelevance cannot be certified efficiently, integrity retains uncertified coordinates rather than excluding them.[C:P3.60;R:AR] When exact competence is available in the active regime (e.g., Theorem 6.1 and the exact-identifiability criterion), proven-irrelevant coordinates are removable; persistent over-specification is irrational for that same objective.[C:T6.1;P3.60;R:E,STR] Proposition 3.60 makes this explicit: in the integrity-preserving matrix, one cell is rational and three are irrational, so irrationality is the default verdict.[C:P3.60;R:AR] (L: IC18, IC25, IC7, CC71, S2P1.)
Remark 7.1 (Regime Contract for Engineering Corollaries).
All claims in this section are formal corollaries under the declared model assumptions.
  • Competence claims are indexed by the regime tuple of Definition 3.9; prescriptions are meaningful only relative to feasible resources under that regime (bounded-rationality feasibility discipline) [71, 5].[C:P3.58;R:AR]
  • Integrity (Definition 3.8) forbids overclaiming beyond certifiable outputs; $\mathsf{ABSTAIN}$$\mathsf{UNKNOWN}$ is first-class when certification is unavailable.[C:P3.58;R:AR]
  • Therefore, hardness results imply a regime-conditional trilemma: abstain, weaken guarantees (heuristics/approximation), or change encoding/structural assumptions to recover competence.[C:T4.6,5.1;P3.60;R:S,S+ETH]
7.1 Configuration Simplification is SUFFICIENCY-CHECK
Because SUFFICIENCY-CHECK is a meta-problem parameterized by an arbitrary decision problem $\mathcal{D}$, real engineering problems with factored configuration spaces are instances of the hardness landscape established above.
Theorem 7.2 (C1–C4: Configuration Simplification Reduces to SUFFICIENCY-CHECK).
Given a software system with configuration parameters $P=\{p_{1},\ldots,p_{n}\}$ and observed behaviors $B=\{b_{1},\ldots,b_{m}\}$, the problem of determining whether parameter subset $I\subseteq P$ preserves all behaviors is equivalent to SUFFICIENCY-CHECK. (L: CR1)
Proof.
Construct decision problem $\mathcal{D}=(A,X_{1},\ldots,X_{n},U)$ where:
  • Actions $A=B\cup\{\bot\}$, where $\bot$ is a sentinel “no-observed-behavior” action
  • Coordinates $X_{i}$ = domain of parameter $p_{i}$
  • State space $S=X_{1}\times\cdots\times X_{n}$
  • For $b\in B$, utility $U(b,s)=1$ if behavior $b$ occurs under configuration $s$, else 0
  • Sentinel utility $U(\bot,s)=1$ iff no behavior in $B$ occurs under configuration $s$, else 0
Then
[…]
So the optimizer map exactly encodes observed-behavior equivalence classes, including the empty-behavior case.
Coordinate set $I$ is sufficient iff:
$s_{I}=s^{\prime}_{I}\implies\text{Opt}(s)=\text{Opt}(s^{\prime})$
This holds iff configurations agreeing on parameters in $I$ exhibit identical behaviors.
Therefore, “does parameter subset $I$ preserve all behaviors?” is exactly SUFFICIENCY-CHECK for the constructed decision problem.  $\blacksquare$
Remark 7.3 (Reduction Scope).
The reduction above requires only:
  1. 1.
    a finite behavior set,
  2. 2.
    parameters with finite domains, and
  3. 3.
    an evaluable behavior map from configurations to achieved behaviors.
These are exactly the model-contract premises C1–C3 instantiated for configuration systems.
Theorem 7.4 (S: Over-Modeling as Boundary-Signal Corollary).
By contraposition of Theorem 7.2, if no coordinate set can be certified as exactly relevance-identifying (Definition 2.16) for the modeled system, then the decision boundary is not completely characterized by the current parameterization. (L: CC40, CC6)
Proof.
Assume the decision boundary were completely characterized by the current parameterization. Then, via Theorem 7.2, the corresponding sufficiency instance admits exact relevance membership, hence a coordinate set that satisfies Definition 2.16. Contraposition gives the claim: persistent failure of exact relevance identification signals incomplete characterization of decision relevance in the model.  $\blacksquare$
7.2 Cost Asymmetry Under ETH
We now prove a cost asymmetry result under the stated cost model and complexity constraints. [^1] Naive subset enumeration still gives an intuitive baseline of $O(2^{n})$ checks, but that is an algorithmic upper bound; the theorem below uses ETH for the lower-bound argument.
Theorem 7.5 (S+ETH: Cost-Asymmetry Consequence).
Consider an engineer specifying a system configuration with $n$ parameters. Let:
  • $C_{\text{over}}(k)$ = cost of maintaining $k$ extra parameters beyond minimal
  • $C_{\text{find}}(n)$ = cost of finding minimal sufficient parameter set
  • $C_{\text{under}}$ = expected cost of production failures from underspecification
Assume ETH in the succinct encoding model of Section 2.4. Then:
  1. 1.
    Exact identification of minimal sufficient sets has worst-case finding cost $C_{\text{find}}(n)=2^{\Omega(n)}$. (Under ETH, SUFFICIENCY-CHECK has a $2^{\Omega(n)}$ lower bound in the succinct model, and exact minimization subsumes this decision task.)
  2. 2.
    Maintenance cost is linear: $C_{\text{over}}(k)=O(k)$.
  3. 3.
    Under ETH, exponential finding cost dominates linear maintenance cost for sufficiently large $n$.
Therefore, there exists $n_{0}$ such that for all $n>n_{0}$, the finding-vs-maintenance asymmetry satisfies:
$C_{\text{over}}(k)<C_{\text{find}}(n)+C_{\text{under}}$
Within [S+ETH], persistent over-specification is unresolved boundary characterization, not proof that all included parameters are intrinsically necessary. Conversely, when exact competence is available in the active regime, certifiably irrelevant coordinates are removable; persistence is irrational for the same cost-minimization objective.[C:T7.5;P3.60;R:S+ETH,TA] (L: CC12, HD13.)
Proof.
Under ETH, the TAUTOLOGY reduction used in Theorem 4.6 yields a $2^{\Omega(n)}$ worst-case lower bound for SUFFICIENCY-CHECK in the succinct encoding. Any exact algorithm that outputs a minimum sufficient set can decide whether the optimum size is 0 by checking whether the returned set is empty; this is exactly the SUFFICIENCY-CHECK query for $I=\emptyset$. Hence exact minimal-set finding inherits the same exponential worst-case lower bound.
Maintaining $k$ extra parameters incurs:
  • Documentation cost: $O(k)$ entries
  • Testing cost: $O(k)$ test cases
  • Migration cost: $O(k)$ parameters to update
Total maintenance cost is $C_{\text{over}}(k)=O(k)$.
The eventual dominance step is mechanized in HD8: for fixed linear-overhead parameter $k$ and additive constant $C_{\text{under}}$ there is $n_{0}$ such that $k<2^{n}+C_{\text{under}}$ for all $n\geq n_{0}$. Therefore:
$C_{\text{over}}(k)\ll C_{\text{find}}(n)$
For any fixed nonnegative $C_{\text{under}}$, the asymptotic dominance inequality remains and only shifts the finite threshold $n_{0}$.  $\blacksquare$
Corollary 7.6 (Impossibility of Automated Configuration Minimization).
[C:C7.6;R:H=succinct-hard] Assuming $\textsf{P}\neq\textsf{coNP}$, there exists no polynomial-time algorithm that:
  1. 1.
    Takes an arbitrary configuration file with $n$ parameters
  2. 2.
    Identifies the minimal sufficient parameter subset
  3. 3.
    Guarantees correctness (no false negatives)
(L: CC37)
Proof.
Such an algorithm would solve MINIMUM-SUFFICIENT-SET in polynomial time, contradicting Theorem 4.7 (assuming $\textsf{P}\neq\textsf{coNP}$).  $\blacksquare$
Remark 7.7.
Corollary 7.6 is a formal boundary statement: an always-correct polynomial-time minimizer for arbitrary succinct inputs would collapse P and coNP.
Remark 7.8 (FPT Scope Caveat).
The practical force of worst-case hardness depends on instance structure, especially $k^{*}$. If SUFFICIENCY-CHECK is FPT in parameter $k^{*}$, then small-$k^{*}$ families can remain tractable even under succinct encodings. The strengthened mechanized gadget (all_coords_relevant_of_not_tautology) still proves existence of hard families with $k^{*}=n$; what is typical in deployed systems is an empirical question outside this formal model.
7.3 Regime-Conditional Operational Corollaries
Theorems 7.4 and 7.5 yield the following conditional operational consequences:
1. Conservative retention under unresolved relevance. If irrelevance cannot be certified efficiently under the active regime, the admissible policy is to retain a conservative superset of parameters.[C:T7.4;R:S]
2. Heuristic selection as weakened-guarantee mode. Under [S+ETH], exact global minimization can be exponentially costly in the worst case (Theorem 7.5); methods such as AIC/BIC/cross-validation instantiate the “weaken guarantees” branch of Definition 3.9 [1, 56, 66, 25, 50].[C:T7.5;R:S+ETH]
3. Full-parameter inclusion as an $O(n)$ upper-bound strategy. Under [S+ETH], if exact minimization is unresolved, including all $n$ parameters is an $O(n)$ maintenance upper-bound policy that avoids false irrelevance claims.[C:T7.5;R:S+ETH]
4. Irrationality outside attempted-competence-failure conditions. If the active regime admits exact competence (tractable structural-access conditions or exact relevance identifiability), or if exact competence was never actually attempted, continued over-specification is irrational; hardness no longer justifies it for the stated objective.[C:T6.1;P3.60;R:TA] (L: CC71, IC6, S2P1.)
These corollaries are direct consequences of the hardness/tractability landscape: over-specification is an attempted-competence-failure policy, not a default optimum. Outside that cell, the admissible exits are to shift to tractable regimes from Theorem 6.1 or adopt explicit approximation commitments.[C:T6.1;P3.60;R:RG]
8 Dominance Theorems for Hardness Placement
Regime for this section: the mechanized Boolean-coordinate model [S_bool] plus the architecture cost model defined below. The centralization-vs-distribution framing follows established software-architecture criteria about modular decomposition and long-run maintenance tradeoffs [49, 10, 6, 60]. Although phrased in software terms, the formal objects are substrate-agnostic: the same hardness-placement and amortization equations apply to physical architectures (for example, centralized vs distributed sensing/actuation computation) once mapped into the same coordinate/cost model. [C:P2.34;T2.37,8.16,9.3,9.6;R:AR,LC] (L: CT6,CT4,HD1,HD4,HD3.)
8.1 Over-Specification as Diagnostic Signal
Corollary 8.1 (Persistent Over-Specification).
[C:C8.1;R:H=succinct-hard] In the mechanized Boolean-coordinate model, if a coordinate is relevant and omitted from a candidate set $I$, then $I$ is not sufficient. (L: S2P6)
Proof.
This is the contrapositive of S2P3.  $\blacksquare$
Corollary 8.2 (Exact Relevance Identifiability Criterion).
[C:C8.2;R:H=succinct-hard] In the mechanized Boolean-coordinate model, for any candidate set $I$:
$I\text{ is exactly relevance-identifying}\iff\bigl(I\text{ is sufficient and }I\subseteq R_{\mathrm{rel}}\bigr),$
where $R_{\mathrm{rel}}$ is the full relevant-coordinate set. (L: S2P1)
Proof.
This is exactly S2P1, with $R_{\mathrm{rel}}=\texttt{relevantFinset}$.  $\blacksquare$
Remark 8.3 (Over-Specification Is Regime-Conditional, Not Default).
In this paper’s formal typing, persistent over-specification is admissible only under attempted competence failure: after an attempted exact procedure, exact relevance competence remains unavailable in the active regime, so integrity retains uncertified coordinates (Section 3; Section 7).[C:P3.60;R:AR] Once exact competence is available in the active regime (Corollaries 8.78.9 together with Corollary 8.2), certifiably irrelevant coordinates are removable; persistent over-specification is irrational for the same objective (verified total-cost minimization).[C:T6.1;P3.60;R:TA] (L: IC18, IC25, IC7, L15, L16, L17, S2P1.)
8.2 Architectural Decision Quotient
Definition 8.4 (Architectural Decision Quotient).
For a software system with configuration space $S$ and behavior space $B$:
$\text{ADQ}(I)=\frac{|\{b\in B:b\text{ achievable with some }s\text{ where }s_{I}\text{ fixed}\}|}{|B|}$
Proposition 8.5 (ADQ Ordering).
For coordinate sets $I,J$ in the same system, if $\mathrm{ADQ}(I)<\mathrm{ADQ}(J)$, then fixing $I$ leaves a strictly smaller achievable-behavior set than fixing $J$. (L: CC2)
Proof.
The denominator $|B|$ is shared. Thus $\mathrm{ADQ}(I)<\mathrm{ADQ}(J)$ is equivalent to a strict inequality between the corresponding achievable-behavior set cardinalities.  $\blacksquare$
8.3 Corollaries for Practice
Corollary 8.6 (Cardinality Criterion for Exact Minimization).
[C:C8.6;R:H=succinct-hard] In the mechanized Boolean-coordinate model, existence of a sufficient set of size at most $k$ is equivalent to the relevance set having cardinality at most $k$. (L: S2P2)
Proof.
By S2P2, sufficiency of size $\leq k$ is equivalent to a relevance-cardinality bound $\leq k$ in the Boolean-coordinate model.  $\blacksquare$
Corollary 8.7 (Bounded-Regime Tractability).
[C:C8.7;R:H=tractable-structured] When the bounded-action or explicit-state conditions of Theorem 6.1 hold, minimal modeling can be solved in polynomial time in the stated input size. (L: L15)
Proof.
This is the bounded-regime branch of Theorem 6.1, mechanized as L3.  $\blacksquare$
Corollary 8.8 (Separable-Utility Tractability).
[C:C8.8;R:H=tractable-structured] When utility is separable with explicit factors, sufficiency checking is polynomial in the explicit-state regime. (L: L16)
Proof.
This is the separable-utility branch of Theorem 6.1, mechanized as L4.  $\blacksquare$
Corollary 8.9 (Tree-Structured Tractability).
[C:C8.9;R:H=tractable-structured] When utility factors form a tree structure with explicit local factors, sufficiency checking is polynomial in the explicit-state regime. (L: L17)
Proof.
This is the tree-factor branch of Theorem 6.1, mechanized as L5.  $\blacksquare$
Corollary 8.10 (Mechanized Hard Family).
There is a machine-checked family of reduction instances where, for non-tautological source formulas, every coordinate is relevant ($k^{*}=n$), exhibiting worst-case boundary complexity. (L: L1)
Proof.
The strengthened reduction proves that non-tautological source formulas induce instances where every coordinate is relevant; this is mechanized as L1.  $\blacksquare$
8.4 Hardness Distribution: Right Place vs Wrong Place
Definition 8.11 (Hardness Distribution).
Let $P$ be a problem family under the succinct encoding of Section 2.4. In this section, baseline hardness $H(P;n)$ denotes worst-case computational step complexity on instances with $n$ coordinates (equivalently, as a function of succinct input length $L$) in the fixed encoding regime. A solution architecture $S$ partitions this baseline hardness into:
  • $H_{\text{central}}(S)$: hardness paid once, at design time or in a shared component
  • $H_{\text{distributed}}(S)$: hardness paid per use site
For $n$ use sites, total realized hardness is:
$H_{\text{total}}(S)=H_{\text{central}}(S)+n\cdot H_{\text{distributed}}(S)$
Proposition 8.12 (Baseline Lower-Bound Principle).
[C:P8.12;R:H=cost-growth] For any problem family $P$ measured by $H(P;n)$ above, any solution architecture $S$ and any number of use sites $n\geq 1$, if $H_{\text{total}}(S)$ is measured in the same worst-case step units over the same input family, then:
$H_{\text{total}}(S)=H_{\text{central}}(S)+n\cdot H_{\text{distributed}}(S)\geq H(P;n).$
For SUFFICIENCY-CHECK, Theorem 5.1 provides the baseline on the hard succinct family: $H(\text{SUFFICIENCY-CHECK};n)=2^{\Omega(n)}$ under ETH. (L: HD20)
Proof.
By definition, $H(P;n)$ is a worst-case lower bound for correct solutions in this encoding regime and cost metric. Any such solution architecture decomposes total realized work as $H_{\text{central}}+n\cdot H_{\text{distributed}}$, so that total cannot fall below the baseline.  $\blacksquare$
Definition 8.13 (Hardness Efficiency).
The hardness efficiency of solution $S$ with $n$ use sites is:
$\eta(S,n)=\frac{H_{\text{central}}(S)}{H_{\text{central}}(S)+n\cdot H_{\text{distributed}}(S)}$
(L: HD10)
Proposition 8.14 (Efficiency Equivalence).
[C:P8.14;R:H=cost-growth] For fixed $n$ and positive total hardness, larger $\eta(S,n)$ is equivalent to a larger central share of realized hardness. (L: HD10)
Proof.
From Definition 8.13, $\eta(S,n)$ is exactly the fraction of total realized hardness paid centrally.  $\blacksquare$
Definition 8.15 (Right vs Wrong Hardness Placement).
For a solution architecture $S$ in this linear model:
$\text{right hardness}\iff H_{\mathrm{distributed}}(S)=0,\qquad\text{wrong hardness}\iff H_{\mathrm{distributed}}(S)>0.$
(L: HD11, HD12)
Theorem 8.16 (Centralization Dominance).
Let $S_{\mathrm{right}},S_{\mathrm{wrong}}$ be architectures over the same problem family with
$H_{\mathrm{distributed}}(S_{\mathrm{right}})=0,\quad H_{\mathrm{central}}(S_{\mathrm{right}})>0,\quad H_{\mathrm{distributed}}(S_{\mathrm{wrong}})>0,$
and let $n>\max\!\bigl(1,H_{\mathrm{central}}(S_{\mathrm{right}})\bigr)$. Then:
  1. 1.
    Lower total realized hardness:
    $H_{\text{total}}(S_{\mathrm{right}})<H_{\text{total}}(S_{\mathrm{wrong}})$
  2. 2.
    Fewer error sites: errors in centralized components affect 1 location; errors in distributed components affect $n$ locations
  3. 3.
    Quantified leverage: moving one unit of work from distributed to central saves exactly n1n-1 units of total realized hardness
(L: HD1, HD2)
Proof.
(1) and (2) are exactly the bundled theorem HD1. (3) is exactly HD2.  $\blacksquare$
Corollary 8.17 (Right-Place vs Wrong-Place Hardness).
[C:C8.17;R:H=cost-growth] In the linear model, a right-hardness architecture strictly dominates a wrong-hardness architecture once use-site count exceeds central one-time hardness. Formally, for architectures $S_{\mathrm{right}},S_{\mathrm{wrong}}$ over the same problem family, if $S_{\mathrm{right}}$ has right hardness, $S_{\mathrm{wrong}}$ has wrong hardness, and $n>H_{\mathrm{central}}(S_{\mathrm{right}})$, then
$H_{\mathrm{central}}(S_{\mathrm{right}})+n\,H_{\mathrm{distributed}}(S_{\mathrm{right}})<H_{\mathrm{central}}(S_{\mathrm{wrong}})+n\,H_{\mathrm{distributed}}(S_{\mathrm{wrong}}).$
(L: HD16)
Proof.
This is the mechanized theorem HD10.  $\blacksquare$
Proposition 8.18 (Dominance Modes).
[C:P8.18;R:H=cost-growth] This section uses two linear-model dominance modes plus generalized nonlinear dominance and boundary modes:
  1. 1.
    Strict threshold dominance: Corollary 8.17 gives strict inequality once $n>H_{\mathrm{central}}(S_{\mathrm{right}})$.
  2. 2.
    Global weak dominance: under the decomposition identity used in HD3, centralized hardness placement is never worse for all $n\geq 1$.
  3. 3.
    Generalized nonlinear dominance: under bounded-vs-growing site-cost assumptions (Theorem 8.22), right placement strictly dominates beyond a finite threshold without assuming linear per-site cost.
  4. 4.
    Generalized boundary mode: without those growth-separation assumptions, strict dominance is not guaranteed (Proposition 8.24).
Proof.
Part (1) is Corollary 8.17. Part (2) is exactly HD3. Part (3) is Theorem 8.22. Part (4) is Proposition 8.24.  $\blacksquare$
Illustrative Instantiation (Type Systems). Consider a capability $C$ (e.g., provenance tracking) with one-time central cost $H_{\text{central}}$ and per-site manual cost $H_{\text{distributed}}$:
Approach $H_{\text{central}}$ $H_{\text{distributed}}$
Native type system support High (learning cost) Low (type checker enforces)
Manual implementation Low (no new concepts) High (reimplement per site)
The table is schematic; the formal statement is Corollary 8.19. The type-theoretic intuition behind this instantiation is consistent with classic accounts of abstraction and static interface guarantees [12, 53, 51, 36, 62, 22].
Corollary 8.19 (Type-System Threshold).
[C:C8.19;R:H=cost-growth] For the formal native-vs-manual architecture instance, native support has lower total realized cost for all
$n>H_{\mathrm{baseline}}(P),$
where $H_{\mathrm{baseline}}(P)$ corresponds to the Lean identifier intrinsicDOF(P) in module HardnessDistribution. (L: HD14)
Proof.
Immediate from HD9.  $\blacksquare$
8.5 Extension: Non-Additive Site-Cost Models
Definition 8.20 (Generalized Site Accumulation).
Let $C_{S}:\mathbb{N}\to\mathbb{N}$ be a per-site accumulation function for architecture $S$. Define generalized total realized hardness by
$H_{\text{total}}^{\mathrm{gen}}(S,n)=H_{\text{central}}(S)+C_{S}(n).$
Definition 8.21 (Eventual Saturation).
A cost function $f:\mathbb{N}\to\mathbb{N}$ is eventually saturating if there exists $N$ such that for all $n\geq N$, $f(n)=f(N)$.
Theorem 8.22 (Generalized Dominance by Growth Separation).
[C:T8.22;R:H=cost-growth] Let
$H_{\text{total}}^{\mathrm{gen}}(S,n)=H_{\text{central}}(S)+C_{S}(n).$
For two architectures $S_{\mathrm{right}},S_{\mathrm{wrong}}$, suppose there exists $B\in\mathbb{N}$ such that:
  1. 1.
    $C_{S_{\mathrm{right}}}(m)\leq B$ for all $m$ (bounded right-side per-site accumulation),
  2. 2.
    $m\leq C_{S_{\mathrm{wrong}}}(m)$ for all $m$ (identity-lower-bounded wrong-side growth).
Then for every
$n>H_{\text{central}}(S_{\mathrm{right}})+B,$
one has
$H_{\text{total}}^{\mathrm{gen}}(S_{\mathrm{right}},n)<H_{\text{total}}^{\mathrm{gen}}(S_{\mathrm{wrong}},n).$
(L: HD8)
Proof.
This is exactly the mechanized theorem HD6.  $\blacksquare$
Corollary 8.23 (Eventual Generalized Dominance).
[C:C8.23;R:H=cost-growth] If condition (1) above holds and there exists $N$ such that condition (2) holds for all $m\geq N$, then there exists $N_{0}$ such that for all $n\geq N_{0}$,
$H_{\text{total}}^{\mathrm{gen}}(S_{\mathrm{right}},n)<H_{\text{total}}^{\mathrm{gen}}(S_{\mathrm{wrong}},n).$
(L: HD9)
Proof.
Immediate from HD7.  $\blacksquare$
Proposition 8.24 (Generalized Assumption Boundary).
[C:P8.24;R:H=cost-growth] In the generalized model, strict right-vs-wrong dominance is not unconditional. There are explicit counterexamples:
  1. 1.
    If wrong-side growth lower bounds are dropped, right-side strict dominance can fail for all $n$.
  2. 2.
    If right-side boundedness is dropped, strict dominance can fail for all $n$ even when wrong-side growth is linear.
(L: HD7, HD6)
Proof.
This is exactly the pair of mechanized boundary theorems listed above.  $\blacksquare$
Theorem 8.25 (Linear Model: Saturation iff Zero Distributed Hardness).
[C:T8.25;R:H=cost-growth] In the linear model of this section,
$H_{\text{total}}(S,n)=H_{\text{central}}(S)+n\cdot H_{\text{distributed}}(S),$
the function $n\mapsto H_{\text{total}}(S,n)$ is eventually saturating if and only if $H_{\text{distributed}}(S)=0$. (L: HD19)
Proof.
This is exactly the mechanized equivalence theorem above.  $\blacksquare$
Theorem 8.26 (Generalized Model: Saturation is Possible).
[C:T8.26;R:H=cost-growth] There exists a generalized site-cost model with eventual saturation. In particular, for
$C_{K}(n)=\begin{cases}n,&n\leq K\\ K,&n>K,\end{cases}$
both $C_{K}$ and $n\mapsto H_{\text{central}}+C_{K}(n)$ are eventually saturating. (L: HD17, HD5)
Proof.
This is the explicit construction mechanized in Lean.  $\blacksquare$
Corollary 8.27 (Positive Linear Slope Cannot Represent Saturation).
[C:C8.27;R:H=cost-growth] No positive-slope linear per-site model can represent the saturating family above for all $n$. (L: HD15)
Proof.
This follows from the mechanized theorem that any linear representation of the saturating family must have zero slope.  $\blacksquare$
Mechanized strengthening reference.
The strengthened all-coordinates-relevant reduction is presented in Section 4 (“Mechanized strengthening”) and formalized in Reduction_AllCoords.lean via all_coords_relevant_of_not_tautology.
The next section develops the major practical consequence of this framework: the Simplicity Tax Theorem.
9 Conservation Law for Decision-Relevant Coordinates
The load-bearing fact in this section is not the set identity itself; it is the difficulty of shrinking the required set $R(P)$ in the first place. By Theorem 4.6 (and Theorem 4.7 for minimization), exact relevance identification is intractable in the worst case under succinct encoding. The identities below therefore quantify how unresolved relevance is redistributed between central and per-site work.[C:T4.6,4.7;R:S] For encoded physical systems, the same redistribution laws apply under declared physical-to-core transport: unresolved coordinates remain per-site external work, while resolved coordinates are paid once in centralized structure.[C:P2.34;T9.3,9.5,9.6;R:AR,LC] (L: CT6,HD4,HD21,HD3.)
Definition 9.1.
Let $R(P)$ be the required dimensions (those affecting Opt) and $A(M)$ the dimensions model $M$ handles natively. The expressive gap is $\text{Gap}(M,P)=R(P)\setminus A(M)$.
Definition 9.2 (Simplicity Tax).
The simplicity tax is the size of the expressive gap:
$\text{SimplicityTax}(M,P):=|\text{Gap}(M,P)|.$
Theorem 9.3 (Redistribution Identity).
$|\text{Gap}(M,P)|+|R(P)\cap A(M)|=|R(P)|$. The total cannot be reduced, only redistributed between “handled natively” and “handled externally.” (L: HD4)
Proof.
In the finite-coordinate model this is the exact set-cardinality identity
$|R\setminus A|+|R\cap A|=|R|,$
formalized as HD5.  $\blacksquare$
Remark 9.4 (Why this is nontrivial in context).
The algebraic identity in Theorem 9.3 is elementary. Its force comes from upstream hardness: reducing $|R(P)|$ by exact relevance minimization is worst-case intractable under the succinct encoding, so redistribution is often the only tractable lever available.
Theorem 9.5 (Linear Growth).
For $n$ decision sites:
$\text{TotalExternalWork}=n\times\text{SimplicityTax}(M,P).$
(L: HD21)
Proof.
This is by definition of per-site externalization and is mechanized as HD12.  $\blacksquare$
Theorem 9.6 (Amortization).
Let $H_{\text{central}}$ be the one-time cost of using a complete model. There exists
$n^{*}=\left\lfloor\frac{H_{\text{central}}}{\text{SimplicityTax}(M,P)}\right\rfloor$
such that for $n>n^{*}$, the complete model has lower total cost. (L: HD3)
Proof.
For positive per-site tax, the threshold inequality
$n>\left\lfloor\frac{H_{\text{central}}}{\text{SimplicityTax}}\right\rfloor\implies H_{\text{central}}<n\cdot\text{SimplicityTax}$
is mechanized as HD4.  $\blacksquare$
Corollary 9.7 (Gap Externalization).
[C:C9.7;R:H=cost-growth] If $\text{Gap}(M,P)\neq\emptyset$, then external handling cost scales linearly with the number of decision sites. (L: HD21, HD18)
Proof.
The exact linear form is HD12. When the gap is nonempty (positive tax), monotone growth with $n$ is HD11.  $\blacksquare$
Corollary 9.8 (Exact Minimization Criterion).
[C:C9.8;R:H=succinct-hard] For mechanized Boolean-coordinate instances, “there exists a sufficient set of size at most kk” is equivalent to “the relevant-coordinate set has cardinality at most kk.” (L: S2P2)
Proof.
This is S2P2.  $\blacksquare$
Appendix A provides theorem statements and module paths for the corresponding Lean formalization.
10 Related Work
10.1 Formalized Complexity and Mechanized Proof Infrastructure
Formalizing complexity-theoretic arguments in proof assistants remains comparatively sparse. Lean and Mathlib provide the current host environment for this artifact [16, 17, 67]. Closest mechanized precedents include verified computability/complexity developments in Coq and Isabelle [21, 32, 45, 44, 26], and certifying toolchain work that treats proofs as portable machine-checkable evidence [34, 43].
Recent LLM-evaluation work also highlights inference nondeterminism from numerical precision and hardware configuration, reinforcing the value of proof-level claims whose validity does not depend on stochastic inference trajectories [72].
10.2 Computational Decision Theory
The complexity of decision-making has been studied extensively. Foundational treatments of computational complexity and strategic decision settings establish the baseline used here [47, 5]. Our work extends this to the meta-question of identifying relevant information. Decision-theoretic and information-selection framing in this paper also sits in the tradition of statistical sufficiency and signaling/information economics [20, 65, 42, 40, 29].
Large-sample Bayesian-network structure learning and causal identification are already known to be hard in adjacent formulations [13, 61, 30]. Our object differs: we classify coordinate sufficiency for optimal-action invariance, with mechanized reductions and regime typing.
Relation to prior hardness results.
Closest adjacent results are feature-selection/model-selection hardness results for predictive subset selection [9, 3]. Relative to those works, this paper changes two formal objects: (i) the reductions are machine-checked (TAUTOLOGY and $\exists\forall$-SAT mappings with explicit polynomial bounds), and (ii) the output is a regime-typed hardness/tractability map for decision relevance under explicit encoding assumptions. The target predicate here is decision relevance, not predictive compression.
10.3 Succinct Representations and Regime Separation
Representation-sensitive complexity is established in planning and decision-process theory: classical and compactly represented MDP/planning problems exhibit sharp complexity shifts under different input models [48, 41, 37]. The explicit-vs-succinct separation in this paper is the same representation-sensitive phenomenon for the coordinate-sufficiency predicate.
The distinction in this paper is the object and scope of the classification: we classify decision relevance (sufficiency, anchor sufficiency, and minimum sufficient sets) for a fixed decision relation, with theorem-level regime typing and mechanized reduction chains.
10.4 Oracle and Query-Access Lower Bounds
Query-access lower bounds are a standard source of computational hardness transfer [18, 11]. Our [Q_fin] and [Q_bool] results are in this tradition, but specialized to the same sufficiency predicate used throughout the paper: they establish access-obstruction lower bounds while keeping the underlying decision relation fixed across regimes.
[…]
10.5 Energy Landscapes and Chemical Reaction Networks
The geometric interpretation of decision complexity connects to the physics of energy landscapes. Wales established a widely used picture in which molecular configurations form high-dimensional landscapes with basins and transition saddles [68, 69, 70]. Protein-folding theory uses the same landscape language via funnel structure toward native states [46]. In this paper’s formal model, the corresponding core object is coordinate-projection invariance of the optimizer map: sufficiency asks whether fixing coordinates preserves Opt, while insufficiency is witnessed by agreeing projections with different optimal sets (Definition 2.4, Proposition 2.6).[C:D2.4;P2.6;R:DM](L: DP7,DP8.)
Section 2.7 makes this transfer rule theorem-level: encoded physical claims are derived from core claims through an explicit assumption-transfer map, and an encoded physical counterexample refutes the corresponding core rule on that encoded slice.[C:P2.34;C2.35;R:AR](L: CT6,CT5,CT8,CT7.)
Under that mapping, the explicit-vs-succinct split is the mechanized statement of access asymmetry: [E] admits polynomial verification on explicit state access, while [S+ETH] admits worst-case exponential lower bounds on succinct inputs (Theorem 5.1); finite query access already exhibits indistinguishability barriers (Proposition 5.3). The complexity separation is mechanized.[C:T5.1;P5.3;R:E,S+ETH,Q_fin](L: CC16,CC23,CC24,CC50.)
The geometric core is mechanized directly: product-space slice cardinality, hypercube node-count identity, and node-vs-edge separation are proved in Lean at the decision abstraction level, with edge witnesses tied exactly to non-sufficiency.(L: DQ3, L25, L68, DP21.)
The circuit-to-decision instantiation is also mechanized in a separate bridge layer: geometry and dynamics are represented as typed components of a circuit object; decision circuits add an explicit interpretation layer; molecule-level constructors instantiate both views.(L: DQ519, DQ518, DQ517, DQ155, DQ154, DQ159, DQ160.)
Chemical reaction networks provide a concrete physical class where this encoding applies directly. Prior work establishes hardness for output optimization, reconfiguration, and reconstruction in CRN settings [4, 2, 19]. Those domain-specific CRN complexity results are external literature claims, not mechanized in this Lean artifact. The mechanized statement here is the transfer principle: once a CRN decision task is encoded as a C1–C4 decision problem, it inherits the same sufficiency hardness/tractability regime map (Theorems 7.2, 4.6, 6.1).[C:T7.2,4.6,6.1;R:E,S,TA](L: CR1,CC61,CC69,CC70,CC71,CC72.)
The physical bridge is not a single-domain add-on: Theorem 2.37 composes law-objective semantics, reduction equivalence, and hardness lower-bound accounting in one mechanized theorem, so physical instantiation and complexity claims share one proof chain.[C:T2.37;R:AR](L: CT4.)
For matter-only dynamics, the law-induced objective specialization is mechanized for arbitrary universes (arbitrary state/action types and transition-feasibility relations): the utility schema is induced from transition laws, and under a strict allowed-vs-blocked gap with nonempty feasible set, the optimizer equals the feasible-action set; with logical determinism at the action layer, the optimizer is singleton.[C:P2.36;R:AR](L: DQ8, DQ5, DQ7, DQ6.)
The interface-time semantics are also explicit and mechanized: timed states are $\mathbb{N}$-indexed, one decision event is exactly one tick, run length equals elapsed interface time, and this unit-time law is invariant across substrate tags under interface-consistency assumptions.[C:P2.28,2.29,2.30,2.31;R:AR](L: DQ548,DQ537,DQ539,DQ534,DQ541,DQ549.)
Regime-side interface machinery is also formalized as typed access objects with an explicit simulation preorder and certificate-upgrade statements. The access-channel and five-way composition theorems are encoded as assumption-explicit composition laws rather than informal metaphors.[C:P2.34;T2.37;R:AR](L: PC10, PC12, PC15, PC17, PC19, PC20, PC21.)
10.6 Thermodynamic Costs and Landauer’s Principle
The thermodynamic lift in Section 5 converts declared bit-operation lower bounds into declared energy/carbon lower bounds via linear conversion constants. In this artifact, those conversion consequences are mechanized conditionally through Propositions 5.28, 5.29, 5.30, and 5.31. Landauer’s principle and reversible-computation context [33, 7] define the physical assumption set; the constants are declared premises, not Lean-derived constants.[C:P5.28,5.29,5.30,5.31;R:AR](L: CC65,CC66,CC67,CC68,CC64.)
Coupling this to integrity yields a mechanized policy condition: when exact competence is hardness-blocked in the declared regime, integrity requires abstention or explicitly weakened guarantees (Propositions 3.61, 3.60). A stronger behavioral claim such as “abstention is globally utility-optimal for every objective” is not currently mechanized and remains objective-dependent.[C:P3.61,3.60;R:AR](L: IC26,IC24,CC30,IC27,IC7,IC6.)
10.7 Feature Selection
In machine learning, feature selection asks which input features are relevant for prediction. This is known to be NP-hard in general [9, 25]. Our results show the decision-theoretic analog is coNP-complete for both checking and minimization. [C:T4.6,4.7;R:S](L: CC61,CC36.)
10.8 Value of Information
The value of information (VOI) framework [27, 52] quantifies the maximum rational payment for information. Our work addresses a different question: not the value of information, but the complexity of identifying which information has value.
10.9 Model Selection
Statistical model selection (AIC/BIC/cross-validation) provides practical heuristics for choosing among models [1, 56, 66]. Our results formalize the regime-level reason heuristic selection appears: without added structural assumptions, exact optimal model selection inherits worst-case intractability, so heuristic methods implement explicit weakened-guarantee policies for unresolved structure. [C:T4.6,5.1,6.1;R:E,S,S+ETH](L: CC61,CC16,CC69,CC70,CC71.)
10.10 Information Compression vs Certification Length
The paper now makes an explicit two-part information accounting object (Definition 3.45): raw report bits and evidence-backed certification bits. This sharpens the information-theoretic reading of typed claims: compression of report payload is not equivalent to compression of certifiable truth conditions [57, 14, 64].
From a zero-error viewpoint, this distinction is structural rather than stylistic: compressed representations can remain short while exact decodability depends on confusability structure and admissible decoding conditions [58, 31, 38, 15]. The same separation appears here between short reports and evidence-backed exact claims.
The rate-distortion and MDL lines make the same split in different language: there is a difference between achievable compression rate, computational procedure for obtaining/optimizing that rate, and certifiable model adequacy [59, 8, 54, 24, 35].
Formally, Theorem 3.47 and Corollary 3.49 characterize the exact-report boundary: admissible exact reports require a strict certified-bit gap above raw payload, while hardness-blocked exact reports collapse to raw-only accounting. This directly links representational succinctness to certification cost under the declared contract, rather than treating “short reports” as automatically high-information claims. [C:T3.47;P3.46;C3.49;R:AR](L: CC18,CC11,CC21.)
10.11 Certifying Outputs and Proof-Carrying Claims
Our integrity layer matches the certifying-algorithms pattern: algorithms emit candidate outputs together with certificates that can be checked quickly, separating producing claims from verifying claims [39, 43]. In this paper, Definition 3.8 is exactly that soundness discipline.
At the systems level, this is the same architecture as proof-carrying code: a producer ships evidence and a consumer runs a small checker before accepting the claim [43, 39]. Our competence definition adds the regime-specific coverage/resource requirement that certifying soundness alone does not provide.
The feasibility qualifier in Definition 3.9 yields the bounded-rationality constraint used in this paper: admissible policy is constrained by computational feasibility under the declared resource model (Remark 2.25) [71, 5].[C:P3.58;R:AR] (L: IC18.)
Cryptographic-verifiability perspective (scope).
The role of cryptography in this paper is structural verifiability, not secrecy: the relevant analogy is certificate-carrying claims with lightweight verification, not confidential encoding [23, 43]. Concretely, the typed-report discipline is modeled as a game-based reporting interface: a prover emits $(r,\pi)$ and a verifier accepts exactly when $\pi$ certifies the declared report type under the active contract. In that model, Evidence–Admissibility Equivalence gives completeness for admissible report types, while Typed Claim Admissibility plus Exact Claims Require Exact Evidence give soundness against exact overclaiming. Certified-confidence constraints then act as evidence-gated admissibility constraints on top of that verifier interface, so raw payload bits and certified bits represent distinct resources rather than stylistic notation variants. [C:P3.18,3.16,3.21,3.23;R:AR](L: IC2,CC75,IC35,IC38.)
Three-axis integration.
In the cited literature, these pillars are treated separately: representation-sensitive hardness, query-access lower bounds, and certifying soundness disciplines. This paper composes all three for the same decision-relevance object in one regime-typed and machine-checked framework. [C:T5.1;P5.3,3.16;R:S+ETH,Qf,AR](L: CC16,CC50,CC75.)
11 Conclusion
Methodology and Disclosure
Role of LLMs in this work. This paper was developed through human-AI collaboration. The author provided the core intuitions: the connection between decision-relevance and computational complexity, the conjecture that SUFFICIENCY-CHECK is coNP-complete, and the insight that the $\Sigma_{2}^{\textsf{P}}$ structure collapses for MINIMUM-SUFFICIENT-SET. Large language models (Claude, GPT-4) served as implementation partners for proof drafting, Lean formalization, and generation.
The Lean 4 proofs were iteratively refined: the author specified the target statements, the LLM proposed proof strategies, and the Lean compiler served as the arbiter of correctness. The complexity-theoretic reductions required careful human oversight to ensure the polynomial bounds were correctly established.
What the author contributed: The problem formulations (SUFFICIENCY-CHECK, MINIMUM-SUFFICIENT-SET, ANCHOR-SUFFICIENCY), the hardness conjectures, the tractability conditions, and the connection to over-modeling in engineering practice.
What LLMs contributed: drafting, Lean tactic exploration, reduction construction assistance, and prose refinement.
The proofs are machine-checked; their validity is independent of generation method. This methodology is disclosed for academic transparency.
Summary of Results
This paper establishes the computational complexity of coordinate sufficiency problems:
  • SUFFICIENCY-CHECK is coNP-complete (Theorem 4.6)
  • MINIMUM-SUFFICIENT-SET is coNP-complete (Theorem 4.7)
  • ANCHOR-SUFFICIENCY is $\Sigma_{2}^{\textsf{P}}$-complete (Theorem 4.9)
  • An encoding-regime separation contrasts explicit-state polynomial-time (polynomial in $|S|$) with a succinct worst-case ETH lower bound witnessed by a hard family with $k^{*}=n$ (Theorem 5.1)
  • Full intermediate query-access lower bounds are formalized as a finite-state Opt-oracle core ($\Omega(|S|)$, Boolean instantiation $\Omega(2^{n})$) plus Boolean-interface refinements for value-entry and state-batch access, with explicit subproblem-to-full transfer, weighted randomized robustness, and oracle-lattice transfer/strictness closures (Propositions 5.35.16)
  • Tractable subcases exist for explicit-state encoding, separable utility, and tree-structured utility with explicit local factors (Theorem 6.1)
  • Discrete-time interface semantics are formalized: time is $\mathbb{N}$-valued, decision events are exactly one-tick transitions, run length equals elapsed time, and substrate-consistent realizations preserve the same one-unit event law (Propositions 2.282.31)
  • Physical-system transport is formalized: encoded physical claims are derived from core rules through explicit assumption transfer, physical counterexamples map to encoded core failures, and the bundled bridge theorem composes law semantics, reduction equivalence, and required-work lower bounds (Proposition 2.34, Corollary 2.35, Theorem 2.37)
[C:T4.6,4.7,4.9,5.1,6.1,2.37;P5.3,5.7,5.11,5.8,5.15,5.16,2.28,2.29,2.30,2.31,2.34;C2.35;R:E,S,S+ETH,Qf,Qb,AR] (L: CC61,CC36,CC5,CC16,CC69,CC50,L9,L8,L11,L21,L20,DQ548,DQ537,DQ539,DQ541,CT6,CT8,CT4.)
These results place the problem of identifying decision-relevant coordinates at the first and second levels of the polynomial hierarchy.
Beyond classification, the paper contributes a formal claim-typing framework (Section 3): structural complexity is a property of the fixed decision question, while representational hardness is regime-conditional access cost. This is why encoding-regime changes can move practical hardness without changing the underlying semantics. [C:P2.20;T3.4;R:AR] (L: CC62,CC76.)
The reduction constructions and key equivalence theorems are machine-checked in Lean 4 (see Appendix A for proof listings). The formalization verifies that the TAUTOLOGY reduction correctly maps tautologies to sufficient coordinate sets; complexity classifications (coNP-completeness, $\Sigma_{2}^{\textsf{P}}$-completeness) follow by composition with standard complexity-theoretic results (TAUTOLOGY is coNP-complete, $\exists\forall$-SAT is $\Sigma_{2}^{\textsf{P}}$-complete). The strengthened gadget showing that non-tautologies yield instances with all coordinates relevant is also formalized. [C:T4.6,4.7,4.9,5.1;R:S,S+ETH] (L: CC61,CC36,CC5,CC16,CC24.)
Complexity Characterization
The results provide precise complexity characterizations within the formal model:
  1. 1.
    Exact bounds. SUFFICIENCY-CHECK is coNP-complete, both coNP-hard and in coNP.
  2. 2.
    Constructive reductions. The reductions from TAUTOLOGY and $\exists\forall$-SAT are explicit and machine-checked.
  3. 3.
    Encoding-regime separation. Under [E], SUFFICIENCY-CHECK is polynomial in $|S|$. Under [S+ETH], there exist succinct worst-case instances (with $k^{*}=n$) requiring $2^{\Omega(n)}$ time. Under [Q_fin], the Opt-oracle core has $\Omega(|S|)$ worst-case query complexity (Boolean instantiation $\Omega(2^{n})$), and under [Q_bool] the value-entry/state-batch refinements preserve the obstruction with weighted-cost transfer closures.
[C:T4.6,5.1;P5.3,5.7,5.11,5.8;R:E,S,S+ETH,Qf,Qb] (L: CC61,CC16,CC50,L9,L8,L11.)
The Complexity Redistribution Corollary
Section 9 develops a quantitative consequence: when a problem requires $k$ dimensions and a model handles only $j<k$ natively, the remaining kjk-j dimensions must be handled externally at each decision site. For $n$ sites, total external work is $(k-j)\times n$.[C:T9.5;R:LC] (L: HD21.)
The set identity is elementary; its operational content comes from composition with the hardness results on exact relevance minimization. This redistribution corollary is formalized in Lean 4 (HardnessDistribution.lean), proving:
  • Redistribution identity: complexity burden cannot be eliminated by omission, only moved between native handling and external handling
  • Dominance: complete models have lower total work than incomplete models
  • Amortization: there exists a threshold $n^{*}$ beyond which higher-dimensional models have lower total cost
[C:T9.3,8.16,9.6;R:LC] (L: HD4,HD1,HD3,HD21.)
Open Questions
The landscape above is complete for the static sufficiency class under C1–C4 and the declared regimes; the items below are adjacent-class extensions or secondary refinements. Several questions remain for future work:
  • Fixed-parameter tractability (primary): Is SUFFICIENCY-CHECK FPT when parameterized by the minimal sufficient-set size $k^{*}$, or is it W[2]-hard under this parameterization?
[…]
  • Average-case complexity: What is the complexity under natural distributions on decision problems?
  • Typed-confidence calibration: For signaled reports $(r,g,p_{\mathrm{self}},p_{\mathrm{cert}})$, characterize proper scoring and calibration objectives that preserve signal consistency and typed admissibility.
  • Learning cost formalization: Can central cost $H_{\text{central}}$ be formalized as the rank of a concept matroid, making the amortization threshold precisely computable?
Practical Corollaries
The practical corollaries are regime-indexed and theorem-indexed:
  • [E] and structured regimes: polynomial-time exact procedures exist (Theorem 6.1).
  • [Q_fin]/[Q_bool] query-access lower bounds: worst-case Opt-oracle complexity is $\Omega(|S|)$ in the finite-state core (Boolean instantiation $\Omega(2^{n})$), and value-entry/state-batch interfaces satisfy the same obstruction in the mechanized Boolean refinement (Propositions 5.35.16), with randomized weighted robustness and oracle-lattice closures.
  • [S+ETH] hard families: exact minimization inherits exponential worst-case cost (Theorem 5.1 together with Theorem 4.6).
  • [S_bool] mechanized criterion: minimization reduces to relevance-cardinality constraints (Corollary 8.6).
  • Redistribution consequences: omitted native coverage externalizes work with explicit growth/amortization laws (Theorems 9.39.6).
  • Physical-system instantiation: once a physical class is encoded into the declared decision interface, the same regime-typed hardness/tractability map and integrity constraints apply to that class (Section 2.7).[C:P2.34;C2.35;T2.37;R:AR](L: CT6,CT8,CT4.)
Hence the design choice is typed: enforce a tractable regime, or adopt weakened guarantees with explicit verification boundaries.[C:T6.1,5.1;R:RG](L: CC71,CC16.) Equivalently, under the attempted-competence matrix, over-specification is rational only in attempted-competence-failure cells; once exact competence is available in the active regime (or no attempted exact competence was made), persistent over-specification is irrational for the same verified-cost objective.[C:P3.60;R:AR](L: IC27.) By Proposition 3.60, the integrity-preserving matrix contains 3 irrational cells and 1 rational cell.[C:P3.60;R:AR](L: IC27.)
Epistemic Alignment Implications for Agent Training
Within this paper’s formalism, epistemic alignment is a claim-discipline property: outputs are asserted only when certifiable, and otherwise emitted as abstain reports (optionally carrying tentative guesses and self-reflected confidence under zero certified confidence). Formally, this is the integrity/competence split of Section 3, not a full value-alignment claim.[C:D3.8,3.9;P3.58;R:TR](L: IC18.)
Truth-Objective Abstention Requirement.
If a solver-agent declares a truth-preserving objective for claims about relation $\mathcal{R}$ in regime $\Gamma$, uncertified outputs are admissible only as $\mathsf{ABSTAIN}/\mathsf{UNKNOWN}$ rather than asserted answers. Under the signal discipline, this includes abstain-with-guess/self-confidence reports while keeping certification evidence-gated.[C:D3.19,3.20;P3.23,3.29;C3.25;R:AR](L: IC46,IC38,IC41.) Equivalently, training policies that suppress abstention in uncertified regions violate the typed claim discipline under the declared task model.[C:D3.8;P3.58,3.61,3.60;R:AR](L: IC18,CC30,IC27.)
This yields a training doctrine for solver-agents under declared tasks/regimes:
  1. 1.
    Treat $\mathsf{ABSTAIN}/\mathsf{UNKNOWN}$ as first-class admissible outputs, not errors.[C:D3.8;R:TR]
  2. 2.
    For signaled reports, maintain two channels: self-reflected confidence may be emitted, but certified confidence is positive only with evidence.[C:D3.20;P3.23,3.30;R:AR](L: IC38,IC47.)
  3. 3.
    Treat runtime disclosure as scalar-witnessed: always emit exact step-count, and emit completion percentage only when a declared bound exists.[C:P3.26,3.27,3.28;R:AR](L: IC42,IC35,IC44.)
  4. 4.
    Optimize competent coverage subject to integrity constraints, rather than optimizing answer-rate alone.[C:D3.9;P3.58;R:AR](L: IC18.)
  5. 5.
    In hardness-blocked regimes, abstention or explicit guarantee-weakening is structurally required unless regime assumptions are changed.[C:T5.1;P3.61,3.60;R:S,S+ETH](L: CC16,CC30,IC27.)
Operationally, a reporting schema consistent with this framework tracks at least five quantities: integrity violations, competent non-abstaining coverage on the declared scope, abstention quality under regime shift, the calibration gap between self-reflected and certified confidence channels, and scalar runtime witness distributions (with bound-qualified fraction semantics when bounds are declared). This is consistent with bounded-rationality feasibility discipline under declared resource models [71, 5].
Appendix A Lean 4 Proof Listings
The complete Lean 4 formalization is available in the companion artifact (Zenodo DOI listed on the title page). The mechanization consists of 13391 lines across 58 files, with 597 theorem/lemma statements.
Handle IDs. Inline theorem metadata now cites compact IDs (for example, HD6, CC12, IC4) instead of full theorem constants. The full ID-to-handle mapping is listed in Section A.1.
A.1 Lean Handle ID Map
ID Full Lean handle
CC1 DecisionQuotient.ClaimClosure.RegimeSimulation
CC2 DecisionQuotient.ClaimClosure.adq_ordering
CC3 DecisionQuotient.ClaimClosure.agent_transfer_licensed_iff_snapshot
CC4 DecisionQuotient.ClaimClosure.anchor_sigma2p_complete_conditional
CC5 DecisionQuotient.ClaimClosure.anchor_sigma2p_reduction_core
CC6 DecisionQuotient.ClaimClosure.boundaryCharacterized_iff_exists_sufficient_subset
CC7 DecisionQuotient.ClaimClosure.bounded_actions_detectable
CC8 DecisionQuotient.ClaimClosure.bridge_boundary_represented_family
CC9 DecisionQuotient.ClaimClosure.bridge_failure_witness_non_one_step
CC10 DecisionQuotient.ClaimClosure.bridge_transfer_iff_one_step_class
CC11 DecisionQuotient.ClaimClosure.certified_total_bits_split_core
CC12 DecisionQuotient.ClaimClosure.cost_asymmetry_eth_conditional
CC13 DecisionQuotient.ClaimClosure.declaredBudgetSlice
CC14 DecisionQuotient.ClaimClosure.declaredRegimeFamily_complete
CC15 DecisionQuotient.ClaimClosure.declared_physics_no_universal_exact_certifier_core
CC16 DecisionQuotient.ClaimClosure.dichotomy_conditional
CC17 DecisionQuotient.ClaimClosure.epsilon_admissible_iff_raw_lt_certified_total_core
CC18 DecisionQuotient.ClaimClosure.exact_admissible_iff_raw_lt_certified_total_core
CC19 DecisionQuotient.ClaimClosure.exact_certainty_inflation_under_hardness_core
CC20 DecisionQuotient.ClaimClosure.exact_raw_eq_certified_iff_certainty_inflation_core
CC21 DecisionQuotient.ClaimClosure.exact_raw_only_of_no_exact_admissible_core
CC22 DecisionQuotient.ClaimClosure.explicit_assumptions_required_of_not_excused_core
CC23 DecisionQuotient.ClaimClosure.explicit_state_upper_core
CC24 DecisionQuotient.ClaimClosure.hard_family_all_coords_core
CC25 DecisionQuotient.ClaimClosure.horizonTwoWitness_immediate_empty_sufficient
CC26 DecisionQuotient.ClaimClosure.horizon_gt_one_bridge_can_fail_on_sufficiency
CC27 DecisionQuotient.ClaimClosure.information_barrier_opt_oracle_core
CC28 DecisionQuotient.ClaimClosure.information_barrier_state_batch_core
CC29 DecisionQuotient.ClaimClosure.information_barrier_value_entry_core
CC30 DecisionQuotient.ClaimClosure.integrity_resource_bound_for_sufficiency
CC31 DecisionQuotient.ClaimClosure.integrity_universal_applicability_core
CC32 DecisionQuotient.ClaimClosure.meta_coordinate_irrelevant_of_invariance_on_declared_slice
CC33 DecisionQuotient.ClaimClosure.meta_coordinate_not_relevant_on_declared_slice
CC34 DecisionQuotient.ClaimClosure.minsuff_collapse_core
CC36 DecisionQuotient.ClaimClosure.minsuff_conp_complete_conditional
CC37 DecisionQuotient.ClaimClosure.no_auto_minimize_of_p_neq_conp
CC38 DecisionQuotient.ClaimClosure.no_exact_claim_admissible_under_hardness_core
CC39 DecisionQuotient.ClaimClosure.no_exact_claim_under_declared_assumptions_unless_excused_core
CC40 DecisionQuotient.ClaimClosure.no_exact_identifier_implies_not_boundary_characterized
CC41 DecisionQuotient.ClaimClosure.no_uncertified_exact_claim_core
CC42 DecisionQuotient.ClaimClosure.one_step_bridge
CC43 DecisionQuotient.ClaimClosure.oracle_lattice_transfer_as_regime_simulation
CC44 DecisionQuotient.ClaimClosure.physical_crossover_above_cap_core
CC45 DecisionQuotient.ClaimClosure.physical_crossover_core
CC46 DecisionQuotient.ClaimClosure.physical_crossover_hardness_core
CC47 DecisionQuotient.ClaimClosure.physical_crossover_policy_core
CC48 DecisionQuotient.ClaimClosure.process_bridge_failure_witness
CC49 DecisionQuotient.ClaimClosure.query_obstruction_boolean_corollary
CC50 DecisionQuotient.ClaimClosure.query_obstruction_finite_state_core
CC51 DecisionQuotient.ClaimClosure.regime_core_claim_proved
CC52 DecisionQuotient.ClaimClosure.regime_simulation_transfers_hardness
CC53 DecisionQuotient.ClaimClosure.reusable_heuristic_of_detectable
CC54 DecisionQuotient.ClaimClosure.selectorSufficient_not_implies_setSufficient
CC55 DecisionQuotient.ClaimClosure.separable_detectable
CC56 DecisionQuotient.ClaimClosure.snapshot_vs_process_typed_boundary
CC57 DecisionQuotient.ClaimClosure.standard_assumption_ledger_unpack
CC58 DecisionQuotient.ClaimClosure.stochastic_objective_bridge_can_fail_on_sufficiency
CC59 DecisionQuotient.ClaimClosure.subproblem_hardness_lifts_to_full
CC60 DecisionQuotient.ClaimClosure.subproblem_transfer_as_regime_simulation
CC61 DecisionQuotient.ClaimClosure.sufficiency_conp_complete_conditional
CC62 DecisionQuotient.ClaimClosure.sufficiency_conp_reduction_core
CC63 DecisionQuotient.ClaimClosure.sufficiency_iff_dq_ratio
CC64 DecisionQuotient.ClaimClosure.sufficiency_iff_projectedOptCover_eq_opt
CC65 DecisionQuotient.ClaimClosure.thermo_conservation_additive_core
CC66 DecisionQuotient.ClaimClosure.thermo_energy_carbon_lift_core
CC67 DecisionQuotient.ClaimClosure.thermo_eventual_lift_core
CC68 DecisionQuotient.ClaimClosure.thermo_hardness_bundle_core
CC69 DecisionQuotient.ClaimClosure.thermo_mandatory_cost_core
CC70 DecisionQuotient.ClaimClosure.tractable_bounded_core
CC71 DecisionQuotient.ClaimClosure.tractable_separable_core
CC72 DecisionQuotient.ClaimClosure.tractable_subcases_conditional
CC73 DecisionQuotient.ClaimClosure.tractable_tree_core
CC74 DecisionQuotient.ClaimClosure.transition_coupled_bridge_can_fail_on_sufficiency
CC75 DecisionQuotient.ClaimClosure.tree_structure_detectable
CC76 DecisionQuotient.ClaimClosure.typed_claim_admissibility_core
CC77 DecisionQuotient.ClaimClosure.typed_static_class_completeness
CR1 DecisionQuotient.ConfigReduction.config_sufficiency_iff_behavior_preserving
CT1 CT.law_gap_physical_claim_holds
CT2 CT.no_law_gap_counterexample
CT3 CT.no_physical_counterexample_of_core_theorem
CT4 CT.physical_bridge_bundle
CT5 CT.physical_claim_lifts_from_core
CT6 CT.physical_claim_lifts_from_core_conditional
CT7 CT.physical_counterexample_invalidates_core_rule
CT8 CT.physical_counterexample_yields_core_counterexample
DP1 DecisionQuotient.DecisionProblem.minimalSufficient_iff_relevant
DP2 DecisionQuotient.DecisionProblem.relevantSet_is_minimal
DP3 DecisionQuotient.DecisionProblem.sufficient_implies_selectorSufficient
DP4 DecisionQuotient.ClaimClosure.DecisionProblem.epsOpt_zero_eq_opt
DP5 DecisionQuotient.ClaimClosure.DecisionProblem.sufficient_iff_zeroEpsilonSufficient
DP6 DecisionProblem.anchorSufficient
DP7 DecisionProblem.classMonotoneOn
DP8 DecisionProblem.constant_opt_all_sufficient
DP21 DecisionProblem.edgeOnComplement_iff_not_sufficient
DQ3 DecisionQuotient.card_anchoredSlice_eq_pow_sub
DQ5 DecisionQuotient.opt_eq_feasible_of_gap
DQ6 DecisionQuotient.opt_eq_of_allowed_iff
DQ7 DecisionQuotient.opt_singleton_of_logicallyDeterministic
DQ8 DecisionQuotient.universe_sets_objective_schema
DQ154 DecisionQuotient.Physics.Instantiation.DecisionCircuit
DQ155 DecisionQuotient.Physics.Instantiation.DecisionInterpretation
DQ159 DecisionQuotient.Physics.Instantiation.MoleculeAsCircuit
DQ160 DecisionQuotient.Physics.Instantiation.MoleculeAsDecisionCircuit
DQ517 DecisionQuotient.Physics.Instantiation.Circuit
DQ518 DecisionQuotient.Physics.Instantiation.Dynamics
DQ519 DecisionQuotient.Physics.Instantiation.Geometry
DQ533 DecisionQuotient.Physics.DecisionTime.decisionTrace_length_eq_ticks
DQ534 DecisionQuotient.Physics.DecisionTime.decision_count_equals_elapsed_time
DQ535 DecisionQuotient.Physics.DecisionTime.decision_event_iff_eq_tick
DQ537 DecisionQuotient.Physics.DecisionTime.decision_taking_place_is_unit_of_time
DQ539 DecisionQuotient.Physics.DecisionTime.run_elapsed_time_eq_ticks
DQ540 DecisionQuotient.Physics.DecisionTime.run_time_exact
DQ541 DecisionQuotient.Physics.DecisionTime.substrate_step_is_time_unit
DQ542 DecisionQuotient.Physics.DecisionTime.substrate_step_realizes_decision_event
DQ545 DecisionQuotient.Physics.DecisionTime.tick_increments_time
DQ546 DecisionQuotient.Physics.DecisionTime.tick_is_decision_event
DQ547 DecisionQuotient.Physics.DecisionTime.time_coordinate_falsifiable
DQ548 DecisionQuotient.Physics.DecisionTime.time_is_discrete
DQ549 DecisionQuotient.Physics.DecisionTime.time_unit_law_substrate_invariant
DQ560 DecisionQuotient.HardnessDistribution.hardnessLowerBound
DQ561 DecisionQuotient.HardnessDistribution.hardness_is_irreducible_required_work
DQ562 DecisionQuotient.HardnessDistribution.requiredWork
DQ563 DecisionQuotient.HardnessDistribution.requiredWork_eq_affine_in_sites
DQ564 DecisionQuotient.HardnessDistribution.workGrowthDegree
DQ565 DecisionQuotient.HardnessDistribution.workGrowthDegree_zero_iff_eventually_constant
DQ566 DecisionQuotient.Physics.PhysicalCore.admissible_report_requires_strict_certified_work
DQ567 DecisionQuotient.Physics.PhysicalCore.certified_work_lower_bound_raw
DQ568 DecisionQuotient.Physics.PhysicalCore.finite_budget_blocks_canonical_exponential_certification
DQ569 DecisionQuotient.Physics.PhysicalCore.finite_budget_blocks_exact_certification_beyond_threshold
DQ570 DecisionQuotient.Physics.PhysicalCore.rlff_abstain_strict_global_maximizer_of_no_certificates
DQ571 DecisionQuotient.Physics.PhysicalCore.substrate_transport_blocks_encoded_counterexamples
DQ572 DecisionQuotient.Physics.PhysicalCore.substrate_transport_sound_conditional
DQ573 DecisionQuotient.Physics.PhysicalCore.eventual_no_exact_admissibility_under_budget
DQ574 DecisionQuotient.Physics.PhysicalCore.eventual_no_exact_admissibility_under_canonical_requirement
DQ575 DecisionQuotient.PhysicalBudgetCrossover.crossover_eventually_of_eventual_split
DQ576 DecisionQuotient.PhysicalBudgetCrossover.exists_least_crossover_point
DQ577 DecisionQuotient.PhysicalBudgetCrossover.explicit_eventual_infeasibility_of_monotone_and_witness
DQ578 DecisionQuotient.PhysicalBudgetCrossover.no_universal_survivor_without_succinct_bound
DQ579 DecisionQuotient.PhysicalBudgetCrossover.payoff_threshold_explicit_vs_succinct
DQ580 DecisionQuotient.PhysicalBudgetCrossover.policy_closure_at_divergence
DQ581 DecisionQuotient.PhysicalBudgetCrossover.policy_closure_beyond_divergence
DQ582 DecisionQuotient.Physics.PhysicalIncompleteness.under_resolution_implies_collision
DQ583 DecisionQuotient.Physics.PhysicalIncompleteness.under_resolution_implies_decision_collision
DQ584 DecisionQuotient.Physics.HeisenbergStrong.HeisenbergStrongBinding
DQ585 DecisionQuotient.Physics.HeisenbergStrong.strong_binding_implies_core_nontrivial
DQ586 DecisionQuotient.Physics.HeisenbergStrong.strong_binding_implies_nontrivial_opt_via_uncertainty
DQ587 DecisionQuotient.checking_witnessing_duality
DQ588 DecisionQuotient.checking_witnessing_duality_budget
HD1 DecisionQuotient.HardnessDistribution.centralization_dominance_bundle
HD2 DecisionQuotient.HardnessDistribution.centralization_step_saves_n_minus_one
HD3 DecisionQuotient.HardnessDistribution.centralized_higher_leverage
HD4 DecisionQuotient.HardnessDistribution.complete_model_dominates_after_threshold
HD5 DecisionQuotient.HardnessDistribution.gap_conservation_card
HD6 DecisionQuotient.HardnessDistribution.generalizedTotal_with_saturation_eventually_constant
HD7 DecisionQuotient.HardnessDistribution.generalized_dominance_can_fail_without_right_boundedness
HD8 DecisionQuotient.HardnessDistribution.generalized_dominance_can_fail_without_wrong_growth
HD9 DecisionQuotient.HardnessDistribution.generalized_right_dominates_wrong_of_bounded_vs_identity_lower
HD10 DecisionQuotient.HardnessDistribution.generalized_right_eventually_dominates_wrong
HD11 DecisionQuotient.HardnessDistribution.hardnessEfficiency_eq_central_share
HD12 DecisionQuotient.HardnessDistribution.isRightHardness
HD13 DecisionQuotient.HardnessDistribution.isWrongHardness
HD14 DecisionQuotient.HardnessDistribution.linear_lt_exponential_plus_constant_eventually
HD15 DecisionQuotient.HardnessDistribution.native_dominates_manual
HD16 DecisionQuotient.HardnessDistribution.no_positive_slope_linear_represents_saturating
HD17 DecisionQuotient.HardnessDistribution.right_dominates_wrong
HD18 DecisionQuotient.HardnessDistribution.saturatingSiteCost_eventually_constant
HD19 DecisionQuotient.HardnessDistribution.simplicityTax_grows
HD20 DecisionQuotient.HardnessDistribution.totalDOF_eventually_constant_iff_zero_distributed
HD21 DecisionQuotient.HardnessDistribution.totalDOF_ge_intrinsic
IC1 DecisionQuotient.IntegrityCompetence.CertaintyInflation
IC2 DecisionQuotient.IntegrityCompetence.CompletionFractionDefined
IC3 DecisionQuotient.IntegrityCompetence.EvidenceForReport
IC4 DecisionQuotient.IntegrityCompetence.ExactCertaintyInflation
IC5 DecisionQuotient.IntegrityCompetence.Percent
IC6 DecisionQuotient.IntegrityCompetence.RLFFWeights
IC7 DecisionQuotient.IntegrityCompetence.ReportSignal
IC8 DecisionQuotient.IntegrityCompetence.ReportBitModel
IC9 DecisionQuotient.IntegrityCompetence.SignalConsistent
IC10 DecisionQuotient.IntegrityCompetence.admissible_irrational_strictly_more_than_rational
IC11 DecisionQuotient.IntegrityCompetence.admissible_matrix_counts
IC12 DecisionQuotient.IntegrityCompetence.abstain_signal_exists_with_guess_self
IC13 DecisionQuotient.IntegrityCompetence.certaintyInflation_iff_not_admissible
IC14 DecisionQuotient.IntegrityCompetence.certificationOverheadBits
IC15 DecisionQuotient.IntegrityCompetence.certificationOverheadBits_of_evidence
IC16 DecisionQuotient.IntegrityCompetence.certificationOverheadBits_of_no_evidence
IC17 DecisionQuotient.IntegrityCompetence.certifiedTotalBits
IC18 DecisionQuotient.IntegrityCompetence.certifiedTotalBits_ge_raw
IC19 DecisionQuotient.IntegrityCompetence.certifiedTotalBits_gt_raw_of_evidence
IC20 DecisionQuotient.IntegrityCompetence.certifiedTotalBits_of_evidence
IC21 DecisionQuotient.IntegrityCompetence.certifiedTotalBits_of_no_evidence
IC22 DecisionQuotient.IntegrityCompetence.claim_admissible_of_evidence
IC23 DecisionQuotient.IntegrityCompetence.competence_implies_integrity
IC24 DecisionQuotient.IntegrityCompetence.completion_fraction_defined_of_declared_bound
IC25 DecisionQuotient.IntegrityCompetence.epsilon_competence_implies_integrity
IC26 DecisionQuotient.IntegrityCompetence.evidence_nonempty_iff_claim_admissible
IC27 DecisionQuotient.IntegrityCompetence.evidence_of_claim_admissible
IC28 DecisionQuotient.IntegrityCompetence.exact_claim_admissible_iff_exact_evidence_nonempty
IC29 DecisionQuotient.IntegrityCompetence.exact_claim_requires_evidence
IC30 DecisionQuotient.IntegrityCompetence.exactCertaintyInflation_iff_no_exact_competence
IC31 DecisionQuotient.IntegrityCompetence.exact_raw_only_of_no_exact_admissible
IC32 DecisionQuotient.IntegrityCompetence.integrity_forces_abstention
IC33 DecisionQuotient.IntegrityCompetence.integrity_not_competent_of_nonempty_scope
IC34 DecisionQuotient.IntegrityCompetence.integrity_resource_bound
IC35 DecisionQuotient.IntegrityCompetence.no_completion_fraction_without_declared_bound
IC36 DecisionQuotient.IntegrityCompetence.overModelVerdict_rational_iff
IC37 DecisionQuotient.IntegrityCompetence.percentZero
IC38 DecisionQuotient.IntegrityCompetence.rlffBaseReward
IC39 DecisionQuotient.IntegrityCompetence.rlffReward
IC40 DecisionQuotient.IntegrityCompetence.rlff_abstain_strictly_prefers_no_certificates
IC41 DecisionQuotient.IntegrityCompetence.rlff_maximizer_has_evidence
IC42 DecisionQuotient.IntegrityCompetence.rlff_maximizer_is_admissible
IC43 DecisionQuotient.IntegrityCompetence.self_reflected_confidence_not_certification
IC44 DecisionQuotient.IntegrityCompetence.signal_certified_positive_implies_admissible
IC45 DecisionQuotient.IntegrityCompetence.signal_consistent_of_claim_admissible
IC46 DecisionQuotient.IntegrityCompetence.signal_no_evidence_forces_zero_certified
IC47 DecisionQuotient.IntegrityCompetence.signal_exact_no_competence_forces_zero_certified
IV1 DecisionQuotient.InteriorVerification.GoalClass
IV2 DecisionQuotient.InteriorVerification.InteriorDominanceVerifiable
IV3 DecisionQuotient.InteriorVerification.TautologicalSetIdentifiable
IV4 DecisionQuotient.InteriorVerification.agreeOnSet
IV5 DecisionQuotient.InteriorVerification.interiorParetoDominates
IV6 DecisionQuotient.InteriorVerification.interior_certificate_implies_non_rejection
IV7 DecisionQuotient.InteriorVerification.interior_dominance_implies_universal_non_rejection
IV8 DecisionQuotient.InteriorVerification.interior_dominance_not_full_sufficiency
IV9 DecisionQuotient.InteriorVerification.interior_verification_tractable_certificate
L1 ExactPhysicalClaimWellTyped
L2 ExcusedBy
L3 ExplicitHardnessAssumptions
L4 Literal.eval
L5 OneStepSequentialObjective
L6 RegimeSimulation
L7 StandardComplexityAssumptions
L8 StochasticCriterionObjective
L9 StructureDetectable
L10 TransitionCoupledObjective
L11 TwoStepObjective
L12 adq
L13 adq_ordering
L14 agentBridgeClass
L15 agent_transfer_licensed_iff_snapshot
L16 anchor_sigma2p_complete_conditional
L17 anchor_sigma2p_reduction_core
L19 boundaryCharacterized_iff_exists_sufficient_subset
L20 bounded_actions_detectable
L21 bounded_actions_reusable_heuristic
L22 bridgeFailureWitness
L23 bridgeTransferLicensed
L24 bridge_boundary_represented_family
L25 boolHypercube_node_count
L68 node_count_does_not_determine_edge_geometry
PBC1 DecisionQuotient.PhysicalBudgetCrossover.CrossoverAt
PBC2 DecisionQuotient.PhysicalBudgetCrossover.explicit_infeasible_succinct_feasible_of_crossover
PC10 PhysicalComplexity.AccessRegime.AccessRegime
PC12 PhysicalComplexity.AccessRegime.RegimeEval
PC15 PhysicalComplexity.AccessRegime.RegimeWithCertificate
PC17 PhysicalComplexity.AccessRegime.certificate_upgrades_regime
PC19 PhysicalComplexity.AccessRegime.regime_upgrade_with_certificate
PC20 PhysicalComplexity.AccessRegime.AccessChannelLaw
PC21 PhysicalComplexity.AccessRegime.FiveWayMeet
S2P1 DecisionQuotient.Sigma2PHardness.exactlyIdentifiesRelevant_iff_sufficient_and_subset_relevantFinset
S2P2 DecisionQuotient.Sigma2PHardness.min_representationGap_zero_iff_relevant_card
S2P3 DecisionQuotient.Sigma2PHardness.min_sufficient_set_iff_relevant_card
S2P6 DecisionQuotient.Sigma2PHardness.representationGap_eq_zero_iff
A.2 What Is Machine-Checked
The Lean formalization establishes:
  1. 1.
    Correctness of the TAUTOLOGY reduction: The theorem tautology_iff_sufficient proves that the mapping from Boolean formulas to decision problems preserves the decision structure (accept iff the formula is a tautology).
  2. 2.
    Decision problem definitions: Formal definitions of sufficiency, optimality, and the decision quotient.
  3. 3.
    Economic theorems: Simplicity Tax redistribution identities and hardness distribution results.
  4. 4.
    Query-access lower-bound core: Formalized Boolean query-model indistinguishability theorem for the full problem via the $I=\emptyset$ subproblem (emptySufficiency_query_indistinguishable_pair), plus obstruction-scale identities (queryComplexityLowerBound, exponential_query_complexity).
Complexity classifications (coNP-completeness, $\Sigma_{2}^{P}$-completeness) follow by conditional composition with standard results (e.g., TAUTOLOGY coNP-completeness and $\exists\forall$-SAT $\Sigma_{2}^{P}$-completeness), represented explicitly as hypotheses in the conditional transfer theorems listed below. The Lean proofs verify the reduction constructions and the transfer closures under those hypotheses. The assumptions themselves are unpacked by an explicit ledger projection theorem (CC5) so dependency tracking is machine-auditable.
A.3 Assumption Ledger (Auto)
Source files.
  • DecisionQuotient/ClaimClosure.lean
Assumption bundles and fields.
  • (No assumption bundles parsed.)
Conditional closure handles.
  • DQ.anchor_sigma2p_complete_conditional
  • DQ.cost_asymmetry_eth_conditional
  • DQ.dichotomy_conditional
  • DQ.minsuff_collapse_to_conp_conditional
  • DQ.minsuff_conp_complete_conditional
  • DQ.sufficiency_conp_complete_conditional
  • DQ.tractable_subcases_conditional
A.4 Polynomial-Time Reduction Definition
We use Mathlib’s Turing machine framework to define polynomial-time computability:
/-- Polynomial-time computable function using Turing machines -/
def PolyTime {  : Type} (ea : FinEncoding ) (eb : FinEncoding )
    (f :   ) : Prop :=
  Nonempty (Turing.TM2ComputableInPolyTime ea eb f)
/-- Polynomial-time many-one (Karp) reduction -/
def ManyOneReducesPoly {  : Type} (ea : FinEncoding ) (eb : FinEncoding )
    (A : Set ) (B : Set ) : Prop :=
   f :   , PolyTime ea eb f   x, x  A  f x  B
This uses the standard definition: a reduction is polynomial-time computable via Turing machines and preserves membership.
A.5 The Main Reduction Theorem
Theorem A.1 (TAUTOLOGY Reduction Correctness, Lean).
The reduction from TAUTOLOGY to SUFFICIENCY-CHECK is correct:
theorem tautology_iff_sufficient ( : Formula n) :
    .isTautology  (reductionProblem ).isSufficient Finset.empty
This theorem is proven by showing both directions:
  • If $\varphi$ is a tautology, then the empty coordinate set is sufficient
  • If the empty coordinate set is sufficient, then $\varphi$ is a tautology
The proof verifies that the utility construction in reductionProblem creates the appropriate decision structure where:
  • At reference states, accept is optimal with utility 1
  • At assignment states, accept is optimal iff $\varphi(a)=\text{true}$
A.6 Economic Results
The hardness distribution theorems (Section 9) are fully formalized:
theorem simplicityTax_conservation (P : SpecificationProblem)
    (S : SolutionArchitecture P) :
    S.centralDOF + simplicityTax P S  P.intrinsicDOF
theorem simplicityTax_grows (P : SpecificationProblem)
    (S : SolutionArchitecture P) (n n : )
    (hn : n < n) (htax : simplicityTax P S > 0) :
    totalDOF S n < totalDOF S n
theorem native_dominates_manual (P : SpecificationProblem) (n : Nat)
    (hn : n > P.intrinsicDOF) :
    totalDOF (nativeTypeSystem P) n < totalDOF (manualApproach P) n
theorem totalDOF_eventually_constant_iff_zero_distributed
    (S : SolutionArchitecture P) :
    IsEventuallyConstant (fun n => totalDOF S n)  S.distributedDOF = 0
theorem no_positive_slope_linear_represents_saturating
    (c d K : ) (hd : d > 0) :
     ( n, c + n * d = generalizedTotalDOF c (saturatingSiteCost K) n)
Identifier note. Lean identifiers retain internal naming (intrinsicDOF, simplicityTax_conservation); in paper terminology these correspond to baseline hardness and the redistribution lower-bound identity, respectively.
A.7 Complete Claim Coverage Matrix
Paper handle Status Lean support
cor:exact-identifiability Full DQ.Sigma2PHardness.exactlyIdentifiesRelevant_iff_sufficient_and_subset_relevantFinset
cor:exact-no-competence-zero-certified Full DQ.IntegrityCompetence.rlff_maximizer_has_evidence
cor:finite-budget-no-exact-admissibility Full DQ.Physics.PhysicalCore.eventual_no_exact_admissibility_under_budget, DQ.Physics.PhysicalCore.eventual_no_exact_admissibility_under_canonical_requirement
cor:finite-budget-threshold-impossibility Full DQ.Physics.PhysicalCore.eventual_no_exact_admissibility_under_canonical_requirement, DQ.Physics.PhysicalCore.finite_budget_blocks_canonical_exponential_certification, DQ.Physics.PhysicalCore.finite_budget_blocks_exact_certification_beyond_threshold
cor:gap-externalization Full DQ.HardnessDistribution.saturatingSiteCost_eventually_constant, DQ.HardnessDistribution.totalDOF_ge_intrinsic
cor:gap-minimization-hard Full DQ.Sigma2PHardness.min_representationGap_zero_iff_relevant_card
cor:generalized-eventual-dominance Full DQ.HardnessDistribution.generalized_right_dominates_wrong_of_bounded_vs_identity_lower
cor:hardness-exact-certainty-inflation Full DQ.ClaimClosure.exact_certainty_inflation_under_hardness_core
cor:hardness-raw-only-exact Full DQ.ClaimClosure.exact_raw_only_of_no_exact_admissible_core, DQ.IntegrityCompetence.competence_implies_integrity
cor:information-barrier-query Full DQ.ClaimClosure.information_barrier_opt_oracle_core, DQ.ClaimClosure.information_barrier_state_batch_core, DQ.ClaimClosure.information_barrier_value_entry_core
cor:integrity-universal Full DQ.ClaimClosure.integrity_universal_applicability_core
cor:interior-singleton-certificate Full DQ.InteriorVerification.interior_verification_tractable_certificate
cor:linear-positive-no-saturation Full DQ.HardnessDistribution.native_dominates_manual
cor:no-auto-minimize Full DQ.ClaimClosure.no_auto_minimize_of_p_neq_conp
cor:no-uncertified-exact-claim Full DQ.ClaimClosure.no_uncertified_exact_claim_core
cor:no-universal-survivor-no-succinct-bound Full DQ.PhysicalBudgetCrossover.no_universal_survivor_without_succinct_bound
cor:outside-excuses-no-exact-report Full DQ.ClaimClosure.no_exact_claim_under_declared_assumptions_unless_excused_core
cor:overmodel-diagnostic-implication Full DQ.Sigma2PHardness.representationGap_eq_zero_iff
cor:physical-counterexample-core-failure Full CT.no_physical_counterexample_of_core_theorem, CT.physical_counterexample_invalidates_core_rule, CT.physical_counterexample_yields_core_counterexample, DQ.Physics.PhysicalCore.substrate_transport_blocks_encoded_counterexamples
cor:physics-no-universal-exact-claim Full DQ.ClaimClosure.no_exact_claim_admissible_under_hardness_core
cor:practice-bounded Full agent_transfer_licensed_iff_snapshot
cor:practice-diagnostic Full DQ.Sigma2PHardness.min_representationGap_zero_iff_relevant_card
cor:practice-structured Full anchor_sigma2p_complete_conditional
cor:practice-tree Full anchor_sigma2p_reduction_core
cor:practice-unstructured Full ExactPhysicalClaimWellTyped
cor:query-obstruction-bool Full DQ.ClaimClosure.query_obstruction_boolean_corollary, RegimeSimulation
cor:right-wrong-hardness Full DQ.HardnessDistribution.no_positive_slope_linear_represents_saturating
cor:rlff-abstain-no-certs Full DQ.IntegrityCompetence.exactCertaintyInflation_iff_no_exact_competence, DQ.Physics.PhysicalCore.rlff_abstain_strict_global_maximizer_of_no_certificates
cor:type-system-threshold Full DQ.HardnessDistribution.linear_lt_exponential_plus_constant_eventually
prop:abstain-guess-self-signal Full DQ.IntegrityCompetence.signal_no_evidence_forces_zero_certified
prop:adq-ordering Full DQ.ClaimClosure.adq_ordering
prop:attempted-competence-matrix Full DQ.IntegrityCompetence.RLFFWeights, DQ.IntegrityCompetence.ReportSignal, DQ.IntegrityCompetence.evidence_of_claim_admissible
prop:bounded-slice-meta-irrelevance Full DQ.ClaimClosure.meta_coordinate_irrelevant_of_invariance_on_declared_slice, DQ.ClaimClosure.meta_coordinate_not_relevant_on_declared_slice
prop:bridge-failure-horizon Full DQ.ClaimClosure.horizonTwoWitness_immediate_empty_sufficient, DQ.ClaimClosure.horizon_gt_one_bridge_can_fail_on_sufficiency
prop:bridge-failure-stochastic Full DQ.ClaimClosure.standard_assumption_ledger_unpack
prop:bridge-failure-transition Full DQ.ClaimClosure.tractable_tree_core
prop:bridge-transfer-scope Full DQ.ClaimClosure.one_step_bridge
prop:budgeted-crossover Full DQ.ClaimClosure.physical_crossover_core, DQ.PhysicalBudgetCrossover.CrossoverAt
prop:certainty-inflation-iff-inadmissible Full DQ.IntegrityCompetence.ReportBitModel, DQ.IntegrityCompetence.claim_admissible_of_evidence
prop:certified-confidence-gate Full DQ.IntegrityCompetence.rlffBaseReward, DQ.IntegrityCompetence.rlffReward
prop:checking-witnessing-duality Full DQ.checking_witnessing_duality, DQ.checking_witnessing_duality_budget
prop:crossover-above-cap Full DQ.ClaimClosure.physical_crossover_above_cap_core, DQ.PhysicalBudgetCrossover.explicit_infeasible_succinct_feasible_of_crossover
prop:crossover-not-certification Full DQ.ClaimClosure.physical_crossover_hardness_core
prop:crossover-policy Full DQ.ClaimClosure.physical_crossover_policy_core
prop:decision-unit-time Full DQ.Physics.DecisionTime.decision_event_iff_eq_tick, DQ.Physics.DecisionTime.decision_taking_place_is_unit_of_time, DQ.Physics.DecisionTime.tick_increments_time, DQ.Physics.DecisionTime.tick_is_decision_event
prop:declared-contract-selection-validity Full DQ.ClaimClosure.explicit_assumptions_required_of_not_excused_core, DQ.ClaimClosure.no_exact_claim_under_declared_assumptions_unless_excused_core, DQ.ClaimClosure.tree_structure_detectable
prop:dominance-modes Full DQ.HardnessDistribution.centralized_higher_leverage
prop:empty-sufficient-constant Full DP.anchorSufficient
prop:eventual-explicit-infeasibility Full DQ.PhysicalBudgetCrossover.explicit_eventual_infeasibility_of_monotone_and_witness
prop:evidence-admissibility-equivalence Full DQ.IntegrityCompetence.certifiedTotalBits, DQ.IntegrityCompetence.certifiedTotalBits_of_evidence, DQ.IntegrityCompetence.certifiedTotalBits_of_no_evidence
prop:exact-requires-evidence Full DQ.IntegrityCompetence.no_completion_fraction_without_declared_bound, DQ.IntegrityCompetence.overModelVerdict_rational_iff
prop:fraction-defined-under-bound Full DQ.IntegrityCompetence.signal_consistent_of_claim_admissible
prop:generalized-assumption-boundary Full DQ.HardnessDistribution.generalizedTotal_with_saturation_eventually_constant, DQ.HardnessDistribution.generalized_dominance_can_fail_without_right_boundedness
prop:hardness-conservation Full DQ.HardnessDistribution.totalDOF_eventually_constant_iff_zero_distributed
prop:hardness-efficiency-interpretation Full DQ.HardnessDistribution.generalized_right_eventually_dominates_wrong
prop:heisenberg-strong-nontrivial-opt Full DQ.Physics.HeisenbergStrong.HeisenbergStrongBinding, DQ.Physics.HeisenbergStrong.strong_binding_implies_core_nontrivial, DQ.Physics.HeisenbergStrong.strong_binding_implies_nontrivial_opt_via_uncertainty
prop:heuristic-reusability Full DQ.ClaimClosure.bounded_actions_detectable, DQ.ClaimClosure.reusable_heuristic_of_detectable, DQ.ClaimClosure.separable_detectable, DQ.ClaimClosure.transition_coupled_bridge_can_fail_on_sufficiency
prop:insufficiency-counterexample Full DP.classMonotoneOn, DP.constant_opt_all_sufficient
prop:integrity-competence-separation Full DQ.IntegrityCompetence.certifiedTotalBits_ge_raw, DQ.IntegrityCompetence.epsilon_competence_implies_integrity
prop:integrity-resource-bound Full DQ.ClaimClosure.integrity_resource_bound_for_sufficiency, DQ.IntegrityCompetence.completion_fraction_defined_of_declared_bound, DQ.IntegrityCompetence.evidence_nonempty_iff_claim_admissible
prop:interior-one-sidedness Full DQ.InteriorVerification.interior_certificate_implies_non_rejection, DQ.InteriorVerification.interior_dominance_not_full_sufficiency
prop:interior-universal-non-rejection Full DQ.InteriorVerification.interiorParetoDominates
prop:interior-verification-tractable Full DQ.InteriorVerification.agreeOnSet, DQ.InteriorVerification.interior_dominance_implies_universal_non_rejection
prop:law-instance-objective-bridge Full CT.law_gap_physical_claim_holds, CT.no_law_gap_counterexample
prop:least-divergence-point Full DQ.PhysicalBudgetCrossover.exists_least_crossover_point
prop:minimal-relevant-equiv Full DQ.DecisionProblem.relevantSet_is_minimal, DQ.DecisionProblem.sufficient_implies_selectorSufficient
prop:no-evidence-zero-certified Full DQ.IntegrityCompetence.rlff_abstain_strictly_prefers_no_certificates
prop:no-fraction-without-bound Full DQ.IntegrityCompetence.signal_certified_positive_implies_admissible
prop:one-step-bridge Full DQ.ClaimClosure.one_step_bridge
prop:oracle-lattice-strict Full ExplicitHardnessAssumptions, L.eval
prop:oracle-lattice-transfer Full DQ.ClaimClosure.oracle_lattice_transfer_as_regime_simulation, bounded_actions_detectable
prop:outside-excuses-explicit-assumptions Full DQ.ClaimClosure.explicit_assumptions_required_of_not_excused_core
prop:payoff-threshold Full DQ.PhysicalBudgetCrossover.crossover_eventually_of_eventual_split, DQ.PhysicalBudgetCrossover.payoff_threshold_explicit_vs_succinct
prop:physical-claim-transport Full CT.physical_claim_lifts_from_core, CT.physical_claim_lifts_from_core_conditional, DQ.Physics.PhysicalCore.substrate_transport_sound_conditional
prop:physics-no-universal-exact Full DQ.ClaimClosure.declared_physics_no_universal_exact_certifier_core
prop:policy-closure-beyond-divergence Full DQ.PhysicalBudgetCrossover.policy_closure_at_divergence, DQ.PhysicalBudgetCrossover.policy_closure_beyond_divergence
prop:query-finite-state-generalization Full StandardComplexityAssumptions, adq_ordering
prop:query-randomized-robustness Full OneStepSequentialObjective, TwoStepObjective, adq
prop:query-randomized-weighted Full bridgeTransferLicensed, bridge_boundary_represented_family
prop:query-regime-obstruction Full DQ.ClaimClosure.query_obstruction_finite_state_core, StandardComplexityAssumptions, adq_ordering
prop:query-state-batch-lb Full StochasticCriterionObjective, agentBridgeClass
prop:query-subproblem-transfer Full DQ.ClaimClosure.regime_simulation_transfers_hardness, DQ.ClaimClosure.stochastic_objective_bridge_can_fail_on_sufficiency, DQ.ClaimClosure.subproblem_hardness_lifts_to_full
prop:query-tightness-full-scan Full TransitionCoupledObjective
prop:query-value-entry-lb Full StructureDetectable, boundaryCharacterized_iff_exists_sufficient_subset
prop:query-weighted-transfer Full bounded_actions_reusable_heuristic, bridgeFailureWitness
prop:raw-certified-bit-split Full DQ.ClaimClosure.certified_total_bits_split_core, DQ.IntegrityCompetence.admissible_irrational_strictly_more_than_rational, DQ.IntegrityCompetence.admissible_matrix_counts, DQ.IntegrityCompetence.certaintyInflation_iff_not_admissible, DQ.IntegrityCompetence.certificationOverheadBits, DQ.IntegrityCompetence.certificationOverheadBits_of_evidence, DQ.IntegrityCompetence.certificationOverheadBits_of_no_evidence, DQ.Physics.PhysicalCore.certified_work_lower_bound_raw
prop:rlff-maximizer-admissible Full DQ.IntegrityCompetence.exact_raw_only_of_no_exact_admissible, DQ.IntegrityCompetence.integrity_forces_abstention
prop:run-time-accounting Full DQ.Physics.DecisionTime.decisionTrace_length_eq_ticks, DQ.Physics.DecisionTime.decision_count_equals_elapsed_time, DQ.Physics.DecisionTime.run_elapsed_time_eq_ticks, DQ.Physics.DecisionTime.run_time_exact
prop:selector-separation Full DQ.ClaimClosure.selectorSufficient_not_implies_setSufficient
prop:self-confidence-not-certification Full DQ.IntegrityCompetence.signal_exact_no_competence_forces_zero_certified
prop:set-to-selector Full DQ.ClaimClosure.DecisionProblem.sufficient_iff_zeroEpsilonSufficient
prop:snapshot-process-typing Full DQ.ClaimClosure.agent_transfer_licensed_iff_snapshot, DQ.ClaimClosure.process_bridge_failure_witness, DQ.ClaimClosure.snapshot_vs_process_typed_boundary
prop:steps-run-scalar Full DQ.IntegrityCompetence.rlff_maximizer_is_admissible, DQ.IntegrityCompetence.self_reflected_confidence_not_certification
prop:substrate-unit-time Full DQ.Physics.DecisionTime.substrate_step_is_time_unit, DQ.Physics.DecisionTime.substrate_step_realizes_decision_event, DQ.Physics.DecisionTime.time_unit_law_substrate_invariant
prop:sufficiency-char Full DQ.ClaimClosure.sufficiency_conp_reduction_core, DQ.ClaimClosure.sufficiency_iff_dq_ratio
prop:thermo-conservation-additive Full DQ.ClaimClosure.sufficiency_iff_projectedOptCover_eq_opt
prop:thermo-hardness-bundle Full DQ.ClaimClosure.thermo_eventual_lift_core
prop:thermo-lift Full DQ.ClaimClosure.thermo_conservation_additive_core, DQ.ClaimClosure.thermo_energy_carbon_lift_core
prop:thermo-mandatory-cost Full DQ.ClaimClosure.thermo_hardness_bundle_core
prop:time-discrete Full DQ.Physics.DecisionTime.time_coordinate_falsifiable, DQ.Physics.DecisionTime.time_is_discrete
prop:typed-claim-admissibility Full DQ.ClaimClosure.tree_structure_detectable
prop:under-resolution-collision Full DQ.Physics.PhysicalIncompleteness.under_resolution_implies_collision, DQ.Physics.PhysicalIncompleteness.under_resolution_implies_decision_collision
prop:universal-solver-framing Full DQ.ClaimClosure.typed_static_class_completeness
prop:zero-epsilon-competence Full DQ.IntegrityCompetence.certifiedTotalBits_gt_raw_of_evidence, DQ.IntegrityCompetence.integrity_not_competent_of_nonempty_scope
prop:zero-epsilon-reduction Full DQ.ClaimClosure.DecisionProblem.epsOpt_zero_eq_opt, DQ.DecisionProblem.minimalSufficient_iff_relevant
thm:amortization Full DQ.HardnessDistribution.centralized_higher_leverage
thm:bridge-boundary-represented Full DQ.ClaimClosure.bridge_boundary_represented_family, DQ.ClaimClosure.bridge_failure_witness_non_one_step, DQ.ClaimClosure.bridge_transfer_iff_one_step_class
thm:centralization-dominance Full DQ.HardnessDistribution.centralization_dominance_bundle, DQ.HardnessDistribution.centralization_step_saves_n_minus_one
thm:claim-integrity-meta Full DQ.ClaimClosure.explicit_assumptions_required_of_not_excused_core
thm:config-reduction Full DQ.ConfigReduction.config_sufficiency_iff_behavior_preserving
thm:cost-asymmetry-eth Full DQ.ClaimClosure.cost_asymmetry_eth_conditional, DQ.HardnessDistribution.isWrongHardness
thm:dichotomy Full DQ.ClaimClosure.dichotomy_conditional, DQ.ClaimClosure.explicit_state_upper_core, DQ.ClaimClosure.hard_family_all_coords_core
thm:exact-certified-gap-iff-admissible Full DQ.ClaimClosure.epsilon_admissible_iff_raw_lt_certified_total_core, DQ.ClaimClosure.exact_admissible_iff_raw_lt_certified_total_core, DQ.ClaimClosure.exact_raw_eq_certified_iff_certainty_inflation_core, DQ.Physics.PhysicalCore.admissible_report_requires_strict_certified_work
thm:generalized-dominance Full DQ.HardnessDistribution.generalized_dominance_can_fail_without_wrong_growth
thm:generalized-saturation-possible Full DQ.HardnessDistribution.gap_conservation_card, DQ.HardnessDistribution.right_dominates_wrong
thm:linear-saturation-iff-zero Full DQ.HardnessDistribution.simplicityTax_grows
thm:overmodel-diagnostic Full DQ.ClaimClosure.boundaryCharacterized_iff_exists_sufficient_subset, DQ.ClaimClosure.no_exact_identifier_implies_not_boundary_characterized
thm:physical-bridge-bundle Full CT.physical_bridge_bundle
thm:regime-coverage Full DQ.ClaimClosure.declaredRegimeFamily_complete, DQ.ClaimClosure.regime_core_claim_proved
thm:tax-conservation Full DQ.HardnessDistribution.complete_model_dominates_after_threshold
thm:tax-grows Full DQ.HardnessDistribution.totalDOF_ge_intrinsic
thm:tractable Full DQ.ClaimClosure.thermo_mandatory_cost_core, DQ.ClaimClosure.tractable_bounded_core, DQ.ClaimClosure.tractable_separable_core, DQ.ClaimClosure.tractable_subcases_conditional
thm:typed-completeness-static Full DQ.ClaimClosure.typed_claim_admissibility_core
Auto summary: mapped 123/123 (full=123, derived=0, unmapped=0).
A.8 Proof Hardness Index (Auto)
Paper handle Hardness profile Regime tags Lean support
cor:gap-externalization cost-growth H=cost-growth HD18, HD21
cor:generalized-eventual-dominance cost-growth H=cost-growth HD9
cor:linear-positive-no-saturation cost-growth H=cost-growth HD15
cor:right-wrong-hardness cost-growth H=cost-growth HD16
cor:type-system-threshold cost-growth H=cost-growth HD14
prop:dominance-modes cost-growth H=cost-growth HD3
prop:generalized-assumption-boundary cost-growth H=cost-growth HD6, HD7
prop:hardness-conservation cost-growth H=cost-growth HD20
prop:hardness-efficiency-interpretation cost-growth H=cost-growth HD10
prop:thermo-conservation-additive cost-growth H=cost-growth CC64
prop:thermo-hardness-bundle cost-growth H=cost-growth CC67
prop:thermo-lift cost-growth H=cost-growth CC65, CC66
prop:thermo-mandatory-cost cost-growth H=cost-growth CC68
thm:generalized-dominance cost-growth H=cost-growth HD8
thm:generalized-saturation-possible cost-growth H=cost-growth HD5, HD17
thm:linear-saturation-iff-zero cost-growth H=cost-growth HD19
cor:finite-budget-no-exact-admissibility exp-lb-conditional H=exp-lb-conditional DQ573, DQ574
cor:finite-budget-threshold-impossibility exp-lb-conditional H=exp-lb-conditional DQ574, DQ568, DQ569
cor:no-universal-survivor-no-succinct-bound exp-lb-conditional H=exp-lb-conditional DQ578
prop:budgeted-crossover exp-lb-conditional H=exp-lb-conditional CC45, PBC1
prop:crossover-above-cap exp-lb-conditional H=exp-lb-conditional CC44, PBC2
prop:crossover-not-certification exp-lb-conditional H=exp-lb-conditional CC46
prop:crossover-policy exp-lb-conditional H=exp-lb-conditional CC47
prop:eventual-explicit-infeasibility exp-lb-conditional H=exp-lb-conditional DQ577
prop:least-divergence-point exp-lb-conditional H=exp-lb-conditional DQ576
prop:payoff-threshold exp-lb-conditional H=exp-lb-conditional DQ575, DQ579
prop:policy-closure-beyond-divergence exp-lb-conditional H=exp-lb-conditional DQ580, DQ581
cor:information-barrier-query query-lb H=query-lb CC27, CC28, CC29
cor:query-obstruction-bool query-lb H=query-lb CC49, L6
prop:checking-witnessing-duality query-lb H=query-lb DQ587, DQ588
prop:query-finite-state-generalization query-lb H=query-lb L7, L13
prop:query-randomized-robustness query-lb H=query-lb L5, L11, L12
prop:query-randomized-weighted query-lb H=query-lb L23, L24
prop:query-tightness-full-scan query-lb H=query-lb L10
prop:query-weighted-transfer query-lb H=query-lb L21, L22
cor:exact-identifiability succinct-hard H=succinct-hard S2P1
cor:gap-minimization-hard succinct-hard H=succinct-hard S2P2
cor:no-auto-minimize succinct-hard H=succinct-hard CC37
cor:overmodel-diagnostic-implication succinct-hard H=succinct-hard S2P6
cor:practice-diagnostic succinct-hard H=succinct-hard S2P2
prop:sufficiency-char succinct-hard H=succinct-hard CC62, CC63
cor:interior-singleton-certificate tractable-structured H=tractable-structured IV9
cor:practice-bounded tractable-structured H=tractable-structured L15
cor:practice-structured tractable-structured H=tractable-structured L16
cor:practice-tree tractable-structured H=tractable-structured L17
prop:heuristic-reusability tractable-structured H=tractable-structured CC7, CC53, CC55, CC74
prop:interior-verification-tractable tractable-structured H=tractable-structured IV4, IV7
cor:exact-no-competence-zero-certified unspecified AR,DC IC41
cor:hardness-exact-certainty-inflation unspecified CR CC19
cor:hardness-raw-only-exact unspecified AR CC21, IC23
cor:integrity-universal unspecified TR CC31
cor:no-uncertified-exact-claim unspecified AR CC41
cor:outside-excuses-no-exact-report unspecified CR,DC CC39
cor:physical-counterexample-core-failure unspecified AR,E,Qb,Qf,S,S+ETH CT3, CT7, CT8, DQ571
cor:physics-no-universal-exact-claim unspecified AR,CR,DC CC38
cor:practice-unstructured unspecified - L1
cor:rlff-abstain-no-certs unspecified AR IC30, DQ570
prop:abstain-guess-self-signal unspecified AR,DC IC46
prop:adq-ordering unspecified - CC2
prop:attempted-competence-matrix unspecified AR IC6, IC7, IC27
prop:bounded-slice-meta-irrelevance unspecified AR CC32, CC33
prop:bridge-failure-horizon unspecified - CC25, CC26
prop:bridge-failure-stochastic unspecified - CC57
prop:bridge-failure-transition unspecified - CC73
prop:bridge-transfer-scope unspecified - CC42
prop:certainty-inflation-iff-inadmissible unspecified CR IC8, IC22
prop:certified-confidence-gate unspecified AR,DC IC38, IC39
prop:decision-unit-time unspecified AR,E,Qb,Qf,S,S+ETH DQ535, DQ537, DQ545, DQ546
prop:declared-contract-selection-validity unspecified - CC22, CC39, CC75
prop:empty-sufficient-constant unspecified AR,DC,DM DP6
prop:evidence-admissibility-equivalence unspecified AR,DC IC17, IC20, IC21
prop:exact-requires-evidence unspecified AR,DC IC35, IC36
prop:fraction-defined-under-bound unspecified AR,DC IC45
prop:heisenberg-strong-nontrivial-opt unspecified AR DQ584, DQ585, DQ586
prop:insufficiency-counterexample unspecified AR,DC,DM DP7, DP8
prop:integrity-competence-separation unspecified AR IC18, IC25
prop:integrity-resource-bound unspecified AR,DC,S,S+ETH CC30, IC24, IC26
prop:interior-one-sidedness unspecified AR IV6, IV8
prop:interior-universal-non-rejection unspecified AR IV5
prop:law-instance-objective-bridge unspecified AR CT1, CT2
prop:minimal-relevant-equiv unspecified - DP2, DP3
prop:no-evidence-zero-certified unspecified AR,DC IC40
prop:no-fraction-without-bound unspecified AR,DC IC44
prop:one-step-bridge unspecified - CC42
prop:oracle-lattice-strict unspecified AR,E,Qb,Qf,S,S+ETH L3, L4
prop:oracle-lattice-transfer unspecified AR,E,Qb,Qf,S,S+ETH CC43, L20
prop:outside-excuses-explicit-assumptions unspecified CR,DC CC22
prop:physical-claim-transport unspecified AR,E,LC,Qb,Qf,S,S+ETH,TA CT5, CT6, DQ572
prop:physics-no-universal-exact unspecified AR,CR,DC CC15
prop:query-regime-obstruction unspecified AR,E,Q_fin,Qb,Qf,S,S+ETH CC50, L7, L13
prop:query-state-batch-lb unspecified AR,E,Qb,Qf,S,S+ETH L8, L14
prop:query-subproblem-transfer unspecified AR,E,Qb,Qf,S,S+ETH CC52, CC58, CC59
prop:query-value-entry-lb unspecified AR,E,Qb,Qf,S,S+ETH L9, L19
prop:raw-certified-bit-split unspecified AR CC11, IC10, IC11, IC13, IC14, IC15, IC16, DQ567
prop:rlff-maximizer-admissible unspecified AR IC31, IC32
prop:run-time-accounting unspecified AR,E,Qb,Qf,S,S+ETH DQ533, DQ534, DQ539, DQ540
prop:selector-separation unspecified - CC54
prop:self-confidence-not-certification unspecified AR,DC IC47
prop:set-to-selector unspecified - DP5
prop:snapshot-process-typing unspecified RA CC3, CC48, CC56
prop:steps-run-scalar unspecified AR,DC IC42, IC43
prop:substrate-unit-time unspecified AR,E,Qb,Qf,S,S+ETH DQ541, DQ542, DQ549
prop:time-discrete unspecified AR,E,Qb,Qf,S,S+ETH DQ547, DQ548
prop:typed-claim-admissibility unspecified AR,DC,Qf,S+ETH CC75
prop:under-resolution-collision unspecified DM DQ582, DQ583
prop:universal-solver-framing unspecified TR CC77
prop:zero-epsilon-competence unspecified - IC19, IC33
prop:zero-epsilon-reduction unspecified - DP4, DP1
thm:amortization unspecified LC HD3
thm:bridge-boundary-represented unspecified - CC8, CC9, CC10
thm:centralization-dominance unspecified AR,LC HD1, HD2
thm:claim-integrity-meta unspecified AR,CR,DC CC22
thm:config-reduction unspecified AR,E,S,TA CR1
thm:cost-asymmetry-eth unspecified S+ETH CC12, HD13
thm:dichotomy unspecified AR,E,Q_fin,Qb,Qf,RG,S,S+ETH CC16, CC23, CC24
thm:exact-certified-gap-iff-admissible unspecified AR CC17, CC18, CC20, DQ566
thm:overmodel-diagnostic unspecified S CC6, CC40
thm:physical-bridge-bundle unspecified AR CT4
thm:regime-coverage unspecified CR,S CC14, CC51
thm:tax-conservation unspecified AR,LC HD4
thm:tax-grows unspecified LC HD21
thm:tractable unspecified AR,E,Qb,Qf,RG,S,S+ETH,STR,TA CC69, CC70, CC71, CC72
thm:typed-completeness-static unspecified AR,CR,S CC76
Auto summary: indexed 123 claims by hardness profile (cost-growth=16; exp-lb-conditional=11; query-lb=8; succinct-hard=6; tractable-structured=6; unspecified=76).
A.9 Claims Not Fully Mechanized
Status: all theorem/proposition/corollary handles in this paper now have Lean backing. Entries marked Full (conditional) are explicitly mechanized transfer theorems that depend on standard external complexity facts (e.g., source-class completeness or ETH assumptions), with those dependencies represented as hypotheses in Lean.
A.10 Module Structure
  • Basic.lean – Core definitions (DecisionProblem, sufficiency, optimality)
  • Sufficiency.lean – Sufficiency checking algorithms and properties
  • Reduction.lean – TAUTOLOGY reduction construction and correctness
  • Hardness/ConfigReduction.lean – Sentinel-action configuration reduction theorem
  • Complexity.lean – Polynomial-time reduction definitions using mathlib
  • HardnessDistribution.lean – Simplicity Tax redistribution and amortization theorems
  • IntegrityCompetence.lean – Solver integrity vs regime competence separation
  • ClaimClosure.lean – Mechanized closure of paper-level bridge and diagnostic claims
  • Tractability/ – Bounded actions, separable utilities, tree structure
A.11 Verification
The proofs compile with Lean 4 and contain no sorry placeholders. Run lake build in the proof directory to verify.